### v5

# Intensity-Efficient Allocations

Abstract
This paper proposes a refinement of Pareto-efficient allocations for situations where, in addition to having ordinal preferences, agents also have ordinal intensities: they can make comparisons such as “I prefer $a$ to $b$ more than I prefer $c$ to $d$, without necessarily being able to quantify them. A rank-based criterion for interpersonal comparisons of such ordinal intensities is introduced for this new analytical environment. Building on this, an allocation is defined to be intensity-efficient if it is Pareto-efficient with respect to the agents’ preferences and also such that when another Pareto-efficient allocation assigns the same pairs of items to the same pairs of agents but in a “flipped” way, the former allocation assigns the commonly preferred item in every such pair to the agent who prefers it more. Conditions are established under which such Pareto-refining allocations exist. The potential usefulness of this theory in matching problems is illustrated with a quadratic-time extension of the Random Priority (RP) algorithm that returns an allocation which intensity-dominates RP’s Pareto-efficient one.
“Suppose I am left with a ticket to a Mozart concert I am unable to attend and decide to give it to one of my closest friends. Which friend should I actually give it to? One thing I will surely consider in deciding this is which friend of mine would enjoy the concert most.”
John C. Harsanyi (1987)
“The problem I have with utilitarianism is not that it is excessively rational, but that the epistemological foundations are weak. My problem is: What are those objects we are adding up? I have no objection to adding them up if there’s something to add.”
Kenneth J. Arrow (1987)
1 Introduction
Distributional principles that are used in the assignment of goods to individuals can be evaluated jointly on the realism of their assumptions and the appeal—according to some justice criterion—of their prescribed allocations. Arguably the most important such economic principle is Pareto-efficiency: assuming only information on agents’ preference rankings over the relevant goods, its notion of optimality is defined by the requirement that no changes can make some individuals better off without hurting others. Yet it is well-known that among the typically many Pareto-efficient allocations there are often some which violate intuitive notions of distributive justice. Appreciation of these facts raises the question of whether one can narrow down and improve upon Pareto’s criterion by considering alternative principles which, in addition to agents’ preferences, also incorporate information on their preference intensities.
Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities. Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered. Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion formalized in Harsanyi (1955) is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Roemer, 1996; Moulin, 2004).
Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Dietrich, 2025; Samuelson, 2025). More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982). This stands in contrast to some contemporary views in the field of cognitive psychology and decision-making, according to which “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” (Vlaev et al., 2011, p. 551). To the extent that comparing options is easier and more reliable than assigning values to them, as the above survey and several studies cited therein suggest, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made. Specifically, instead of cardinal-welfarist or Arrowian profiles that, respectively consist of interpersonally comparable cardinal utility functions—possibly not arising from expected-utility preferences (Nebel, 2024; Weymark, 2005; Harvey, 1999)—or preference orderings, this domain comprises profiles of generally ordinal preference intensity orderings, i.e. collections of quaternary relations that reflect the agents’ preference intensity comparisons. This analytical choice assumes (and this assumption is validated in Section 5, Proposition 3) that, similar to the way in which preferences are elicited, e.g. in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
This new and general framework where both preference and preference-intensity information is available to the social planner naturally raises the question of how such information might be used—as per the goals that we set out above—to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ preference intensities and mitigate the multiplicity of Pareto-efficient allocations. Similar to utilitarianism and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons. Unlike those notions, however, where such comparisons are based on agents’ utilities (Baccelli, 2023; Echenique et al., 2023; d’Aspremont, 2018; Fleurbaey and Hammond, 2004), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful. In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more. In Section 3 we formalize and defend this Interpersonal Comparisons of Ordinal Intensities new postulate/informational basis for social choice and welfare. We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (Baccelli, 2023; d’Aspremont, 2018).
In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just. Specifically, we define an allocation $x$ to be intensity-efficient if it is Pareto-efficient and also such that, whenever another Pareto-efficient allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more. Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile, without further restrictions, when there are three agents and alternatives. Yet we also show that with four or more existence is not guaranteed because the underlying dominance relation may be cyclic without further assumptions. For the case of $n$ agents and items we then establish a sufficient condition which requires profiles to be (effectively) homeo-monotonic and whose key part could be summarized informally thus: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any two consecutively-ranked items in this subset, then this ordering cannot be violated in how much they prefer any non-consecutively ranked items. Furthermore, although this condition is not necessary for existence, we prove that a weaker version of it is indeed satisfied by every profile that has an intensity-efficient allocation.
Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible. Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide, although they do so under an intuitive additional condition that we introduce. The potential discrepancy between these two intensity-incorporating welfare concepts that may arise in such cardinal environments is itself of potential usefulness in applications, because either notion can be used as a secondary selection criterion in those situations where the other one predicts multiple allocations.
In Section 5, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon Pareto-efficiency in the direction of intensity-efficiency. We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited efficiently (specifically, in quadratic time) by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class we study. Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) one and produces—in quadratic time in the number of items/agents—an allocation that intensity-dominates the Pareto-efficient allocation that is found by RP. While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm. We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and which focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
1.1 Related Literature
This paper’s contribution lies at the intersection of welfare economics, social choice and matching theory. Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Kaneko and Nakamura, 1979; Roemer, 1996; Moulin, 2004), Pareto-efficient egalitarian-equivalence (Pazner and Schmeidler, 1978) or relative egalitarianism (Kalai and Smorodinsky, 1975; Imai, 1983; Sprumont, 2013), the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process. Unlike these refinements, the hereby proposed one is the first that is operational in the general class of ordinal preference-intensity profiles, and builds on a novel interpersonal-comparability criterion that is distinct from interpersonal comparisons of utility differences, which the above notions rely on. Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain. The latter problem is important and deserves separate treatment.
Intensity-efficient allocations and their domain of application are also distinct from those dictated by the Borda scoring rule (Borda, 1781; Young, 1974; Maskin, 2025). Although this rule is defined by a particular normalization of agents’ ordinal utility functions over alternatives, in practice it is often used with a cardinal interpretation due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where every agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart. The class of ordinal intensity relations that we consider does not impose this restriction. Yet at the same time, for the main results of this paper we assume a strictness condition on agents’ intensities which rules out such “linear” intensity profiles.
In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs in the allocation process toward maximizing utilitarian welfare. Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer. While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
2 Notation, Preliminaries and Measurement-Theoretic Foundations
By $A:=\{a_{1},\ldots,a_{n}\}$ and $V:=\{v_{1},\ldots,v_{n}\}$ we denote, respectively, the finite sets of choice alternatives and agents. To simplify the assignment problem under study we assume that these sets have equal sizes. When subscripts are unnecessary we also write $a,b\in A$. Similarly, when no confusion arises we will refer to agent $v_{l}$ by their subscript, $l$. The preference intensity relation of agent $l\leq n$ on $A$ is denoted by $\dot{\succsim}_{l}$, with $\dot{\succ}_{l}$ and $\dot{\sim}_{l}$ its asymmetric and symmetric parts. In line with extensive-measurement theory (Krantz et al., 1971; Pfanzagl, 1971; Roberts, 1979), these are quaternary relations on $A$ or, equivalently, binary relations on $A\times A$. The statement $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ will be interpreted as “$a$ is preferred to $b$ at least as much as $c$ is to dd” when the first option in each pair is the (possibly weakly) preferred one at that pair, and as “$b$ is preferred to $a$ no more than $d$ is preferred to cc” when the converse is true (naturally, $(a,b)\dot{\succsim}_{l}(c,d)$ will also hold when $a$ is preferred to $b$ and $d$ is preferred to $c$). When no ambiguity arises, we will interpret $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ more succinctly as suggesting that the intensity difference between $a$ and $b$ is weakly larger than that between $c$ and $d$.
In line with these interpretations and the literature, agent $l$’s binary preference relation $\succsim_{l}$ on $A$ is derived from their intensity relation $\dot{\succsim}_{l}$ by $a\succsim_{l}b\Leftrightarrow(a,b)\,\dot{\succsim}_{l}\,(b,a)$. We will refer to $\succsim_{l}$ as the preference relation that is induced by $\dot{\succsim}_{l}$. As usual, the asymmetric and symmetric parts of $\succsim_{l}$ are denoted by $\succ_{l}$ and $\sim_{l}$. Under the structure that $\dot{\succsim}_{l}$ will be endowed with below, both this relation and its induced $\succsim_{l}$ are weak orders on their respective domains.
For $a,b\in A$, the intensity-equivalence class of $(a,b)\in A\times A$ under $\dot{\succsim}_{l}$ is defined by
$[a,b]$ $:=$ $\{(a^{\prime},b^{\prime})\in A\times A:(a^{\prime},b^{\prime})\,\dot{\sim}_{l}\,(a,b)\}$ (1)
Furthermore, the quotient set of $A\times A$ under $\dot{\sim}_{l}$ is defined by
$(A\times A)_{\dot{\sim}_{l}}$ $:=$ $\{[a,b]:[a,b]\text{ is an intensity-equivalence class under }\dot{\sim}_{l}\}$ (2)
The definitions that follow omit existential quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
Definition 1.
A relation $\dot{\succsim}_{l}$ has a utility-difference representation if there is a function $u_{l}:A\rightarrow\mathbb{R}$ that satisfies
$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $u_{l}(a)-u_{l}(b)\geq u_{l}(c)-u_{l}(d)$ (3)
In an influential article, psychologist S. S. Stevens (1946) distinguished between four scales of measurement that are afforded by some numerical assignment on a set of items: nominal (qualitative/categorical measurement); ordinal (invariance up to a strictly increasing transformation); interval (invariance up to positive affine transformation); and ratio (invariance up to a positive linear transformation). Utility-difference representations were added to this list by psychologist C. H. Coombs (1950) as those that correspond to ordered metric scale measurement. These are more general than neoclassical cardinal utility representations (Lange, 1934; Alt, 1936, 1971; Samuelson, 1938) because $u_{l}$ in the latter case is an interval scale whereas in the former it “falls logically between an interval scale and an ordinal scale” (p. 145). Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a subset of utility-difference comparisons whose order is indeed preserved by all ordinal transformations of the original representation. In our environment, whenever $u_{l}$ in (1) is not cardinal it has the complex uniqueness property of additive representations on finite sets.
Scott (1964), completing and extending the analysis that was initiated in Scott and Suppes (1958), characterized the existence of a utility-difference representation over a finite set by means of the following three axioms on $\dot{\succsim}_{l}$:
$(\mathbf{1_{\mathit{D}}})$ [Completeness]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ or $(c,d)\,\dot{\succsim}_{l}\,(a,b)$.
$(\mathbf{2_{\mathit{D}}})$ [Cancellation]. $(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
$(\mathbf{3_{\mathit{D}}})$ [Reversal]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ implies $(d,c)\,\dot{\succsim}_{l}\,(b,a)$.
The first imposes universal comparability on $\dot{\succsim}_{l}$. The third is essential for $\dot{\succsim}_{l}$ to be interpreted as a preference intensity relation rather than as an arbitrary preference relation over pairs. The second is known to have “no simple way of summarizing what it says about choice behavior” (Luce and Suppes, 1965, p. 277). Scott (1964, p. 244-5) explained how $(\mathbf{2_{\mathit{D}}})$ implies transitivity of $\dot{\succsim}_{l}$, but also remarked that $(\mathbf{2_{\mathit{D}}})$ is an infinite bundle of conditions (for each $t>0$, each $\pi,\sigma$), and it was shown in Scott and Suppes (1958) that no finite number of them could be sufficient”. Once the number $n$ of alternatives in $A$ is fixed, there are obviously finitely many such conditions to be satisfied. Scott’s remark, however, clarifies that it is impossible to reduce this condition to a finite number of statements/axioms that, together with $(\mathbf{1_{\mathit{D}}})$ and $(\mathbf{3_{\mathit{D}}})$, would be sufficient to characterize (3) for any finite $n$.
In Gerasimou (2021, 2022) this author recently proposed and characterized the class of ordinal intensity relations that admit the following more general numerical representation.
Definition 2.
A relation $\dot{\succsim}_{l}$ is representable by a preference intensity function if there is a mapping $s_{l}:A\times A\rightarrow\mathbb{R}$, unique up to a strictly increasing transformation, such that
$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(c,d)$ (4)
$(a,b)\,\dot{\succsim}_{l}\,(b,a)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(e,e)\geq s_{l}(b,a)$ (5)
$\min\{s_{l}(a,b),s_{l}(b,c)\}\geq s_{l}(e,e)$ $\Longrightarrow$ $s_{l}(a,c)\geq\max\{s_{l}(a,b),s_{l}(b,c)\},$ (6)
with a strict inequality on the left hand side of (6) implying a strict inequality on the right.
In words, $s_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)]. Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
$(\mathbf{4_{\mathit{D}}})$ [Translocation Consistency]. $(a,c)\,\dot{\succsim}_{l}\,(b,c)$ if and only if $(a,b)\,\dot{\succsim}_{l}\,(b,a)$.
We denote by $\mathcal{I}$ the class of intensity relations that admit this ordinal representation.
The (pseudo-)cardinal utility-difference model is nested by the ordinal model in the special case where “lateral consistency” [(6)] is strengthened to “additivity” [(7)]:
Lemma 1.
A relation $\dot{\succsim}_{l}$ that is represented by $s_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widehat{s}_{l}$ of $s_{l}$ satisfies
$\widehat{s}_{l}(a,c)$ $=$ $\widehat{s}_{l}(a,b)+\widehat{s}_{l}(b,c)$ (7)
Proof.
Upon viewing $\widehat{s}_{l}$ as a general function of two variables, (7) is known as Sincov’s functional equation, whose solution is some $f:A\rightarrow\mathbb{R}$, unique up to an additive constant, such that $\widehat{s}_{l}(a,b)\equiv f(a)-f(b)$ (Aczél, 1966). This establishes the “if” part. The “only if” part is immediate upon defining $\widehat{s}_{l}(a,b):=u_{l}(a)-u_{l}(b)$ for the postulated function $u_{l}$ with the requisite representation property. ∎
The normalization of $s_{l}$ introduced next is important for the rest of this paper.
Definition 3.
A preference intensity function $s_{l}$ represents $\dot{\succsim}_{l}$ canonically if
$s_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
with $k$ the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
Lemma 2.
Any relation $\dot{\succsim}_{l}\in\mathcal{I}$ admits a canonical preference intensity representation.
Proof.
Let $s_{l}$ be an arbitrary preference intensity function for $\dot{\succsim}_{l}\in\mathcal{I}$. Define the level set of $s_{l}$ at $(a,b)\in A\times A$ by $\mathbf{s}_{l}[a,b]:=\{(a^{\prime},b^{\prime})\in A\times A:s_{l}(a,b)=s_{l}(a^{\prime},b^{\prime})\}$. Further, define $r_{l}:A\times A\rightarrow\mathbb{R}$ by
$r_{l}(a,b)$ $:=$ $\left\{\begin{array}[]{ll}|\{\mathbf{s}_{l}[a^{\prime},b^{\prime}]:s_{l}(a,b)>s_{l}(a^{\prime},b^{\prime})\geq s_{l}(e,e)\}|,&\text{if }s_{l}(a,b)>s_{l}(e,e)\\ s_{l}(a,b),&\text{if }s_{l}(a,b)=s_{l}(e,e)\\ -r_{l}(b,a),&\text{if }s_{l}(a,b)<s_{l}(e,e)\end{array}\right$.
By construction, $r_{l}$ is canonical and $r_{l}(a,b)\geq r_{l}(c,d)\Leftrightarrow s_{l}(a,b)\geq s_{l}(c,d)$. ∎
We assume throughout that every $\dot{\succsim}_{l}$ belongs to $\mathcal{I}$, hence is representable canonically. Our subsequent analysis will further impose the following additional property on $\dot{\succsim}_{l}$:
$(\mathbf{5_{\mathit{D}}})$ [Strictness]. $(a,b)\,\mathord{\not\mathrel{\dot{\sim}_{l}}}\,(c,d)$ for all distinct off-diagonal pairs $(a,b),(c,d)\in A\times A$.
This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$. It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies.
As we elaborate in the next section, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$. In particular, the only non-trivial intensity-equivalence class that exists when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
Table 1: The number of distinct intensity relations of different classes for small values of $n$.
Strict & utility-difference Strict Strict ordinal
Strict representable ordinally representable intensity
$n=$ orders intensity relations intensity relations profiles
3 6 12 12 $12^{3}$
4 24 240 384 $384^{4}$
5 120 13,680 92,160 $92,160^{5}$
Finally, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit, respectively, a (pseudo-cardinal) utility-difference and an ordinal representation when $n\leq 5$, and was produced with Minizinc (Nethercote et al., 2007; Gecode Team, 2006).
3 The Comparability Postulate
[…]
Interpersonal Comparisons of Ordinal Intensities
Given an intensity profile $S\in\mathcal{\widehat{S}}$ and the corresponding canonical profile $s$, the statement
$s_{l}(a,b)$ $>$ $s_{m}(c,d)\;>\;0$ (10)
is interpreted as suggesting that agent $l$ prefers $a$ to $b$ more than agent $m$ prefers $c$ to $d$.
Towards defending this new concept we first recall that preference intensities at the level of the individual decision maker are not assumed here to be quantifiable beyond the level of an ordinal ranking. Yet some information about the generally different welfare effects of having $a$ or $b$ vs $c$ or $d$ is available here. Suppose, in particular, that agent $l$ prefers $a$ to $b$ to $c$ to $d$, and the first over the second more than the third over the fourth. This is equivalent to saying that her intensity difference at pair $(a,b)$ is greater than at $(c,d)$. Consider agent $m$ next, and suppose for simplicity that the same applies to him too: $s_{m}(a,b)>s_{m}(c,d)$. Considering this, and that both have preferences and intensities over the same set of alternatives, how might (10) be interpreted?
It is clearly a statement of fact that the intensity difference within pair $(a,b)$ is ranked higher by agent $l$ in her intensity ordering than the corresponding difference within pair $(c,d)$ is ranked by agent $m$ in his. Now recall that, by the Strictness assumption, $l$ and $m$ have the same number of possible rank positions in their respective intensity orderings which pairs $(a,b)$ and $(c,d)$ can occupy. In other words, the agents’ ordinal intensity scales coincide: if $(a,b)$ and $(c,d)$ are ranked 5th and 8th in $l$’s and $m$’s scales, respectively, then they are ranked 5th and 8th out of the same $k$ possible positions. As we mentioned in the Introduction, this logic resembles the way scoring rules such as the well-known Borda-count (Borda, 1781) normalization of ordinal utilities or scoring-like (see d’Aspremont, 2018) rules such as the relative-utilitarian (Dhillon and Mertens, 1999) normalization of cardinal utilities perform interpersonal comparisons in their respective environments.
The question now emerges: Should the different agents’ intensity orderings be treated equally by the social planner? While equal treatment may or may not be an appropriate approach to follow in practice, depending on what else is known about agents and the problem at hand, since the intensity orderings are assumed here to encode all available welfare-relevant information, treating them in any way other than equal would call for a justification that appears elusive. Against this background, we view the suggested interpretation of (10) as a reasonable starting point for interpersonal comparisons in this informational environment.
On the analytical side, another relevant question also arises: in the special case where the ordinal intensity ordering of every agent is defined by how the pairwise differences of a cardinal utility function are ordered, and these differences in turn define a strict intensity profile, does interpersonal comparisons of ordinal intensities reduce to the interpersonal comparability of utility differences that is required by utilitarian aggregation? As the next example shows, the answer is no.
Example 1.
Let $n=3$ and consider the utility profile $U:=\big{(}(53,38,9),(46,33,21),$ $(20,30,50)\bigr{)}$ over $A=\{a_{1},a_{2},a_{3}\}$, where each $u_{l}$ in $U$, $l\leq 3$, is constructed so its values add up to 100 (this is without loss). It is readily seen that the intensity profile defined by the intra-personal utility differences in $U$ is in $\mathcal{\widehat{S}}$. Also, defining $\dot{\succ}_{i}$ by $(a_{i},a_{j})\dot{\succ}_{l}(a_{k},a_{m})$ $\Leftrightarrow$ $u_{l}(a_{i})-u_{l}(a_{j})>u_{l}(a_{k})-u_{l}(a_{m})$, one observes that $(a_{2},a_{3})\,\dot{\succ}_{1}\,(a_{1},a_{2})$ and $(a_{1},a_{2})\,\dot{\succ}_{2}\,(a_{2},a_{3})$ hold, hence $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$ too. However, $u_{1}(a_{1})-u_{1}(a_{2})>u_{2}(a_{1})-u_{2}(a_{2})$ is also true. $\blacklozenge$
Thus, even if there is some ordinal transformation $\widehat{s}_{i}$ of the canonical intensity function $s_{i}$ such that $\widehat{s}_{i}(a,b)\equiv u_{i}(a)-u_{i}(b)$ for some cardinally unique $u_{i}$ [cf. (7)], the above clarifies that (10) neither implies nor is implied by the inequality $u_{l}(a)-u_{l}(b)>u_{m}(c)-u_{m}(d)$. For this reason, and in light of the preceding discussion, in such rich special environments we must generally distinguish between the relative intensity difference that is captured by $s_{l}(a,b)>s_{m}(c,d)$ and the absolute intensity difference that is captured by the above interpersonal utility-difference inequality, which is preserved under any positive affine transformation of agents’ utilities that involves a common multiplicative factor (i.e. the Cardinal Unit Comparability informational basis—see also Section 4.2).
4 Intensity Dominance and Intensity-Efficient Allocations
4.1 Definitions
Since the number of agents and items are assumed to coincide, an allocation of the $n$ goods is a permutation on $A$. The set of all allocations is denoted by $\mathcal{A}$. As we discussed in the Introduction, we are interested in the assignment of the $n$ objects in $A$ to the $n$ agents in $V$ in a way that satisfies Pareto-efficiency but also improves upon it in delivering a narrower set of acceptable allocations that satisfy some additional desideratum that is appealing from a distributive-justice perspective and is operational in the present analytical environment where information on ordinal preference intensities and interpersonal comparisons thereof is assumed to be available. To this end, we proceed with the introduction of the following novel notions of dominance and efficiency.
Definition 4.
Let $S=\big{(}\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n}\bigr{)}\in\widehat{\mathcal{S}}$ be a strict intensity profile and $s$ its canonical representation. Given allocations $x$ and $y$, the former is said to intensity-dominate the latter if, for every pair of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$ it holds that $s_{i}(x_{i},x_{j})\geq s_{j}(y_{j},y_{i})$, and there is at least one pair $(i,j)$ where this inequality is strict.
Definition 5.
An allocation is intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ if it is Pareto-efficient at the induced preference profile $P_{S}\in\mathcal{P}^{V}$ and is not intensity-dominated by another Pareto-efficient allocation at $P_{S}$. It is intensity-dominant if it intensity-dominates every other Pareto-efficient allocation at $P_{S}$.
Definition 6.
Two allocations $x$ and $y$ that are intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ are equivalent if $s_{i}(x_{i},x_{j})=s_{j}(y_{j},y_{i})$ for all pairs of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$, and $x_{k}=y_{k}$ for every other agent $k\leq n$.
If $x$ and $y$ are Pareto-efficient allocations and $x$ intensity-dominates $y$, then in every pair of agents that is “flipped” by $x$ and $y$ in the sense that both allocations assign the same two alternatives $a$ and $b$ to the two agents in that pair but do so in opposite ways, the agent receiving $a$ (which, under the postulated Pareto-efficiency, is the mutually preferred alternative) under $x$ prefers it to $b$ weakly more than the agent receiving it under $y$ and, in at least one case, strictly more. An allocation that is both Pareto efficient and not intensity-dominated is intensity-efficient. Two intensity-efficient allocations $x$ and $y$ are equivalent if the agents in every “flipping” pair $(i,j)$ who are assigned $a$ and $b$ by $x$ and $y$ but in opposite ways have the same preference intensity for $a$ over $b$ in the sense that the intensity difference at $(a,b)$ is ranked the same way in the two agents’ intensity orderings, while every agent who is not in such a pair is treated identically by $x$ and $y$.
4.2 Intensity-Efficient and Utilitarian Allocations Are Distinct
To enable a comparison between the concept introduced above and the benchmark Pareto-efficiency refinement afforded by classical utilitarianism, we assume throughout this subsection that, in addition to satisfying the strictness condition $(\mathbf{5_{\mathit{D}}})$, each $\dot{\succsim}_{l}$ is utility-difference representable as in (3) under some cardinal utility index $u_{l}$. This is the case, for example, when $u_{l}:A\rightarrow\mathbb{R}$ is the utility function associated with an expected-utility representation of preferences over lotteries over $A$ and the intensity order defined by the $u_{l}$-differences is strictly convex or strictly concave relative to the agent’s induced (strict) preferences. With this assumption in place, we finally invoke the Cardinal Unit Comparability (Roemer, 1996; Baccelli, 2023) informational basis to social welfare. More specifically, we assume that: (i) $u_{l}:A\rightarrow\mathbb{R}$ is unique up to a positive affine transformation and features utility differences that define an intensity relation $\dot{\succsim}_{l}$ that abides with $(\mathbf{5_{\mathit{D}}})$; (ii) interpersonal utility-difference comparisons are possible under such transformations of any utility profile $U=(u_{1},\ldots,u_{n})$ where each $u_{l}$ is of this kind, as long as the scale/multiplicative parameter of such transformations is the same for all $u_{l}$. As is well-known, this informational basis is sufficient [^1] The reason why it is not necessary is closely related to the complex uniqueness properties of general utility-difference representations in the sense of (3). See Bossert (1991) for more on this. for the classical utilitarian social welfare functional—recalled below—to be meaningful.
Fixing a utility profile $U$ with the above properties, we refer to it as strict if the ordering of every agent’s utility differences is strict (see Example 1). In this case, clearly, the ordinal intensity profile $S$ that is induced by $U$ is also strict, i.e. an element of $\mathcal{\widehat{S}}$. Finally, recall that $x_{cu}(U)\in\mathcal{A}$ is a classical utilitarian allocation under profile $U$ if it holds that
$x_{cu}(U)$ $\in$ $\arg\max_{y\in\mathcal{A}}\sum_{l=1}^{n}u_{l}(y_{l})$ (11)
Proposition 1.
The sets of intensity-efficient and classical utilitarian allocations associated with strict utility profiles are generally distinct.
Proof.
Consider the strict utility profile of Example 1. Observe that the Pareto-efficient allocations are $(a_{1},a_{2},a_{3})$ and $(a_{2},a_{1},a_{3})$, and that the former allocation is the utilitarian one. Yet this is intensity-dominated by the latter since $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$. ∎
This potential divergence between intensity efficiency and the hallmark cardinal-welfarist allocation criterion is important because it shows that the hereby proposed notion is indeed distinct even under the richest possible informational assumptions. Furthermore, while perhaps not obvious ex ante, as is clarified by Example 1 and the discussion at the end of in Section 3, the reason why such a divergence may occur is that the classical utilitarian criterion is guided by the utility differences across agents. By contrast, in this special environment where intensity orderings are defined by the relative ordering rather than the absolute magnitude of utility differences, the Interpersonal Comparisons of Ordinal Intensities postulate upon which the intensity efficiency criterion is built compares the relative ranking of those differences across agents’ intensity orderings.
Naturally, one may inquire about the conditions under which classical utilitarian allocations are intensity-efficient. To this end, and motivated by the preceding remarks, we introduce the following condition that a utility profile may satisfy, which is sufficient—though not necessary—for utilitarian allocations to be intensity-efficient.
Definition 7.
A utility profile $U=(u_{1},\ldots,u_{n})$ is balanced if, for any two agents $i,j$ and alternatives $a,b$,
$u_{i}(a)-u_{i}(b)>u_{j}(a)-u_{j}(b)$ $\Longleftrightarrow$ $s_{i}(a,b)>s_{j}(a,b),$ (12)
where $s=(s_{1},\ldots,s_{n})$ is the profile of canonical preference intensity functions that represent the profile $S=(\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n})\in\widehat{\mathcal{S}}$ that is defined by the utility differences in $U$.
In words, a cardinal utility profile is balanced if interpersonal differences in the levels of utility differences are perfectly aligned with interpersonal differences in their rank orders.
Proposition 2.
Classical utilitarian allocations of strict and balanced utility profiles are intensity-efficient.
Proof.
Under the postulated conditions, suppose to the contrary that $x(U)$ is a utilitarian but not intensity-efficient allocation. Then, given that agent $l$’s intensity relation $\dot{\succsim}_{l}$ is defined by $(a,b)\,\dot{\succsim}_{l}\,(c,d)\Leftrightarrow u_{l}(a)-u_{l}(b)\geq u_{l}(c)-u_{l}(d)$, from the definition of intensity-dominance and intensity-efficiency there exist at least one pair of agents $(i,j)$ and of objects $(a,b)$ such that $x_{i}(U)=a$, $x_{j}(U)=b$ and $s_{j}(a,b)>s_{i}(a,b)$. By balancedness, this implies
$u_{j}(a)-u_{j}(b)$ $>$ $u_{i}(a)-u_{i}(b)$ (13)
Consider allocation $x^{\prime}$ that is identical to $x(U)$ except that $x^{\prime}_{i}=b$ and $x^{\prime}_{j}=a$. From (13) we get $u_{j}(a)-u_{i}(a)>u_{j}(b)-u_{i}(b)$. This shows that the utilitarian benefit from the $i\leftrightarrow j$ swap in the transition from $x(U)$ to $x^{\prime}$, $u_{j}(a)-u_{i}(a)$, exceeds the corresponding loss, $u_{j}(b)-u_{i}(b)$. It follows that $\sum_{l=1}^{n}u_{l}(x^{\prime}_{l})>\sum_{l=1}^{n}u_{l}\big{(}x_{l}(U)\bigr{)}$, which contradicts the postulated utilitarian-efficiency of $x(U)$. ∎
It is well-known that utilitarian allocations are Pareto-efficient. Proposition 2 shows that, when the utility profile is strict and balanced, this benchmark criterion has the additional normative status conveyed by intensity-efficiency. Furthermore, Proposition 2 implies the following existence result that is perhaps worth stating explicitly:
Corollary 1.
An intensity-efficient allocation exists for every intensity profile that is induced by a strict and balanced utility profile.
4.3 Existence without Cardinality Assumptions
We now return to the general environment in which intensity profiles come from the class of relations that were specified in Section 4.1, without imposing any further restrictions such as cardinal or even utility-difference representability of the individuals’ intensities.
Theorem 1.
An intensity-efficient allocation exists for every profile $S\in\mathcal{\widehat{S}}$ when $n=3$ but not in general when $n\geq 4$.
Proof.
The argument proceeds by considering the possible ways in which an arbitrary profile $S\in\mathcal{\widehat{S}}$ when $n=3$ might generate a sequence of distinct Pareto-efficient allocations that are implicated in an intensity-dominance cycle. To this end, let $D$ be the intensity-dominance relation that is introduced in Definition 4. Suppose to the contrary that
$w^{1}Dw^{2}D\ldots Dw^{k}Dw^{1}$ (14)
for Pareto efficient allocations $w^{1},\ldots,w^{k}$ on $X:=\{a,b,c\}$.
Observation 1.
$n=3$ implies that for any two allocations $w^{i},w^{i+1}$ such that $w^{i}Dw^{i+1}$ it must be that $w^{i}_{l}=w^{i+1}_{l}$ for exactly one agent $l\in\{1,2,3\}$ and $(w^{i}_{j},w^{i}_{k})=(w^{i+1}_{k},w^{i+1}_{j})$ for $j,k\neq l$.
Observation 2.
$n=3$ implies $k\leq 6$.
Observation 3.
Pareto efficiency of $w^{i}=(a^{\prime},b^{\prime},c^{\prime})$ and Strictness together imply
$s_{2}(a^{\prime},b^{\prime})>0$ $\Longrightarrow$ $s_{1}(a^{\prime},b^{\prime})>0$ (15)
$s_{3}(b^{\prime},c^{\prime})>0$ $\Longrightarrow$ $s_{2}(b^{\prime},c^{\prime})>0$ (16)
$s_{3}(a^{\prime},c^{\prime})>0$ $\Longrightarrow$ $s_{1}(a^{\prime},c^{\prime})>0$ (17)
Observation 4.
Strictness and canonicality of $s_{i}$ imply $s_{i}(a^{\prime},b^{\prime})=s_{i}(c^{\prime},d^{\prime})>0\Leftrightarrow(a^{\prime},b^{\prime})=(c^{\prime},d^{\prime})$ and, jointly with $n=3$, also $s_{i}(a^{\prime},b^{\prime})>0\Leftrightarrow s_{i}(a^{\prime},b^{\prime})\in\{1,2,3\}$.
Notice that (14) is impossible for $k=2$ because $D$ is asymmetric by construction. Suppose $k=3$. Without loss of generality, write $w^{1}:=(a,b,c)$ and $w^{2}:=(b,a,c)$. Then, by Observation 1, either $w^{3}=(b,c,a)$ or $w^{3}=(c,a,b)$. Since, in both cases, $w^{1}$ and $w^{3}$ are $D$-incomparable by construction, the $w^{3}Dw^{1}$ postulate in (14) is contradicted.
Now suppose $k=4$. By (14) and the above implications, we may take $w^{1}$, $w^{2}$, $w^{3}$ to be as in the $k=3$ case, from which it then follows that allocation $w^{4}$ must be either $(c,b,a)$ or $(a,c,b)$. Notice that either possibility is compatible with $w^{3}=(b,c,a)$ and with $w^{3}=(c,a,b)$. We therefore have the following 4 cases to consider:
[[Note: in what follows we make repeated use—often without explicit reference—of the Pareto-efficiency implications (15)–(17), the skew-symmetry property $s_{l}(a,b)=-s_{l}(b,a)$ of canonical representations and, whenever exact values of the $s_{i}$ functions are asserted, of the lateral-consistency property (6) together with the assumption that every $s_{i}$ is canonical and strict (cf Observations 34).]]
Case 1. $w^{3}=(b,c,a)$, $w^{4}=(c,b,a)$. By the definition of $D$, and by the above assumptions:
$w^{1}Dw^{2}$ $\Longrightarrow$ $s_{1}(a,b)>s_{2}(a,b)$ (18)
$w^{2}Dw^{3}$ $\Longrightarrow$ $s_{2}(a,c)>s_{3}(a,c)$ (19)
$w^{3}Dw^{4}$ $\Longrightarrow$ $s_{1}(b,c)>s_{2}(b,c)$
$w^{4}Dw^{1}$ $\Longrightarrow$ $s_{3}(a,c)>s_{1}(a,c)$
Therefore,
$s_{2}(a,c)>s_{3}(a,c)>s_{1}(a,c)$ (20)
Since $w^{i}$ is Pareto efficient for $i\leq 4$, it follows from (15)–(17) that there are 4 subcases to consider:
Subcase 1-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{2}(b,c)>0$. By (6) and the fact that $s_{1},s_{2}$ are canonical, this implies $s_{1}(a,c)=3$, which contradicts (20).
Subcase 1-ii. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{2}(c,b)>s_{1}(c,b)>0$. Suppose $s_{1}(a,c)>0$ is also true. Then, by (6), (20) and the fact that $s_{1}$ is canonical, $s_{1}(a,c)=1$ and $s_{1}(a,b)=3$. If $s_{2}(a,c)>0$ is also true, then $s_{2}(a,b)=3$. This contradicts $s_{1}(a,b)>s_{2}(a,b)$. So, it must be that $s_{2}(c,a)>0$ instead. But in this case $s_{2}(c,a)>0$, $s_{2}(a,b)>0$ and (6) together imply $s_{2}(c,b)=3$. This contradicts $s_{2}(a,c)=3$ which is now implied by (20) and the fact that the profile $s$ is canonical. Thus, it must be that $s_{1}(c,a)>0$ instead. So now we have $s_{1}(c,a)>0$, $s_{1}(a,b)>0$, which implies $s_{1}(c,b)=3$. But since, by assumption, $s_{2}(c,b)>s_{1}(c,b)$ and $s_{2}$ is canonical, this is a contradiction.
Subcase 1-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{2}(b,c)>0$. Suppose first that $s_{2}(a,c)>0$ is also true. Then, $s_{2}(b,a)>0$ and $s_{2}(a,c)>0$ implies $s_{2}(b,c)=3$. If $s_{1}(a,c)>0$ is also true, then (20) and the fact that $s$ is canonical together imply $s_{2}(a,c)=s_{2}(b,c)$, which contradicts (6) and Strictness. So, it must be that $s_{1}(c,a)>0$. From $s_{1}(b,c)>0$ and $s_{1}(c,a)>0$ we now get $s_{1}(b,a)=3$. In view of $s$ being canonical, this contradicts $s_{2}(b,a)>s_{1}(b,a)$.
Subcase 1-iv. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{2}(c,b)>s_{1}(c,b)>0$. Because $s$ is canonical, this and (6) readily imply $s_{1}(c,a)=s_{2}(c,a)=3$. But since (20) is equivalent to $s_{1}(c,a)>s_{3}(c,a)>s_{2}(c,a)$, this is a contradiction.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$.
Case 2. $w^{3}=(c,a,b)$, $w^{4}=(c,b,a)$.
We again have $w^{1}Dw^{2}\Longrightarrow s_{1}(a,b)>s_{2}(a,b)$ and $w^{2}Dw^{3}\Longrightarrow s_{1}(b,c)>s_{3}(b,c)$ and, in addition,
$w^{3}Dw^{4}$ $\Longrightarrow$ $s_{2}(a,b)>s_{3}(a,b)$ (21)
$w^{4}Dw^{1}$ $\Longrightarrow$ $s_{3}(a,c)>s_{1}(a,c)$ (22)
Therefore,
$s_{1}(a,b)>s_{2}(a,b)>s_{3}(a,b)$ (23)
In view of (15)–(17), we can now consider the following 4 possible subcases:
Subcase 2-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. By (23) and (15)–(17), the former postulate implies $s_{3}(a,b)>0$. Since $s$ is canonical, this further implies $s_{3}(a,b)=1$, $s_{2}(a,b)=2$ and $s_{1}(a,b)=3$. This, canonicality of $s$ and $s_{1}(b,c)>s_{3}(b,c)>0$ together imply $s_{1}(b,c)=2$ and $s_{3}(b,c)=1=s_{3}(a,b)$, which contradicts Strictness.
[…]
Subcase 2-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The former postulate, together with (23) and canonicality, implies $s_{3}(b,a)=3$, $s_{2}(b,a)=2$, $s_{1}(b,a)=1$. By (21), either $s_{3}(a,c)>s_{1}(a,c)>0$ or $s_{1}(c,a)>s_{3}(c,a)>0$ also holds. Consider the first possibility. From $s_{1}(b,a)=1$, $s_{1}(a,c)>0$, (6) and canonicality we get $s_{1}(a,c)=2$. This and (22) implies $s_{3}(a,c)=3$. Since $s_{3}(b,a)=3$ is also true, this contradicts Strictness. Hence, it must be that $s_{1}(c,a)>s_{3}(c,a)>0$. But in this case $s_{1}(b,c)>0$, $s_{1}(c,a)>0$, (6) and canonicality imply $s_{1}(b,a)=3$, which contradicts (23).
Subcase 2-iv. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{3}(c,b)>s_{1}(c,b)>0$. As in 2-iii, we have $s_{3}(b,a)=3$, $s_{2}(b,a)=2$, $s_{1}(b,a)=1$. But $s_{3}(c,b)>0$, $s_{3}(b,a)>0$ and (6) imply $s_{3}(c,a)>s_{3}(b,a)=3$ which, by canonicality, is impossible.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$ too.
Case 3. $w^{3}=(b,c,a)$, $w^{4}=(a,c,b)$.
We now have
$w^{3}Dw^{4}$ $\Longrightarrow$ $s_{3}(a,b)>s_{1}(a,b)$ (24)
$w^{4}Dw^{1}$ $\Longrightarrow$ $s_{3}(b,c)>s_{2}(b,c)$ (25)
It follows that
$s_{3}(a,b)>s_{1}(a,b)>s_{2}(a,b)$ (26)
We consider the 4 subcases that are now possible:
Subcase 3-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The latter, together with (15)–(17), (25) and canonicality, implies $s_{3}(a,b)=1$. But canonicality, (15)–(17) and (17) also implies $s_{3}(a,b)=3$, which contradicts Strictness.
Subcase 3-ii. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{3}(c,b)>s_{1}(c,b)>0$. The first postulate and (26), together with canonicality, implies $s_{3}(a,b)=3$, $s_{1}(a,b)=2$ and $s_{2}(a,b)=1$. Since $s_{3}(c,b)>s_{1}(c,b)>0$ is also assumed, this and canonicality further imply $s_{3}(c,b)=2$. Now, because $s_{3}(a,b)>0$ and $s_{3}(c,b)>0$, it follows from (6) that $s_{3}(c,a)>0$ too. But (6) in this case further implies $s_{3}(c,a)>s_{3}(a,b)=3$, which is impossible.
Subcase 3-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The first postulate, together with (26) and canonicality, implies $s_{2}(b,a)=3$, $s_{1}(b,a)=2$, $s_{3}(b,a)=1$. The second postulate and $s_{1}(b,a)=2$, together with Strictness, implies $s_{1}(b,c)=3$. This in turn implies $s_{1}(a,c)=1$ or $s_{1}(c,a)=1$. If the latter is true, then $s_{1}(b,c)>0$, $s_{1}(c,a)>0$ and (6), together with canonicality, implies $s_{1}(b,a)=3$, a contradiction. Hence, it must be that $s_{1}(a,c)=1$. We therefore have $s_{1}(b,a)=2$, $s_{1}(a,c)=1$ and, by (6) and canonicality, $s_{1}(b,c)=3$. From (15)–(17), (19), (25) and canonicality we also know that $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)>0$ implies $s_{3}(b,c)=2$ and $s_{2}(b,c)=1$. Thus, we have $s_{3}(b,a)>0$, $s_{3}(b,c)>0$ and, by (15)–(17) and $s_{1}(a,c)>0$, also $s_{3}(a,c)>0$. But $s_{3}(b,a)>0$, $s_{3}(a,c)>0$ together with (6) and canonicality implies $s_{3}(b,c)=3$, a contradiction.
Subcase 3-iv. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{3}(c,b)>s_{1}(c,b)>0$. These readily imply $s_{1}(c,a)=3$. As above, (26) implies $s_{2}(b,a)=3$, $s_{1}(b,a)=2$ and $s_{3}(b,a)=1$. By (6), $s_{1}(c,a)=3$ and $s_{1}(a,b)=2$ implies $s_{1}(c,b)=1$. From the above postulates and from (25), $s_{2}(c,b)>s_{3}(c,b)>s_{1}(c,b)$ further implies $s_{2}(c,b)=3$, which contradicts $s_{2}(b,a)=3$ and Strictness.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$ here as well.
Case 4. $w^{3}=(c,a,b)$, $w^{4}=(a,c,b)$. It is now true that
$w^{3}Dw^{4}$ $\Longrightarrow s_{2}(a,c)>s_{1}(a,c)$ (27)
$w^{4}Dw^{1}$ $\Longrightarrow s_{3}(b,c)>s_{2}(b,c)$ (28)
which, together with (18)–(19), imply
$s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)$ (29)
Suppose first that $s_{2}(b,c)>0$. Then, by (29) and (15)–(17), $s_{1}(b,c)=3$, $s_{2}(b,c)=1$ and $s_{3}(b,c)=2$. From $s_{1}(b,c)=3$ and (6) we also get $s_{1}(b,a)>0$ and $s_{1}(a,c)>0$. Hence, by (18), $s_{2}(b,a)>s_{1}(b,a)>0$ and, by (27), $s_{2}(a,c)>s_{1}(a,c)>0$. These inequalities and (6) together imply $s_{2}(b,c)=3$, which is a contradiction.
Now suppose instead that $s_{2}(b,c)<0$, i.e. $s_{2}(c,b)>0$. It follows from (29) that $s_{2}(c,b)=3$, $s_{3}(c,b)=2$ and $s_{1}(c,b)=1$. Suppose $s_{1}(a,c)>0$. From (27) and (15)–(17), $s_{2}(a,c)>0$. Since $s_{2}(a,c)>0$ and $s_{3}(c,b)>0$, by (6) we get $s_{2}(a,b)>s_{2}(c,b)=3$, which is impossible. Hence, $s_{1}(c,a)>0$ holds instead and, from (27) and (15)–(17), $s_{1}(c,a)>s_{2}(c,a)>0$ is also true. Suppose $s_{2}(a,b)>0$ holds too. By (18), $s_{1}(a,b)>0$. By (6) and $s_{1}(c,a)>0$, $s_{1}(a,b)>0$ we get $s_{1}(c,b)=3$, a contradiction. Hence, $s_{2}(b,a)>0$ must be true instead and, by (18), $s_{2}(b,a)>s_{1}(b,a)>0$ also. So, we have $s_{2}(c,b)>0$ and $s_{2}(b,a)>0$, which, by (6), implies $s_{2}(c,a)>s_{(}c,b)=3$. This too is a contradiction.
Hence, $w^{4}Dw^{1}$ is impossible for such $w^{3}$ and $w^{4}$ also.
Next, suppose $k=5$. Arguing as above, allocations $w^{1},\ldots,w^{4}$ in (14) must be as in one of the four cases considered previously. Combined with the fact that each $w^{i}$ in sequence $(w^{1},\ldots,w^{5})$ must be distinct and the notational convention $w^{1}=(a,b,c)$ and $w^{2}=(b,a,c)$, this gives rise to the following four possibilities:
[…]
Clearly, because either $w^{5}=(b,c,a)$ or $w^{5}=(c,a,b)$ must hold in all four cases, and recalling that $w^{1}=(a,b,c)$, it cannot be that $w^{5}Dw^{1}$.
Finally, suppose $k=6$. With allocations $w^{1},\ldots,w^{5}$ in (14) being as in the $k=5$ case that was just considered above, $w^{6}$ can only coincide with allocation $(a,c,b)$ in each of the four relevant cases. In view of the previous steps, these are as follows:
Case 1: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(b,c,a)$, $w^{4}=(c,b,a)$, $w^{5}=(c,a,b)$, $w^{6}=(a,c,b)$. By definition of $D$, and by the above assumptions: $w^{1}Dw^{2}\Longrightarrow s_{1}(a,b)>s_{2}(a,b)$; $w^{2}Dw^{3}\Longrightarrow s_{2}(a,c)>s_{3}(a,c)$; $w^{3}Dw^{4}\Longrightarrow s_{1}(b,c)>s_{2}(b,c)$; $w^{4}Dw^{5}\Longrightarrow s_{3}(a,b)>s_{2}(a,b)$; $w^{5}Dw^{6}\Longrightarrow s_{2}(a,c)>s_{1}(a,c)$; and $w^{6}Dw^{1}\Longrightarrow s_{3}(b,c)>s_{2}(b,c)$. It follows that
$s_{2}(a,c)>s_{1}(a,c)>s_{1}(a,b)>s_{2}(a,b)$ (30)
Suppose $s_{2}(a,b)>0$. Then, (30) implies $s_{2}(a,c)>3$, which contradicts canonicality. If $s_{2}(b,a)>0$ instead, then (30) together with skew-symmetry of $s_{2}$ implies $s_{2}(b,a)>3$ and results in the same contradiction.
Case 2: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(c,a,b)$, $w^{4}=(c,b,a)$, $w^{5}=(b,c,a)$, $w^{6}=(a,c,b)$. Notice that the postulated dominance implications $w^{1}Dw^{2}\Longrightarrow s_{1}(a,b)>s_{2}(a,b)$, $w^{3}Dw^{4}\Longrightarrow s_{2}(a,b)>s_{3}(a,b)$ and $w^{5}Dw^{6}\Longrightarrow s_{3}(a,b)>s_{1}(a,b)$ lead to $s_{1}(a,b)>s_{2}(a,b)>s_{3}(a,b)>s_{1}(a,b)$, which is absurd.
Case 3: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(b,c,a)$, $w^{4}=(a,c,b)$, $w^{5}=(c,a,b)$, $w^{6}=(a,c,b)$. Observe here that the postulated dominance implications $w^{4}Dw^{5}\Longrightarrow s_{1}(a,c)>s_{2}(a,c)$ and $w^{5}Dw^{6}\Longrightarrow s_{2}(a,c)>s_{1}(a,c)$ directly contradict each other.
Case 4: $w^{1}=(a,b,c)$, $w^{2}=(b,a,c)$, $w^{3}=(c,a,b)$, $w^{4}=(a,c,b)$, $w^{5}=(b,c,a)$, $w^{6}=(a,c,b)$. As in Case 2, the postulated dominance implications $w^{4}Dw^{5}\Longrightarrow s_{1}(a,b)>s_{3}(a,b)$ and $w^{5}Dw^{6}\Longrightarrow s_{3}(a,b)>s_{1}(a,b)$ result in the same contradiction.
It has therefore been shown that $D$ is acyclic when $n=3$, establishing the existence of an intensity-efficient allocation in this case.
For the claimed potential non-existence when $n\geq 4$, consider the example strict intensity profile on $A=\{a,b,c,d\}$ whose canonical representation is shown in the table below:
$s_{i}(a,b)=$ $i=$ 1 2 3 4
6 $(a,d)$ $(a,d)$ $(a,d)$ $(a,d)$
5 $(b,d)$ $(a,c)$ $(a,c)$ $(b,d)$
4 $(a,c)$ $(b,d)$ $(b,d)$ $(a,c)$
3 $(b,c)$ $(b,c)$ $(b,c)$ $(b,c)$
2 $(a,b)$ $(c,d)$ $(c,d)$ $(a,b)$
1 $(c,d)$ $(a,b)$ $(a,b)$ $(c,d)$
We start by observing that $s_{1}=s_{4}$, $s_{2}=s_{3}$ and $a\succ_{i}b\succ_{i}c\succ_{i}d$, $i=1,\ldots,4$. This implies that all 24 possible allocations, defined and listed below, are Pareto efficient.
$\begin{array}[]{llllllllllllllllll}x^{1}&=&(a,b,c,d)&&&x^{2}&=&(a,b,d,c)&&&x^{3}&=&(a,c,b,d)&&&x^{4}&=&(a,c,d,b)\\ x^{5}&=&(a,d,b,c)&&&x^{6}&=&(a,d,c,b)&&&x^{7}&=&(b,a,c,d)&&&x^{8}&=&(b,a,d,c)\\ x^{9}&=&(b,c,a,d)&&&x^{10}&=&(b,c,d,a)&&&x^{11}&=&(b,d,a,c)&&&x^{12}&=&(b,d,c,a)\\ x^{13}&=&(c,a,b,d)&&&x^{14}&=&(c,a,d,b)&&&x^{15}&=&(c,b,a,d)&&&x^{16}&=&(c,b,d,a)\\ x^{17}&=&(c,d,a,b)&&&x^{18}&=&(c,d,b,a)&&&x^{19}&=&(d,a,b,c)&&&x^{20}&=&(d,a,c,b)\\ x^{21}&=&(d,b,a,c)&&&x^{22}&=&(d,b,c,a)&&&x^{23}&=&(d,c,a,b)&&&x^{24}&=&(d,c,b,a)\end{array}$
The following comparisons, whose validity can be readily established by the reader, demonstrate that for each $x^{i}$, $i=1,\ldots,24$, in this set there is a distinct $x^{j}$ that intensity-dominates $x^{i}$—a situation denoted by $x^{j}Dx^{i}$.
$\begin{array}[]{llllllll}x^{6}Dx^{1}&x^{1}Dx^{2}&x^{4}Dx^{3}&x^{13}Dx^{4}&x^{3}Dx^{5}&x^{17}Dx^{6}&x^{1}Dx^{7}&x^{2}Dx^{8}\\ x^{3}Dx^{9}&x^{8}Dx^{10}&x^{9}Dx^{11}&x^{11}Dx^{12}&x^{18}Dx^{13}&x^{13}Dx^{14}&x^{17}Dx^{15}&x^{22}Dx^{16}\\ x^{23}Dx^{17}&x^{24}Dx^{18}&x^{8}Dx^{19}&x^{22}Dx^{20}&x^{11}Dx^{23}&x^{12}Dx^{22}&x^{24}Dx^{23}&x^{10}Dx^{24}\end{array}$
Therefore, no intensity-efficient allocation exists in this profile. ∎
The somewhat surprising fact that a plausible dominance concept may by cyclic and prevent an optimal entity to emerge invites an informal analogy between intensity-dominance cycles over allocations with at least 4 agents and Condorcet cycles over alternatives in pairwise majority-based preference aggregation with at least 3 agents (Condorcet, 1785). Importantly, however, although intensity-dominance cycles here may prevent refining the Pareto set, unlike the Condorcet social-welfare theoretic framework these do not lead to a “policy paralysis” problem because the Pareto set is always non-empty and, absent any distributively juster suggestions, one of them might be promoted by the social planner. That said, it is naturally of interest to understand better the conditions under which our proposed refinement of Pareto-efficiency is well-defined. This is the focus of the next subsection.
4.4 Existence Condition for General $n$: Homeo-Monotonicity
For any set $A^{\prime}\subseteq A$ and preference relation $\succ_{i}$ that is induced by an intensity relation $\dot{\succsim}_{i}\in\mathcal{I}$ we write $\succ_{i}^{A^{\prime}}$ for the restriction of $\succ_{i}$ on $A^{\prime}$.
Definition 8.
Given a profile $S\in\mathcal{\widehat{S}}$, a set of agents $V^{\prime}\subseteq V$ is said to be:
  1. 1.
    Envy-free with respect to a set of items $A^{\prime}\subseteq A$ if each $v_{i}\in V^{\prime}$ has a distinct $\succ_{i}$-best item in $A$, and this item (denoted $b_{i}^{\prime}$) belongs to $A^{\prime}$.
  2. 2.
    Top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ if $\succ_{i}^{A^{\prime}}=\succ_{j}^{A^{\prime}}$ for all $v_{i},v_{j}\in V^{\prime}$ and, when $A^{\prime}\subset A$, $a^{\prime}\in A^{\prime}$, $a\in A\setminus A^{\prime}$ holds, then $a^{\prime}\succ_{i}a$ for all $v_{i}\in V^{\prime}$.
An envy-free set $V^{\prime}$ with respect to $A^{\prime}$ consists of agents in $V^{\prime}$ who have distinct overall-best items in $A^{\prime}$. This terminology is motivated by the fact that, if $V^{\prime}\equiv V$ and $A^{\prime}\equiv A$, then each agent receiving their most preferred option results in an allocation that is Pareto-efficient and also envy-free (Varian, 1974; Moulin, 2004). A top-preference-coincident pair consists of agents whose preferences over $A^{\prime}$ are identical and also such that every item in $A^{\prime}$ is preferred to everything in $A\setminus A^{\prime}$, with no restrictions on preferences in the latter set.
Definition 9.
A set of agents $V^{\prime}\subseteq V$ that is top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ at profile $S\in\mathcal{\widehat{S}}$ is monotonic with respect to $A^{\prime}$ if $|V^{\prime}|=|A^{\prime}|$ and
$s_{1}(a_{1},a_{2})$ >> $s_{1}(a_{2},a_{3})$ >> $\cdots$ >> $s_{1}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
$\geq$ $\geq$ $\cdots$ $\geq$
$s_{2}(a_{1},a_{2})$ >> $s_{2}(a_{2},a_{3})$ >> $\cdots$ >> $s_{2}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
$\geq$ $\geq$ $\cdots$ $\geq$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
$\geq$ $\geq$ $\cdots$ $\geq$
$s_{|V^{\prime}|}(a_{1},a_{2})$ >> $s_{|V^{\prime}|}(a_{2},a_{3})$ >> $\cdots$ >> $s_{|V^{\prime}|}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
It is strictly monotonic if every weak inequality above is replaced by a strict inequality.
Finally, a (strictly) monotonic such set is homeo-monotonic with respect to $A^{\prime}$ if
$s_{i}(a_{i},a_{j})$ $\geq$ $s_{j}(a_{i},a_{j})$ for all $1\leq i<j\leq|V^{\prime}|$.
Verbally, when $k$ agents have identical preferences over their top $k$ items, and agents can be (strictly) ordered in how much they prefer any two consecutively-ranked items in this subset [cf (strict) monotonicity], then this ordering cannot be violated in how much they prefer any non-consecutively ranked items in this subset (cf homeo-monotonicity).
Definition 10.
A profile $S\in\mathcal{\widehat{S}}$ is homeo-monotonic if, whenever $V$ and $A$ can be partitioned into subsets $V^{1},\ldots,V^{m}$ and $A^{1},\ldots,A^{m}$ where $V^{1}$ is envy-free with respect to $A^{1}$ and $V^{i>1}$ is monotonically top-preference-coincident with respect to $A^{i>1}$, then $V^{i>1}$ is homeo-monotonic with respect to $A^{i>1}$.
Definition 11.
A profile $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic if it induces partitions of agents and items relative to which it is homeo-monotonic.
Example 2.
The table below describes an effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where $n=6$ and where the envy-free and top-preference-coincident sets that partition $V$ and $A$ are $(V^{1},A^{1})=(\{v_{1}\},\{a_{1}\})$, $(V^{2},A^{2})=(\{v_{2},v_{3},v_{4}\},\{a_{2},a_{3},a_{4}\})$, $(V^{3},A^{3})=(\{v_{5},v_{6}\},\{a_{5},a_{6}\})$:
$\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$ $\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{3}$}\hskip 20.0pt}$
1 2 3 4 5 6
$s_{i}(a,b)=$ $i=$ ${\color[rgb]{1,0,1}a_{1}}\succ_{1}\cdots$ ${\color[rgb]{1,0,0}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}}\succ_{i}\cdots$ ${\color[rgb]{0,0,1}a_{5}\succ_{i}a_{6}}\succ_{i}\cdots$
15 $(a_{1},a_{6})$ $(a_{2},a_{1})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{5},a_{1})$ $(a_{5},a_{3})$
14 $(a_{2},a_{6})$ $(a_{2},a_{6})$ $(a_{3},a_{5})$ $(a_{3},a_{6})$ $(a_{5},a_{4})$ $(a_{5},a_{1})$
13 $(a_{1},a_{5})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{3})$ $(a_{5},a_{4})$
12 $(a_{3},a_{6})$ $(a_{2},a_{4})$ $(a_{3},a_{6})$ $(a_{2},a_{1})$ $(a_{5},a_{2})$ $(a_{5},a_{2})$
11 $(a_{2},a_{5})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$ $(a_{3},a_{1})$ $(a_{6},a_{1})$ $(a_{6},a_{3})$
10 $(a_{1},a_{4})$ $(a_{3},a_{6})$ $(a_{3},a_{1})$ $(a_{4},a_{1})$ $(a_{6},a_{4})$ $(a_{6},a_{1})$
9 $(a_{4},a_{6})$ $(a_{3},a_{5})$ $(a_{4},a_{5})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$
8 $(a_{3},a_{5})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{6},a_{3})$ $(a_{6},a_{4})$
7 $(a_{2},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$
6 $(a_{1},a_{3})$ $(a_{4},a_{1})$ $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{2},a_{4})$ $(a_{4},a_{3})$
5 $(a_{5},a_{6})$ $(a_{4},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{1})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$
4 $(a_{4},a_{5})$ $(a_{5},a_{1})$ $(a_{1},a_{5})$ $(a_{1},a_{6})$ $(a_{6},a_{2})$ $(a_{6},a_{2})$
3 $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{4},a_{1})$ $(a_{2},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$
2 $(a_{2},a_{3})$ $(a_{5},a_{6})$ $(a_{1},a_{6})$ $(a_{2},a_{3})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$
1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
$\blacklozenge$
Remark 1.
If $V$ is envy-free with respect to $A$, then $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
Clearly, any set of agents can be partitioned into groups according to whether distinct or identical alternatives are the most preferred within each group. The additional structure imposed by effectively homeo-monotonic profiles amounts to requiring that: (i) the preferences of agents with the same top-ranked item also coincide in how they rank their second-, third-, $\ldots$, $k$th-most preferred item for some $k\leq n$ that coincides with the number of agents in that group; (ii) the preferences of distinct such groups are sufficiently heterogeneous. While this condition is restrictive and one can easily construct profiles where intensity-efficient allocations exist but the condition is violated, the proof of the next result shows that it is helpful toward clarifying some important mechanisms for the existence of such allocations in the general case. Moreover, a weaker version of this condition that we consider later turns out to be necessary for such existence.
Theorem 2.
Every effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation.
Proof.
Consider a profile $S\in\mathcal{\widehat{S}}$ with the postulated structure. Let $V^{1}:=\{v^{1}_{1},\ldots,v^{1}_{|V^{1}|}\}$ be the envy-free set with respect to $A^{1}:=\{a^{1}_{1},\ldots,a^{1}_{|A^{1}|}\}$, and let $V^{i>1}:=\{i_{1},\ldots,i_{|V^{i}|}\}$ be the $i$th top-preference-coincident set with respect to $A^{i>1}:=\{i_{1},\ldots,i_{|A^{i}|}\}$, $i=2,\ldots,m$. By the profile assumption, $|A^{i}|=|V^{i}|$ for all $i\leq m$, and the collections $(V^{i})_{i=1}^{m}$ and $(A^{i})_{i=1}^{m}$ partition the sets $V$ and $A$, respectively. Without loss of generality we may assume that $a^{1}_{j}$ is the top-ranked item of agent $v^{1}_{j}$, $j=1,\ldots,|A^{1}|$ (for the general case we apply a permutation over $\{1,\ldots,|A^{1}|\}$ on the $a^{1}$-subscripts). Recalling our earlier notation, this amounts to writing $a_{1}^{1}\equiv b_{1}^{1}$, $a_{2}^{1}\equiv b_{2}^{1}$, etc. Also without loss, for $A^{i>1}$, $V^{i>1}$ we may assume that $a^{i}_{1}\succ_{j}a^{i}_{2}\succ_{j}\cdots\succ_{j}a^{i}_{|A^{i}|}$ for all $v_{j}\in V^{i>1}$. (Note: the special cases where $V^{1}=\emptyset$, $V^{1}=V$ or $V^{i}=V$ for some $i>1$ are allowed by the postulated $S\in\mathcal{\widehat{S}}$, and so will they be in the proof.)
Lemma 3.
If $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic, then an allocation $x$ is Pareto-efficient with respect to $P_{S}$ if and only if it satisfies the following:
  1. (i)
    $v_{j}\in V^{1}$ $\Rightarrow$ $x_{j}=b_{j}^{1}$.
  2. (ii)
    $v_{j}\in V^{i>1}$ $\Rightarrow$ $x_{j}\in A^{i>1}$.
Proof.
For the “if” part, let $x$ be an arbitrary allocation that satisfies (i) and (ii). Assume to the contrary that $x$ is not Pareto efficient. Let $y$ be another allocation that Pareto-dominates $x$. This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$. Let $V^{*}\subsetneq V$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$. By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$. Suppose $V^{*}=\emptyset$. From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$. If the envy-free set $V^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in V^{1}$. If $V^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $V^{i>1}$ is non-empty. Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in V^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction. Now suppose $V^{*}\neq\emptyset$. Since $y$ Pareto-dominates $x$, the set $V^{**}:=V\setminus V^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty. From the preceding argument we know that for every $v_{l}\in V^{**}$ it must be that $v_{l}\not\in V^{1}$ and $v_{l}\neq v_{1}^{i}\in V^{i>1}$, for all $i>1$. By $s_{l}(y_{l},x_{l})>0$ and $|A|=|V|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$. By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$. Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in V^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $V^{i>1}$ and $V^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$. By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$. Since, by the definition of top-preference-coincident sets, $s_{l}(a,a^{\prime})>0$ and $s_{k}(b,b^{\prime})>0$ holds for all $a\in A^{i>1},a^{\prime}\in A\setminus A^{i>1}$ and $b\in A^{j>1},b^{\prime}\in A\setminus A^{j>1}$, and since, by (ii), it holds that $x_{l}\in A^{i>1}$ and $x_{k}\equiv y_{l}\in A^{j>1}$, the postulate $s_{l}(y_{l},x_{l})>0$ leads to a contradiction. Therefore, $x$ is Pareto-efficient.
For the “only if” part, consider an arbitrary Pareto-efficient allocation $x$. Suppose to the contrary that (i) is false. That is, $x_{j}\neq b_{j}^{1}$ for some $v_{j}\in V^{1}$. Let $b_{j}^{1}=x_{l}$ for the relevant distinct agent $v_{l}\in V$ (such an agent exists because $|V|=|A|$). Either $v_{l}\in V^{1}$ or $v_{l}\in V^{i>1}$. In the former case, swapping $x_{j}$ and $x_{l}$ between the two agents Pareto-improves upon $x$, contradicting its postulated Pareto optimality. In the latter case, from the assumption that $V^{i>1}$ is top-preference-coincident with respect to $A^{i>1}$, and that $S$ is homeo-monotonic, we have $a^{i}_{t}\succ_{l}x_{l}\equiv b^{1}_{j}$ for every $a^{i}_{t}\in A^{i>1}$. Furthermore, in this case (again because $|V|=|A|$) there must also exist another agent $k\in V^{p}$, $p\neq i$, such that $x_{k}\in A^{i>1}$. There are two possibilities: $p=1$ or $p>1$. If $p=1$, then the above contradiction is obtained now for $x_{l}$ and $x_{k}$. If $p>1$, then, by homeo-monotonicity of $S$ and the implied structure of $(V^{l>1},A^{l>1})$, $(V^{p>1},A^{p>1})$, we have $s_{l}(x_{k},x_{l})>0$ and $s_{k}(x_{l},x_{k})>0$, which also contradicts the postulated Pareto-efficiency of $x$. The argument establishing necessity of (ii) is analogous and omitted. ∎
Next, consider a top-preference-coincident set $V^{i>1}$ with respect to $A^{i>1}$. By effective homeo-monotonicity, either such a non-empty set exists or the envy-free set $V^{1}$ coincides with $V$. In the latter case, the allocation $x^{*}\in\mathcal{A}$ defined by $x^{*}_{i}=b^{1}_{i}$ is obviously the unique Pareto-efficient allocation at $S$ and therefore, trivially, intensity-efficient. So, in the sequel we assume that $V^{i>1}\neq\emptyset$.
Denote by $a_{1}^{i}\succ_{i}a_{2}^{i}\succ_{i}\cdots\succ_{i}a_{|A^{i}|}^{i}$ the induced preference order on $A^{i>1}$ which, by the definition of $(V^{i>1},A^{i>1})$, is common across all agents in $V^{i>1}$. Define the binary relation $\trianglerighteq^{i}$ on $V^{i>1}$ by
$v_{j}\trianglerighteq^{i}v_{k}$ $\Longleftrightarrow$ $s_{j}(a^{i}_{l},a^{i}_{l+1})\geq s_{k}(a^{i}_{l},a^{i}_{l+1})\text{ for all }l=1,\ldots,|A^{i>1}|-1$
By the postulated monotonicity of $(V^{i>1},A^{i>1})$, $\trianglerighteq^{i}$ is a weak order on $V^{i>1}$. Let $\trianglerighteq_{i}^{*}$ be some extension of $\trianglerighteq^{i}$ into a linear order over $V^{i>1}$, and denote by $\trianglerighteq^{*i}(j)$ the $\trianglerighteq^{*i}$-rank of agent $v_{j}\in V^{i>1}$. By construction, $\trianglerighteq^{*i}$ coincides with $\trianglerighteq^{i}$ except where it breaks any equivalence ties that the latter relation may feature.
Now define $x^{*}\in\mathcal{A}$ by
$x_{j}^{*1}:=b_{j}^{1},\;j=1,\ldots,|V^{1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{1}$
$x_{j}^{*i>1}:=a^{i}_{l},\;j=1,\ldots,|V^{i>1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{i>1}\text{ and }\trianglerighteq^{*i}(j)=l$
That is, each agent in the set $V^{1}$ (which could be empty) is assigned their most preferred item, while the $\trianglerighteq^{*i}$-$l$th-ranked agent in $V^{i>1}$ is assigned their $l$th most preferred item. Recalling that $V^{1}$ is envy-free and each $V^{i>1}$ is top-preference-coincident, it follows from Lemma 3 that $x^{*}$ is Pareto-efficient. We will prove that it is also intensity-efficient.
To this end, suppose to the contrary that some Pareto-efficient allocation $y\in\mathcal{A}$ intensity-dominates $x^{*}$. Then, there are agents $v_{i},v_{j}\in V$ such that $(x_{i}^{*},x_{j}^{*})=(y_{j},y_{i})$ and $s_{j}(y_{j},y_{i})\equiv s_{j}(x_{i}^{*},x_{j}^{*})>s_{i}(x^{*}_{i},x_{j}^{*})>0$. By construction of $x^{*}$, $v_{i},v_{j}$ cannot belong to the same $V^{l\geq 1}$. For $l=1$ this is obvious because $x^{*i}=b^{1}_{i}$ and $x^{*j}=b^{1}_{j}$, contradicting the above postulate. For $l>1$, moreover, one has $x^{*}_{i},x_{j}^{*}\in A^{l>1}$ while items in $A^{l>1}$ are assigned to agents in $V^{l>1}$ according to the linear order $\trianglerighteq^{*l}$, which, by homeo-monotonicity of $S$, ensures that $s_{i}(x_{i}^{*},x_{j}^{*})\geq s_{j}(x_{i}^{*},x_{j}^{*})$, thereby contradicting that postulate. The remaining possibility is $v_{i}\in V^{l>1}$ and $v_{j}\in V^{k>1}$ for some $k\neq l$. By construction of $x^{*}$, however, we have $x_{i}^{*}\in A^{l>1}$ and $x_{j}^{*}\in A^{k>1}$ which, because $V^{l>1}$, $V^{k>1}$ are top-preference-coincident with respect to $A^{l>1}$, $A^{k>1}$ and $A^{l>1}\cap A^{k>1}=\emptyset$, implies $s_{i}(x_{i}^{*},x_{j}^{*})\equiv s_{i}(y_{j},y_{i})>0$ and $s_{j}(x_{i}^{*},x_{j}^{*})\equiv s_{j}(y_{j},y_{i})<0$. Thus, the postulate that $y$ intensity-dominates $x^{*}$ is contradicted in all situations. ∎
Example 2 (continued).
By Lemma 3, there are 12 Pareto-efficient allocations in this profile, namely those which assign $a_{1}$ to $v_{1}$ and alternate items $a_{2},a_{3},a_{4}$ among agents $v_{2},v_{3},v_{4}$ and $a_{5},a_{6}$ among $v_{5},v_{6}$. Out of these 12 allocations, only the following 2 are the (in fact, equivalent) intensity-efficient ones: $(a_{1},a_{2},a_{3},a_{4},a_{5},a_{6})$; $(a_{1},a_{2},a_{3},a_{4},a_{6},a_{5})$. $\blacklozenge$
The way in which the intensity-efficient allocation $x^{*}$ is constructed in the proof of Theorem 2 readily suggests the following stronger implication for a class of special cases:
Corollary 2.
An intensity-dominant—up to equivalence—allocation exists in any effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where every top-preference-coincident set has two agents.
Importantly, although effective homeo-monotonicity is a sufficient but not necessary condition for the existence of intensity-efficient allocations, we show next that the restriction it imposes within monotonic top-preference-coincident sets—if the profile induces such sets—is also necessary whenever these sets are strictly monotonic.
Theorem 3.
If a profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation, then every strictly monotonic top-preference-coincident set in it is homeo-monotonic.
Proof.
[…]
Lemma 4.
[…]
Proof.
The only difference between allocations $x$ and $x^{\prime}$ lies in how they permute items in $A^{\prime}$ to agents in $V^{\prime}$. Denote their restrictions on these subsets by $\widehat{x}$ and $\widehat{x}^{\prime}$, respectively. By Lemma 3 and the postulated intensity-efficiency (hence Pareto-efficiency) of $x$ it follows that $x^{\prime}$ is Pareto-efficient too. Since $x$ is intensity-efficient, it is not intensity-dominated by $x^{\prime}$. Hence, there is no pair $(a^{\prime},a^{\prime\prime})\in A^{\prime}\times A^{\prime}$ such that $(a^{\prime},a^{\prime\prime})=(\widehat{x}_{i},\widehat{x}_{j})=(\widehat{x}^{\prime}_{j},\widehat{x}^{\prime}_{i})$ for some $v^{\prime}_{i},v^{\prime}_{j}\in V^{\prime}$, making $x$ and $x^{\prime}$ incomparable by intensity-dominance. The argument in the proof of Theorem 2 following Lemma 3 establishes that there is no other Pareto-efficient allocation that intensity-dominates $x^{\prime}$ either. Thus, $x^{\prime}$ is intensity-efficient. ∎
[…]
$\begin{array}[]{lllllllll}s^{\prime}_{1}(a^{\prime}_{i},a^{\prime}_{i+1})&>&s^{\prime}_{2}(a^{\prime}_{i},a^{\prime}_{i+1})&>&\cdots&>&s^{\prime}_{|V^{\prime}|}(a^{\prime}_{i},a^{\prime}_{i+1})&&\text{for all }i\leq|A^{\prime}|-1\end{array}$
Now recall that the allocation $x^{\prime}$ constructed above is intensity-efficient. This assigns $x^{\prime}_{m}=a^{\prime}_{m}\in A^{\prime}$ to each $v^{\prime}_{m}\in V^{\prime}$. Since $(V^{\prime},A^{\prime})$ is strictly monotonic but not homeo-monotonic, there are $v^{\prime}_{j},v^{\prime}_{l}\in V^{\prime}$ such that $v^{\prime}_{j}\trianglerighteq^{\prime}v^{\prime}_{l}$, jl>1j-l>1 and $s^{\prime}_{l}(a^{\prime}_{j},a^{\prime}_{l})>s^{\prime}_{j}(a^{\prime}_{j},a^{\prime}_{l})$. Since $(x^{\prime}_{j},x^{\prime}_{l})=(a^{\prime}_{j},a^{\prime}_{l})$, this implies that the allocation $x^{\prime\prime}$ which is identical to $x^{\prime}$ except that $(x^{\prime\prime}_{j},x^{\prime\prime}_{l})=(a^{\prime}_{l},a^{\prime}_{j})$ intensity-dominates $x^{\prime}$, contradicting the postulated intensity-efficiency of the latter allocation. ∎
The next example explains why “strictly monotonic” cannot be replaced by “monotonic” in the statement of Theorem 3.
Example 3.
For $n=4$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as follows:
$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$
1 2 3 4
$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}$ $a_{4}\succ_{4}\cdots\succ_{4}a_{1}$
6 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{4},a_{1})$
5 $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{1},a_{3})$
4 $(a_{1},a_{3})$ $(a_{1},a_{3})$ $(a_{2},a_{4})$
3 $(a_{1},a_{2})$ $(a_{1},a_{2})$ $(a_{3},a_{4})$
2 $(a_{2},a_{3})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
1 $(a_{3},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$
The Pareto-efficient allocations here are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$. Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$ and $z$ (these are dominated by $p$ and $y$, respectively). Yet $V^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$. $\blacklozenge$
Our last example, finally, shows that a strictly monotonic profile that is not homeo-monotonic may fail to have an intensity-efficient allocation.
Example 4.
For $n=5$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4},v_{5}\},\{a_{4},a_{5}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\}$, $\{a_{1},a_{2},a_{3}\})$, as follows:
$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{1}$\hskip 20.0pt}}$
1 2 3    4    5
$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}\succ_{i}a_{5}$ $a_{4}\succ_{4}\cdots$ $|$ $a_{5}\succ_{5}\cdots$
10 $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{4},\cdot)$ $(a_{5},\cdot)$
9 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $\vdots$ $\vdots$
8 $(a_{2},a_{5})$ $(a_{2},a_{5})$ $(a_{1},a_{3})$
7 $(a_{2},a_{4})$ $(a_{1},a_{3})$ $(a_{2},a_{5})$
6 $(a_{1},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$
5 $(a_{3},a_{5})$ $(a_{3},a_{5})$ $(a_{3},a_{5})$
4 $(a_{1},a_{2})$ $(a_{3},a_{4})$ $(a_{3},a_{4})$
3 $(a_{2},a_{3})$ $(a_{1},a_{2})$ $(a_{4},a_{5})$
2 $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
1 $(a_{4},a_{5})$ $(a_{4},a_{5})$ $(a_{2},a_{3})$
The Pareto-efficient allocations here are
$\begin{array}[]{lllllll}\underbrace{(c,a,b,d,e)}_{o},&\underbrace{(c,b,a,d,e)}_{t},&\underbrace{(a,b,c,d,e)}_{x},&\underbrace{(a,c,b,d,e)}_{y},&\underbrace{(b,c,a,d,e)}_{w},&\underbrace{(b,a,c,d,e)}_{z}\end{array}$
Yet the intensity-dominance relation $D$ is cyclic over this set:
[…]
$\blacklozenge$
Together, Theorems 2-3 and Examples 2-4 highlight the importance of the intuitive structure imposed by homeo-monotonicity for the general existence of intensity-efficient allocations, both as a sufficient condition and, in a weaker sense, as a necessary one too. As the counterexample for $n=4$ in the proof of Theorem 1 shows, moreover, a profile may fail to have an intensity-efficient allocation even when homeo-monotonicity as defined above is satisfied, yet a subtle variation of it is not. More specifically, that profile features: (i) $n=4$; (ii) $a\succ_{i}b\succ_{i}c\succ_{i}d$ for $i\leq 4$ (hence a top-preference-coincident set $V$ with respect to $A$); (iii) $s_{1}=s_{4}$, $s_{2}=s_{3}$. Although the four agents’ intensities over their common preferences do not define a monotonic top-preference-coincident set, one observes that emergent here is a variant of this notion whereby each set in the partition $\{\{v_{1},v_{2}\},\{v_{3},v_{4}\}\}$ of $V$ is monotonically top-preference-coincident with respect to $\{a,b,c\}\subset A$: $s_{1}(a,b)>s_{2}(a,b)$, $s_{1}(b,c)=s_{2}(b,c)$ and $s_{4}(a,b)>s_{3}(a,b)$, $s_{4}(b,c)=s_{3}(b,c)$. However, both these pairs of agents violate what homeo-monotonicity would require here: $s_{1}(a,c)<s_{2}(a,c)$ and $s_{4}(a,c)<s_{3}(a,c)$. Thus, while the problem of fully characterizing the conditions under which strict profiles have intensity-efficient allocations remains open, the preceding analysis hopefully provides non-trivial and useful guidance in that direction.
Figure 1: Simulations-based histograms showing the proportions of Pareto-efficient allocations that are also intensity-efficient when the number of agents and items is small.
Additional motivation for further study is also contributed by the analysis of uniform-randomly drawn intensity profiles when $n=3,4,5$. Figure 1 summarizes this output and displays how the distribution of the per-profile proportion of Pareto-efficient allocations that are also intensity-efficient varies with $n$. Interestingly, while the average number of Pareto-efficient allocations increases rapidly from 2.8 to 5.9 to 14.1 when $n=3,4,5$, respectively, the corresponding averages for intensity-efficient allocations are 1.6, 1.9 and 2.4. The refinement gains afforded by intensity-efficiency are therefore significant and appear to be increasing in $n$, with the mean proportion of intensity-efficient out of all Pareto-efficient allocations falling from 0.63 ($n=3$) to 0.2 ($n=5$). Finally, non-existence of intensity-efficient allocations is a rare occurrence in these simulations, with only 1 and 15 out of the respective 100,000 randomly drawn profiles being associated with a cyclic intensity-dominance relation over Pareto-efficient allocations when $n=4$ and $n=5$, respectively.
We end this section by remarking that the paper’s new analytical environment necessitated novel arguments in the proofs of this section’s main results. More specifically, viewed as an abstract mathematical problem, existence of intensity-efficient allocations amounts to identifying sufficient conditions for acyclicity of the intensity-dominance binary relation defined on the set of permutations over alternatives. The mathematical literature on the combinatorics of permutations has studied partial orders over such sets, which may even have the richer structure of a lattice that additionally ensures the existence of both a greatest (hence maximal) and a smallest element. These results, however, are not applicable in our environment. The reason is that those partial orders, namely the strong and weak Bruhat orders (Stanley, 2012, pp. 399-400), relate two permutations if and only if there is a single transposition (“flip”) between elements in the two permutations, and this transposition corresponds to an inversion (Stanley, 2012, p. 30) relative to some postulated fixed linear order over its elements. The weak Bruhat order (see Ceballos and Pons (2024) for a recent generalization) gives rise to a lattice—whose graph is known as the permut(a)(o)hedron—but under the additional assumption that the said single transposition occurs between adjacent elements of the two relevant permutations. Our problem on the other hand does not assume a fixed linear order over items and is defined by intensity-dominance comparisons that generally feature possibly multiple transpositions of non-adjacent elements, which are defined by dominance between specific entries in the two relevant columns of an integer matrix, defined in turn by the corresponding agents’ intensity orderings.
5 Eliciting Ordinal Intensities and Finding Intensity-Dominating Allocations
We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems. In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency in the intensity-dominating direction. Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational. To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
Proposition 3.
Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in exactly $\frac{1}{2}(n-1)(n-2)$ steps, hence in $O(n^{2})$ time.
Proof.
Define $f(n):=\frac{1}{2}(n-1)(n-2)$. Observe that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains
$|\dot{\succsim}_{l}|$ $=$ $2{n\choose 2}+n$ (31)
comparisons between pairs in $A\times A$. We proceed by induction.
Let $n=3$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$. By (31), $|\dot{\succsim}_{l}|=9$. By (6) and the assumption that $\succ_{l}$ is induced by $\dot{\succsim}_{l}$ we readily get $(a_{1},a_{3})\dot{\succsim}_{l}(a_{1},a_{2})$ and $(a_{1},a_{3})\dot{\succsim}_{l}(a_{2},a_{3})$. By (5), moreover, we have $(a_{1},a_{1})\dot{\sim}_{l}(a_{2},a_{2})\dot{\sim}_{l}(a_{3},a_{3})$. By (5), $(\mathbf{5}_{D})$ and $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$ we also get $(a_{1},a_{2})\dot{\succ}_{l}(a_{i},a_{i})$ and $(a_{2},a_{3})\dot{\succ}_{l}(a_{i},a_{i})$, $i\leq 3$. By (5) and the above comparisons, finally, we also have $(a_{i},a_{i})\dot{\succ}_{l}(a_{2},a_{1})\dot{\succ}_{l}(a_{3},a_{1})$ and $(a_{i},a_{i})\dot{\succ}_{l}(a_{3},a_{2})\dot{\succ}_{l}(a_{3},a_{1})$. Together, these account for 7 out of the 9 comparisons in $\dot{\succsim}_{l}$, with those between $(a_{1},a_{2})$, $(a_{2},a_{3})$ and between $(a_{3},a_{2})$, $(a_{2},a_{1})$ being the two missing ones. By (5) once again, knowing either the first or the second such comparison also pins down the other. Hence, $f(3)=1$ is indeed the number of steps needed to elicit $\dot{\succsim}_{l}$ from $\succ_{l}$ when $n=3$.
Now let $n=4$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}\succ_{l}a_{4}$. By the preceding arguments we know that, out of the $|\dot{\succsim}_{l}|=16$ total comparisons here, missing are the following 3: $(a_{1},a_{2})$ vs $(a_{2},a_{3})$; $(a_{2},a_{3})$ vs $(a_{3},a_{4})$; $(a_{1},a_{3})$ vs $(a_{2},a_{4})$. Hence, $f(4)=3$ is the number of necessary steps here too.
For the induction step of the proof when $n>4$, suppose $f(m)$ is the correct number for every $n\leq m$. We will show that $f(m+1)$ is the correct number when $n=m+1$. To this end, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m}$ and notice that, from the induction hypothesis and the information derived from the $n=3$ and $n=4$ cases above, we can decompose the $f(m)$ comparisons as follows:
(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-2$ (m2m-2 comparisons);
(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-3$ (m3m-3 comparisons);
(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-4$ (m4m-4 comparisons);
$\cdots$
(m2m-2):  $(a_{i},a_{i+m-1})$ vs $(a_{i+1},a_{i+m})$ (1 comparison).
Indeed, summing the terms inside brackets across the right hand sides of $(1)\cdots(m-2)$ we get
$\sum\limits_{i=2}^{m-1}(m-i)$ $=$ $\sum\limits_{i=1}^{m-1}(m-i)-(m-1)$
$=$ $\frac{1}{2}(m-1)m-(m-1)$
$=$ $(m-1)\left(\frac{1}{2}m-1\right)$
$=$ $\frac{1}{2}(m-1)(m-2)$
$=$ $f(m)$
Consider now $n=m+1$ and, once again without loss, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m+1}$. From the preceding decomposition of the required comparisons when $n=m$ we can deduce that, for $n=m+1$,
(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-1$ (m1m-1 comparisons);
(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-2$ (m2m-2 comparisons);
(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-3$ (m3m-3 comparisons);
$\cdots$
(m1m-1):  $(a_{i},a_{i+m})$ vs $(a_{i+1},a_{i+m+1})$ (1 comparison).
Summing across categories now yields $\sum\limits_{i=1}^{m}(m-i)=\frac{1}{2}(m-1)m=f(m+1)$. ∎
Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available. The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm. The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run. TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment. Shapley and Scarf (1974) proved that TTC leads to a core allocation. Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique. Extending the results in Roth (1982), Ma (1994) later established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes. This mechanism is not strategy-proof. Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism. The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available. These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority under some ranking over agents.
Bogomolnaia and Moulin (2001), finally, demonstrated that, despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, these are generally ex ante inefficient in the sense that, based on the stochastic-dominance relation defined by all agents’ preferences, they may be dominated by another random allocation. In response to this realization, the authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
Informed by the main insights of the above-cited literature, we now consider the following algorithm for the problem at hand:
Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
Step 1: Choose a linear order $\gg$ over $V$ from the uniform distribution with support $n!$.
Step 2: Apply RP under $\gg$ on $P_{S}$ to arrive at a Pareto-efficient allocation $x$.
Step 3: Reorder the entries in $x$ according to $\gg$, with $j>i$ for $x_{i},x_{j}$ if and only if $i\gg j$.
Step 4: Search for the $\gg$-lowest agent $v_{j}\in\{v_{2},\ldots,v_{n}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$. If such an agent exists, swap $x_{1}\leftrightarrow x_{j}$ and remove $v_{1},v_{j}$ from $V$ and $x_{1},x_{j}$ from $A$.
If no such agent exists, remove $v_{1}$, $x_{1}$.
Repeat Step 4 until no agents remain.
Proposition 4.
If $S\in\mathcal{\widehat{S}}$ and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm terminates in $O(n^{2})$ time and either produces the Pareto-efficient allocation that it finds in its RP-implementing Step 2 or an allocation that intensity-dominates it.
Proof.
Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step. In the worst-case scenario, Step 4 always—with the possible exception of the agent ranked second-last by $\gg$—performs exhaustive search and removes one agent and one object in every iteration: Starting with $v_{1},x_{1}$, after n1n-1 searches it removes $v_{1},x_{1}$. Continuing with $v_{2},x_{2}$, after n2n-2 searches it removes $v_{2},x_{2}$. $\cdots$ Continuing with $v_{n-1},x_{n-1}$, after one search it removes both $v_{n-1},x_{n-1}$ and $v_{n},x_{n}$, either by implementing the swap $x_{n-1}\leftrightarrow x_{n}$ or otherwise. Worst-case performance is therefore associated with a total number of searches equalling
$\sum_{m=1}^{n-1}(n-m)$ $=$ $n^{2}-n-\sum_{m=1}^{n-1}m$
$=$ $n^{2}-\sum_{m=1}^{n}m$
Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in $O(n^{2})$ time.
It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$. Suppose $n=3$ first. Write $x:=(x_{1},x_{2},x_{3})$. By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap. Thus, $y$ must be one of the following:
(i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$
(ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$
(iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$
In each case, $yDx$. Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic. Denote by $(V^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $V$ and $A$ into a (possibly empty) top-preference diverse set $V^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $V^{i>1}$ with respect to $A^{i>1}$, $i\leq k$. By the properties of these sets profiles, and by the definition of algorithm’s Step 4, irrespective of the $\gg$-order on $V$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in V^{1}$. Thus, $y_{i}=x_{i}$ for all $v_{i}\in V^{1}$. By the properties of $(V^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in V^{l>1}$ and $v_{j}\in V^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct. It follows that $y$ differs from $x$ by at most $\frac{n}{2}$ or $\frac{n-1}{2}$ pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such swap $x_{i}\leftrightarrow x_{j}$ occurring within the same top-preference-coincident pair $(V^{l>1},A^{l>1})$. By Step 4, finally, every such swap reassigns the jointly preferred (by Pareto-efficiency of $x$) item within the relevant pair in the direction of intensity-dominance. Hence, $yDx$. ∎
The allocation delivered by this algorithm is not intensity-efficient in general. This is so because the forced exit of agents and items in the fourth step—a crucial feature for the algorithm’s manageable quadratic-time complexity—means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation. Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP. It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.
In light of the discussion preceding the statement of Proposition 4, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency. Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 3, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully. The answer is negative. To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$. The possible orderings over $V$ are: 1. $v_{1}\gg v_{2}\gg v_{3}$; 2. $v_{1}\gg v_{3}\gg v_{2}$; 3. $v_{2}\gg v_{1}\gg v_{3}$; 4. $v_{2}\gg v_{3}\gg v_{1}$; 5. $v_{3}\gg v_{1}\gg v_{2}$; 6. $v_{3}\gg v_{2}\gg v_{1}$. The two Pareto-efficient allocations are $(a,b,c)$ and $(b,a,c)$, and the former intensity-dominates the latter. Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with probability $\frac{1}{2}$). If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it. However, if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e. with probability $\frac{1}{2}$) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem. Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up. Abdulkadiroğlu et al. (2011), however, followed by Pycia (2011); Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for. In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools. In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible. In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.

### v6

@@
-Intensity-Efficient Allocations
+Intensity-Efficient and Intensity-Positional Allocations
 Abstract
-This paper proposes a refinement of Pareto-efficient allocations for situations where, in addition to having ordinal preferences, agents also have ordinal intensities: they can make comparisons such as “I prefer $a$ to $b$ more than I prefer $c$ to $d$, without necessarily being able to quantify them.
-A rank-based criterion for interpersonal comparisons of such ordinal intensities is introduced for this new analytical environment.
-Building on this, an allocation is defined to be intensity-efficient if it is Pareto-efficient with respect to the agents’ preferences and also such that when another Pareto-efficient allocation assigns the same pairs of items to the same pairs of agents but in a “flipped” way, the former allocation assigns the commonly preferred item in every such pair to the agent who prefers it more.
-Conditions are established under which such Pareto-refining allocations exist.
-The potential usefulness of this theory in matching problems is illustrated with a quadratic-time extension of the Random Priority (RP) algorithm that returns an allocation which intensity-dominates RP’s Pareto-efficient one.
+This paper studies the allocation problem in environments where, in addition to having ordinal preferences, agents also have ordinal intensities: they can say “I prefer $a$ to $b$ more than I prefer $c$ to $d$ without necessarily being able to claim how much more.
+A rank-based criterion for interpersonal comparisons of ordinal intensities is introduced for this analytical environment.
+Building on it, the concepts of intensity-dominated, intensity-efficient and intensity-positional allocations are defined.Two allocations are related by intensity dominance if they assign the same two items to the same two agents but in a “flipped” way, and one allocation assigns the commonly preferred item in any such pair to the agent who prefers it more.
+An allocation is intensity-efficient if it is Pareto-efficient and not intensity-dominated.
+Distinctly, an allocation is intensity-positional if it maximizes the agents’ total intensity-rank scores.
+The latter extend standard Borda scores to environments where consecutively-ranked items do not necessarily feature identical intensity differences.
+The paper analyses intensity-efficient and intensity-positional allocations, and clarifies how they differ from those based on the utilitarian or Borda rules.
 “Suppose I am left with a ticket to a Mozart concert I am unable to attend and decide to give it to one of my closest friends.
@@
 Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered.
-Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion formalized in Harsanyi (1955) is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Roemer, 1996; Moulin, 2004).
-Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Dietrich, 2025; Samuelson, 2025).
+Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion, formalized in Harsanyi (1955), is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Moulin, 1988, 2004).
+Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Alaoui and Penta, 2025; Samuelson, 2025).
 More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982).
@@
 551).
-To the extent that comparing options is easier and more reliable than assigning values to them, as the above survey and several studies cited therein suggest, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
+To the extent that comparing options is easier and more reliable than assigning values to them, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
 Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made.
@@
 collections of quaternary relations that reflect the agents’ preference intensity comparisons.
-This analytical choice assumes (and this assumption is validated in Section 5, Proposition 3) that, similar to the way in which preferences are elicited, e.g.
+This analytical choice assumes that, similar to the way in which preferences are elicited, e.g.
 in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
-This new and general framework where both preference and preference-intensity information is available to the social planner naturally raises the question of how such information might be used—as per the goals that we set out above—to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ preference intensities and mitigate the multiplicity of Pareto-efficient allocations.
-Similar to utilitarianism and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons.
-Unlike those notions, however, where such comparisons are based on agents’ utilities (Baccelli, 2023; Echenique et al., 2023; d’Aspremont, 2018; Fleurbaey and Hammond, 2004), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
+This framework where both preference and—generally only ordinal—intensity information is available to the social planner naturally raises the question of how it might be used, as per the goals that we set out above, to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ intensities and mitigate the multiplicity of Pareto-efficient allocations.
+Similar to utilitarian and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons.
+Unlike those notions, however, where such comparisons are based on agents’ utilities (Fleurbaey and Hammond, 2004; d’Aspremont, 2018; Echenique et al., 2023; Baccelli, 2023), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
 To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful.
 In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more.
-In Section 3 we formalize and defend this Interpersonal Comparisons of Ordinal Intensities new postulate/informational basis for social choice and welfare.
-We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (Baccelli, 2023; d’Aspremont, 2018).
+In Section 3 we formalize and motivate this interpretational assumption concerning interpersonal comparisons of ordinal intensities in the present analytical environment.
+We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781; Young, 1974; Mas-Colell et al., 1995; Maskin, 2025) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (d’Aspremont, 2018; Baccelli, 2023).
 In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just.
@@
 when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more.
-Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile, without further restrictions, when there are three agents and alternatives.
-Yet we also show that with four or more existence is not guaranteed because the underlying dominance relation may be cyclic without further assumptions.
-For the case of $n$ agents and items we then establish a sufficient condition which requires profiles to be (effectively) homeo-monotonic and whose key part could be summarized informally thus: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any two consecutively-ranked items in this subset, then this ordering cannot be violated in how much they prefer any non-consecutively ranked items.
-Furthermore, although this condition is not necessary for existence, we prove that a weaker version of it is indeed satisfied by every profile that has an intensity-efficient allocation.
+Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile when there are three agents and alternatives.
+We also show that existence is not guaranteed when there are more agents and items, because the underlying dominance relation may be then be cyclic.
+Motivated by this, for the general case of $n$ agents and as many items we establish a sufficient condition for existence, requiring profiles to be (effectively) homeo-monotonic.
+The key part of this condition could be summarized informally as follows: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any item in this subset over that item’s immediate successor, then this ordering cannot be violated when it comes to comparing the agents’ intensities between any non-consecutively ranked items in this subset.
+Importantly, although the presence of such groups—and the corresponding restrictions within them—is not necessary for existence, we show that if a profile does give rise to such a group, and if the above-mentioned ordering over its agents is strict, then homeo-monotonicity in that group is indeed necessary for an intensity-efficient allocation to exist.
 Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible.
-Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide, although they do so under an intuitive additional condition that we introduce.
-The potential discrepancy between these two intensity-incorporating welfare concepts that may arise in such cardinal environments is itself of potential usefulness in applications, because either notion can be used as a secondary selection criterion in those situations where the other one predicts multiple allocations.
-In Section 5, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon Pareto-efficiency in the direction of intensity-efficiency.
-We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited efficiently (specifically, in quadratic time) by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class we study.
-Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) one and produces—in quadratic time in the number of items/agents—an allocation that intensity-dominates the Pareto-efficient allocation that is found by RP.
+Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons do emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide.
+They do so, however, when the said utility profile satisfies an intuitive additional condition that we introduce.
+This requires any interpersonal dominance that emerges by comparing two agents’ utility differences to be accompanied by an interpersonal dominance in the same direction in the rank-order positions of these utility differences in the respective agents’ intensity orderings.
+In Section 5 we build on a stronger interpersonal comparisons assumption to propose and analyse the concept of intensity-positional allocations.
+Toward defining them, we recall first than an item’s Borda score for some agent is the number of items that are ranked below it by that agent.
+Although Borda scores give rise to a specific normalized ordinal utility representation of the agent’s preferences, in practice they are often used with a cardinal interpretation (see, for example, Maskin, 2025) due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where the agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart.
+We will refer to these intensity relations as “linear”.
+Similar to Borda allocations, which are defined as those that maximize the aggregate Borda score, intensity-positional allocations are defined as the maximisers of the aggregate intensity-rank score.
+Unlike an item’s Borda score, however, which reveals the position of that item on the agent’s preference ordering, its intensity-rank score corresponds to the highest rank-order position in the agent’s intensity ordering where this item appears as preferred to something else.
+Hence, it encodes information both about the number of alternatives that are ranked below this item and about the relative intensity differences between it and the items it is preferred to.
+This information, in turn, is reflected in the prescribed allocation process.
+In the paper’s focal case of strict intensity profiles we show that intensity-positional allocations refine standard Borda allocations.
+We further show that, in a larger domain where a like-for-like comparison remains meaningful, intensity-rank and Borda-score aggregation diverge in general but coincide at linear profiles.
+The intuitive reason is that, unlike intensity-positional aggregation, Borda aggregation treats every profile as it if were linear.
+Thus, the intensity-positional criterion is a genuine generalization of the standard Borda-count criterion in this extended domain.
+Intensity-positional and intensity-efficient allocations are logically distinct.
+Two advantages of the former relative to the latter are their generic existence and their incorporation of a way to resolve intensity-pertinent welfare trade-offs across agents.
+A disadvantage is that making intensity-positional allocations operational implicitly invokes stronger interpersonal comparability assumptions.
+In Section 6, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon the Pareto-efficiency desideratum.
+We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited in quadratic time by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class that we focus on.
+Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) mechanism’s algorithm and produces—in quadratic time in the number of items/agents—an allocation which intensity-dominates the Pareto-efficient one that is found by RP.
 While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm.
-We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and which focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
-1.1 Related Literature
-This paper’s contribution lies at the intersection of welfare economics, social choice and matching theory.
-Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Kaneko and Nakamura, 1979; Roemer, 1996; Moulin, 2004), Pareto-efficient egalitarian-equivalence (Pazner and Schmeidler, 1978) or relative egalitarianism (Kalai and Smorodinsky, 1975; Imai, 1983; Sprumont, 2013), the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process.
-Unlike these refinements, the hereby proposed one is the first that is operational in the general class of ordinal preference-intensity profiles, and builds on a novel interpersonal-comparability criterion that is distinct from interpersonal comparisons of utility differences, which the above notions rely on.
+We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
+As far as the broader related literature is concerned, this paper’s contribution lies at the intersection of welfare economics, social choice and matching theory, in all three cases building bridges between the existing ordinal- and cardinal-welfarist paradigms.
+Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Moulin, 2004) and others, the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process.
+Unlike these refinements, the hereby proposed ones are the first that are operational in the general class of ordinal preference-intensity profiles, and build on novel interpersonal-comparability criteria that are distinct from interpersonal comparisons of utility differences which the above notions rely on.
 Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain.
-The latter problem is important and deserves separate treatment.
-Intensity-efficient allocations and their domain of application are also distinct from those dictated by the Borda scoring rule (Borda, 1781; Young, 1974; Maskin, 2025).
-Although this rule is defined by a particular normalization of agents’ ordinal utility functions over alternatives, in practice it is often used with a cardinal interpretation due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where every agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart.
-The class of ordinal intensity relations that we consider does not impose this restriction.
-Yet at the same time, for the main results of this paper we assume a strictness condition on agents’ intensities which rules out such “linear” intensity profiles.
-In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs in the allocation process toward maximizing utilitarian welfare.
+Intensity-positional allocations on the other hand are indeed derived from a social welfare function.
+The problem of characterizing this function axiomatically is important and deserves separate treatment.
+In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs into the allocation process toward maximizing utilitarian welfare.
 Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer.
 While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
-2 Notation, Preliminaries and Measurement-Theoretic Foundations
-By $A:=\{a_{1},\ldots,a_{n}\}$ and $V:=\{v_{1},\ldots,v_{n}\}$ we denote, respectively, the finite sets of choice alternatives and agents.
+2 Measurement-Theoretic Foundations
+By $A:=\{a_{1},\ldots,a_{n}\}$ and $N:=\{1,\ldots,n\}$ we denote, respectively, the finite sets of choice alternatives and agents.
 To simplify the assignment problem under study we assume that these sets have equal sizes.
@@
 $(A\times A)_{\dot{\sim}_{l}}$ $:=$ $\{[a,b]:[a,b]\text{ is an intensity-equivalence class under }\dot{\sim}_{l}\}$ (2)
-The definitions that follow omit existential quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
+The definitions that follow omit universal quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
 Definition 1.
-A relation $\dot{\succsim}_{l}$ has a utility-difference representation if there is a function $u_{l}:A\rightarrow\mathbb{R}$ that satisfies
+[…]
 $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $u_{l}(a)-u_{l}(b)\geq u_{l}(c)-u_{l}(d)$ (3)
@@
 145).
-Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a subset of utility-difference comparisons whose order is indeed preserved by all ordinal transformations of the original representation.
-In our environment, whenever $u_{l}$ in (1) is not cardinal it has the complex uniqueness property of additive representations on finite sets.
+[^1] More specifically, it has the complex uniqueness property of additive representations on finite sets, articulated in Krantz et al.
+(1971, Theorem 2, p.
+431).
+Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a certain subset of utility-difference comparisons whose order is, in fact, preserved by all ordinal transformations of the original representation.
 Scott (1964), completing and extending the analysis that was initiated in Scott and Suppes (1958), characterized the existence of a utility-difference representation over a finite set by means of the following three axioms on $\dot{\succsim}_{l}$:
@@
 $(\mathbf{2_{\mathit{D}}})$ [Cancellation].
-$(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
+$(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for all $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
 $(\mathbf{3_{\mathit{D}}})$ [Reversal]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ implies $(d,c)\,\dot{\succsim}_{l}\,(b,a)$.
@@
 Definition 2.
-A relation $\dot{\succsim}_{l}$ is representable by a preference intensity function if there is a mapping $s_{l}:A\times A\rightarrow\mathbb{R}$, unique up to a strictly increasing transformation, such that
-$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(c,d)$ (4)
-$(a,b)\,\dot{\succsim}_{l}\,(b,a)$ $\Longleftrightarrow$ $s_{l}(a,b)\geq s_{l}(e,e)\geq s_{l}(b,a)$ (5)
-$\min\{s_{l}(a,b),s_{l}(b,c)\}\geq s_{l}(e,e)$ $\Longrightarrow$ $s_{l}(a,c)\geq\max\{s_{l}(a,b),s_{l}(b,c)\},$ (6)
-with a strict inequality on the left hand side of (6) implying a strict inequality on the right.
-In words, $s_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)].
+[…]
+$(a,b)\,\dot{\succsim}_{l}\,(c,d)$ $\Longleftrightarrow$ $\widehat{s}_{l}(a,b)\geq\widehat{s}_{l}(c,d)$ (4)
+$(a,b)\,\dot{\succsim}_{l}\,(b,a)$ $\Longleftrightarrow$ $\widehat{s}_{l}(a,b)\geq\widehat{s}_{l}(e,e)\geq\widehat{s}_{l}(b,a)$ (5)
+$\min\{\widehat{s}_{l}(a,b),\widehat{s}_{l}(b,c)\}\geq\widehat{s}_{l}(e,e)$ $\Longrightarrow$ $\widehat{s}_{l}(a,c)\geq\max\{\widehat{s}_{l}(a,b),\widehat{s}_{l}(b,c)\},$ (6)
+with a strict inequality on the left part of (6) implying a strict inequality on the right.
+In words, $\widehat{s}_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)].
 Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
@@
 We denote by $\mathcal{I}$ the class of intensity relations that admit this ordinal representation.
-The (pseudo-)cardinal utility-difference model is nested by the ordinal model in the special case where “lateral consistency” [(6)] is strengthened to “additivity” [(7)]:
+The ordinal model nests the (pseudo-)cardinal utility-difference model in the special case where “lateral consistency” [(6)] is strengthened to “additivity” [(7)]:
 Lemma 1.
-A relation $\dot{\succsim}_{l}$ that is represented by $s_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widehat{s}_{l}$ of $s_{l}$ satisfies
-$\widehat{s}_{l}(a,c)$ $=$ $\widehat{s}_{l}(a,b)+\widehat{s}_{l}(b,c)$ (7)
-Proof.
-Upon viewing $\widehat{s}_{l}$ as a general function of two variables, (7) is known as Sincov’s functional equation, whose solution is some $f:A\rightarrow\mathbb{R}$, unique up to an additive constant, such that $\widehat{s}_{l}(a,b)\equiv f(a)-f(b)$ (Aczél, 1966).
-This establishes the “if” part.
-The “only if” part is immediate upon defining $\widehat{s}_{l}(a,b):=u_{l}(a)-u_{l}(b)$ for the postulated function $u_{l}$ with the requisite representation property.
-∎
-The normalization of $s_{l}$ introduced next is important for the rest of this paper.
+A relation $\dot{\succsim}_{l}$ that is represented by $\widehat{s}_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widetilde{s}_{l}$ of $\widehat{s}_{l}$ satisfies
+$\widetilde{s}_{l}(a,c)$ $=$ $\widetilde{s}_{l}(a,b)+\widetilde{s}_{l}(b,c)$ (7)
+The proof of this and most other results appear in the Appendix.
+The normalization of $\widehat{s}_{l}$ introduced next is important for the rest of this paper.
 Definition 3.
-A preference intensity function $s_{l}$ represents $\dot{\succsim}_{l}$ canonically if
-$s_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
+A preference intensity function $\widehat{s}_{l}$ represents $\dot{\succsim}_{l}$ canonically if
+$\widehat{s}_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
 with $k$ the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
+We will denote by $s_{l}$ the canonical intensity function that represents $\dot{\succsim}_{l}$.
 The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
@@
 Any relation $\dot{\succsim}_{l}\in\mathcal{I}$ admits a canonical preference intensity representation.
-Proof.
-Let $s_{l}$ be an arbitrary preference intensity function for $\dot{\succsim}_{l}\in\mathcal{I}$.
-Define the level set of $s_{l}$ at $(a,b)\in A\times A$ by $\mathbf{s}_{l}[a,b]:=\{(a^{\prime},b^{\prime})\in A\times A:s_{l}(a,b)=s_{l}(a^{\prime},b^{\prime})\}$.
-Further, define $r_{l}:A\times A\rightarrow\mathbb{R}$ by
-$r_{l}(a,b)$ $:=$ $\left\{\begin{array}[]{ll}|\{\mathbf{s}_{l}[a^{\prime},b^{\prime}]:s_{l}(a,b)>s_{l}(a^{\prime},b^{\prime})\geq s_{l}(e,e)\}|,&\text{if }s_{l}(a,b)>s_{l}(e,e)\\ s_{l}(a,b),&\text{if }s_{l}(a,b)=s_{l}(e,e)\\ -r_{l}(b,a),&\text{if }s_{l}(a,b)<s_{l}(e,e)\end{array}\right$.
-By construction, $r_{l}$ is canonical and $r_{l}(a,b)\geq r_{l}(c,d)\Leftrightarrow s_{l}(a,b)\geq s_{l}(c,d)$. ∎
 We assume throughout that every $\dot{\succsim}_{l}$ belongs to $\mathcal{I}$, hence is representable canonically.
@@
 It is therefore analogous to the standard preference-strictness postulate which, in fact, it implies.
-As we elaborate in the next section, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$.
-In particular, the only non-trivial intensity-equivalence class that exists when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
+As we elaborate in Section 3, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$.
+In particular, the only non-trivial intensity-equivalence class when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
+For illustration, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit,respectively, a (pseudo-cardinal) utility-difference representation and an ordinal representation when $n\leq 5$.
 Table 1: The number of distinct intensity relations of different classes for small values of $n$.
@@
 Strict representable ordinally representable intensity
-$n=$ orders intensity relations intensity relations profiles
+$n$ orders intensity relations intensity relations profiles
 3 6 12 12 $12^{3}$
@@
 5 120 13,680 92,160 $92,160^{5}$
-Finally, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit, respectively, a (pseudo-cardinal) utility-difference and an ordinal representation when $n\leq 5$, and was produced with Minizinc (Nethercote et al., 2007; Gecode Team, 2006).
-3 The Comparability Postulate
+Source: author’s computations with the Minizinc constraint-programming tool (Nethercote et al., 2007).
+We conclude this section by commenting on possible sources of intensity information.
+Section 2 in Gerasimou (2021) discusses various non-choice variables that, according to different literatures, might be thought of as having intensity-revealing potential.
+Prominent examples include willingness to pay; 50/50 lottery choices; survey-based ratings; and response times.
+Section 5 of that paper, moreover, shows how such data can be analysed alongside choices by means of otherwise standard revealed-preference theoretic techniques, providing testable foundations for the hypothesis that an agent’s overall behaviour complies with utility maximization and ordinal intensity revelation.
+The present paper, in contrast, operates within the environment of the allocation problem, where agents are traditionally assumed to submit their full ordinal or cardinal preference reports directly.
+This is different from a lab, field or other environment that might, in principle, provide the analyst with agents’ choices at several menus at different points in time, together with intensity-revealing information at those menus.
+As we elaborate formally in Section 6, for this specific environment we assume that agents’ ordinal preferences are available, and that their ordinal intensities are elicited by asking them directly—and conditional on their submitted preferences—the kinds of simple binary questions that were mentioned earlier.
+3 Interpersonal (Non-)Comparability of Ordinal Intensities and Intensity-Rank Scores
 […]
-Interpersonal Comparisons of Ordinal Intensities
+Assumption 1 (for Sections 4 & 5)
 Given an intensity profile $S\in\mathcal{\widehat{S}}$ and the corresponding canonical profile $s$, the statement
@@
 is interpreted as suggesting that agent $l$ prefers $a$ to $b$ more than agent $m$ prefers $c$ to $d$.
-Towards defending this new concept we first recall that preference intensities at the level of the individual decision maker are not assumed here to be quantifiable beyond the level of an ordinal ranking.
-Yet some information about the generally different welfare effects of having $a$ or $b$ vs $c$ or $d$ is available here.
-Suppose, in particular, that agent $l$ prefers $a$ to $b$ to $c$ to $d$, and the first over the second more than the third over the fourth.
+Towards defending this assumption—whose importance, we stress, is interpretational—recall first that an individual’s preference intensities are not assumed to be quantifiable beyond the level of an ordinal ranking.
+Yet some information about the generally different welfare effects of having $a$ or $b$ vs $c$ or $d$ is available.
+Suppose, in particular, that $a\succ_{l}b\succ_{l}c\succ_{l}d$, and that the agent prefers $a$ to $b$ more than she does $c$ to $d$.
 This is equivalent to saying that her intensity difference at pair $(a,b)$ is greater than at $(c,d)$.
 Consider agent $m$ next, and suppose for simplicity that the same applies to him too: $s_{m}(a,b)>s_{m}(c,d)$.
-Considering this, and that both have preferences and intensities over the same set of alternatives, how might (10) be interpreted?
-It is clearly a statement of fact that the intensity difference within pair $(a,b)$ is ranked higher by agent $l$ in her intensity ordering than the corresponding difference within pair $(c,d)$ is ranked by agent $m$ in his.
-Now recall that, by the Strictness assumption, $l$ and $m$ have the same number of possible rank positions in their respective intensity orderings which pairs $(a,b)$ and $(c,d)$ can occupy.
+Given this, and given also that both agents have preferences and intensities over the same set of alternatives, how might (9) be interpreted?
+We know that the intensity difference within pair $(a,b)$ is ranked higher by agent $l$ in her intensity ordering than the corresponding difference within pair $(c,d)$ is ranked by agent $m$ in his.
+By the Strictness assumption, moreover, $l$ and $m$ have the same number of possible rank positions in their respective intensity orderings, which pairs $(a,b)$ and $(c,d)$ could possibly occupy.
 In other words, the agents’ ordinal intensity scales coincide: if $(a,b)$ and $(c,d)$ are ranked 5th and 8th in $l$’s and $m$’s scales, respectively, then they are ranked 5th and 8th out of the same $k$ possible positions.
 As we mentioned in the Introduction, this logic resembles the way scoring rules such as the well-known Borda-count (Borda, 1781) normalization of ordinal utilities or scoring-like (see d’Aspremont, 2018) rules such as the relative-utilitarian (Dhillon and Mertens, 1999) normalization of cardinal utilities perform interpersonal comparisons in their respective environments.
-The question now emerges: Should the different agents’ intensity orderings be treated equally by the social planner?
-While equal treatment may or may not be an appropriate approach to follow in practice, depending on what else is known about agents and the problem at hand, since the intensity orderings are assumed here to encode all available welfare-relevant information, treating them in any way other than equal would call for a justification that appears elusive.
-Against this background, we view the suggested interpretation of (10) as a reasonable starting point for interpersonal comparisons in this informational environment.
-On the analytical side, another relevant question also arises: in the special case where the ordinal intensity ordering of every agent is defined by how the pairwise differences of a cardinal utility function are ordered, and these differences in turn define a strict intensity profile, does interpersonal comparisons of ordinal intensities reduce to the interpersonal comparability of utility differences that is required by utilitarian aggregation?
+Importantly, neither of these social welfare functions formally requires interpersonal utility comparisons in order for the respective social outcome to be immune to permissible transformations of the agents’ utility functions.
+Clearly, this is so because both welfare criteria are defined in terms of specific normalizations which reduce, respectively, any ordinally or cardinally equivalent utility function that represents an agent’s preferences into the same normalized ordinal or cardinal utility function.
+At the same time, however, applications of the Borda and relative-utilitarian normalizations do introduce interpersonal utility comparisons.
+We elaborate on this in the case of the Borda method, which is conceptually closer to our setting.
+This takes agent $l$’s preference relation $\succ_{l}$ and defines the Borda score of every item $a$ in $A$ by
+$b_{l}(a)$ $:=$ $|\{a^{\prime}\in A:a\succ_{l}a^{\prime}\}|,$ (10)
+i.e.
+by the number of alternatives that $l$ considers inferior to $a$.
+Once these normalised utility functions are constructed for all agents, the Borda social welfare function maps any such profile to a social ordering that prioritises items with higher aggregate Borda scores.
+Although interpersonal comparability assumptions are not part of the axioms that characterize the Borda method, the motivation for its very existence does implicitly invoke such comparisons.
+This is articulated clearly in Maskin (2025, f.17), as follows: “Notice that the Borda count implies a particular way of making interpersonal comparisons, e.g., if individual 1 ranks $x$ two positions above $y$, that preference is exactly canceled by two individuals who rank $y$ one position above $x$.
+Observe, however, that none of our axioms speaks directly to interpersonal comparisons at all; such comparisons are an emergent property of the joint imposition of the axioms”.
+Returning to Assumption 1, we note the analogy between the canonically-normalized values that possibly different pairs of items may occupy in the intensity orderings of different agents, and the possibly distinct positions (Borda scores) of items across agents’s preference orderings.
+However, unlike the additional cancellation assumptions that are implicit in the aggregation of Borda scores, Assumption 1 merely requires that an ordinal comparison be made about $l$’s intensity between $a$ and $b$ vs $m$’s intensity between $c$ and $d$.
+The question now emerges: Assuming that it is desirable in the first place for this additional information that is contained in the different agents’ ordinal intensity orderings to be accounted for, should these be treated equally by the social planner?
+While equal treatment may or may not be an appropriate approach to follow in practice, depending on what else is known about agents and the problem at hand, since the intensity orderings are assumed here to encode all available welfare-relevant information, accepting their relevance but treating them in any way other than equal would call for a justification that appears elusive.
+Against this background, we view the suggested interpretation of (9) as a reasonable starting point for interpersonal comparisons of ordinal intensities in this informational environment.
+Another relevant question arises, now on the analytical side: in the special case where the ordinal intensity ordering of every agent is defined by how the pairwise differences of a cardinal utility function are ordered, and these differences in turn define a strict intensity profile, does interpersonal comparability of ordinal intensities in the sense of Assumption 1 reduce to the interpersonal comparability of utility differences that is required by utilitarian aggregation?
 As the next example shows, the answer is no.
 Example 1.
-Let $n=3$ and consider the utility profile $U:=\big{(}(53,38,9),(46,33,21),$ $(20,30,50)\bigr{)}$ over $A=\{a_{1},a_{2},a_{3}\}$, where each $u_{l}$ in $U$, $l\leq 3$, is constructed so its values add up to 100 (this is without loss).
+Let $n=3$ and consider the utility profile $U:=\big((53,38,9),(46,33,21),$ $(20,30,50)\bigr)$ over $A=\{a_{1},a_{2},a_{3}\}$, where each $u_{l}$, $l\leq 3$, is constructed so its values add up to 100 (this is without loss).
 It is readily seen that the intensity profile defined by the intra-personal utility differences in $U$ is in $\mathcal{\widehat{S}}$.
-Also, defining $\dot{\succ}_{i}$ by $(a_{i},a_{j})\dot{\succ}_{l}(a_{k},a_{m})$ $\Leftrightarrow$ $u_{l}(a_{i})-u_{l}(a_{j})>u_{l}(a_{k})-u_{l}(a_{m})$, one observes that $(a_{2},a_{3})\,\dot{\succ}_{1}\,(a_{1},a_{2})$ and $(a_{1},a_{2})\,\dot{\succ}_{2}\,(a_{2},a_{3})$ hold, hence $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$ too.
-However, $u_{1}(a_{1})-u_{1}(a_{2})>u_{2}(a_{1})-u_{2}(a_{2})$ is also true.
+Also, defining $\dot{\succ}_{i}$ by $(a_{i},a_{j})\dot{\succ}_{l}(a_{k},a_{m})$ $\Leftrightarrow$ $u_{l}(a_{i})-u_{l}(a_{j})>u_{l}(a_{k})-u_{l}(a_{m})$, we observe that $(a_{2},a_{3})\,\dot{\succ}_{1}\,(a_{1},a_{2})$ and $(a_{1},a_{2})\,\dot{\succ}_{2}\,(a_{2},a_{3})$, hence $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$ too.
+Yet $u_{1}(a_{1})-u_{1}(a_{2})>u_{2}(a_{1})-u_{2}(a_{2})$ is also true.
 $\blacklozenge$
@@
 the Cardinal Unit Comparability informational basis—see also Section 4.2).
-4 Intensity Dominance and Intensity-Efficient Allocations
-4.1 Definitions
+Although the analysis of Section 4 makes use of Assumption 1222In fact, it uses a weaker version where interpersonal comparisons are made when $(a,b)=(c,d)$ in (9)., Section 5 revolves around concepts that become meaningful under a richer interpersonal comparability assumption on agents’ ordinal intensities.
+Before stating it we first use an agent’s canonical intensity function to introduce the intensity-rank score (or intensity rank) of item $a$ by
+$r_{l}(a)$ $:=$ $\max_{a^{\prime}\in A}s_{l}(a,a^{\prime})$ (11)
+In Lemma 5 (see Appendix) we show that $r_{l}$ is an ordinal utility function for $l$’s preferences that are induced by $\dot{\succsim}_{l}$.
+Recall now that, under Assumption 1, the comparison $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ is interpreted as suggesting that the intensity difference between $a$ and $b$ is at least as large for agent $l$ as the respective difference between $c$ and $d$.
+Logically, then, $r_{l}^{S}(a)>r_{l}^{S}(b)$ means that the maximum intensity difference between $a$ and some alternative that is inferior to it exceeds the maximum such difference that corresponds to $b$.
+In other words, the intensity-rank score of $a$ reflects the highest position in $l$’s intensity ordering where $a$ finds itself being preferred to another item.
+By its definition, therefore, and unlike the Borda score $b_{l}(a)$, this new score contains information both about how many alternatives are worse than $a$ in $l$’s preference ranking and about the intensity difference between $a$ and its inferior alternatives.
+Assumption 2 (for Section 5)
+Given an intensity profile $S\in\mathcal{\widehat{S}}$ and the corresponding canonical profile $s$, the statement
+$r_{l}(a)$ $>$ $r_{m}(b)$ (12)
+is interpreted as suggesting that agent $l$ prefers $a$ more than agent $m$ prefers $b$.
+To motivate this assumption and connect it to the previous one, suppose $r_{l}(a)\equiv s_{l}(a,a_{w})$ and $r_{m}(b)\equiv r_{m}(b,b_{w})$, where $a_{w}$ and $b_{w}$ are the bottom-ranked items in $A$ for agents $l$ and $m$, respectively.
+Let $r_{l}(a)>r_{m}(b)$.
+By definition, this is equivalent to $s_{l}(a,a_{w})>s_{m}(b,b_{w})$.
+Under Assumption 1, this means that $l$ prefers $a$ to their worst possible item more than $m$ prefers $b$ to their own worst possible item.
+Assumption 2, then, makes this last statement the definition of an interpersonal utility comparison between $a$ and $b$ for agents $l$ and $m$.
+Utility functions defined by Borda scores are onto the same consecutive integral range $\{0,1,\ldots,n-1\}$.
+By contrast, the extended—but still common across agents—range of utility functions defined by intensity-rank scores is $\{0,1,\ldots,{{n}\choose{2}}\}$.
+Those functions are obviously not onto this range, and their images are generally distinct, although, by definition, they all have the same minimum and maximum values.
+We will elaborate further on the relation between $b_{l}$ and $r_{l}$ in Section 5.
+We conclude this section by stressing that every welfare-theoretic concept that we introduce in the sequel relies on the position where a pair of alternatives lies within an agent’s intensity ranking.
+Thus, by introducing canonically normalized intensity functions at this early stage we assign a specific name and piece of notation for these score-like positions.
+Crucially, this means that the welfare concepts in later sections are invariant with respect to arbitrary monotonic transformations of any intensity functions that one may choose to represent the agents’ ordinal intensities by, because—similar to the Borda and relative utilitarian functions—these are automatically translated into canonically normalized values of ordinal intensity functions that are onto the same range for all agents.
+4 Intensity-Dominance and Efficiency
+4.1 Intensity-Dominance and Efficiency
 Since the number of agents and items are assumed to coincide, an allocation of the $n$ goods is a permutation on $A$.
 The set of all allocations is denoted by $\mathcal{A}$.
-As we discussed in the Introduction, we are interested in the assignment of the $n$ objects in $A$ to the $n$ agents in $V$ in a way that satisfies Pareto-efficiency but also improves upon it in delivering a narrower set of acceptable allocations that satisfy some additional desideratum that is appealing from a distributive-justice perspective and is operational in the present analytical environment where information on ordinal preference intensities and interpersonal comparisons thereof is assumed to be available.
+As we discussed in the Introduction, we are interested in the assignment of the $n$ objects in $A$ to the $n$ agents in $N$ in a way that satisfies Pareto-efficiency but also improves upon it in delivering a narrower set of acceptable allocations that satisfy some additional desideratum that is appealing from a distributive-justice perspective and is operational in the present analytical environment where information on ordinal preference intensities and interpersonal comparisons thereof is assumed to be available.
 To this end, we proceed with the introduction of the following novel notions of dominance and efficiency.
@@
 Let $S=\big(\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n}\bigr)\in\widehat{\mathcal{S}}$ be a strict intensity profile and $s$ its canonical representation.
-Given allocations $x$ and $y$, the former is said to intensity-dominate the latter if, for every pair of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$ it holds that $s_{i}(x_{i},x_{j})\geq s_{j}(y_{j},y_{i})$, and there is at least one pair $(i,j)$ where this inequality is strict.
+Given allocations $x$ and $y$, the former is said to intensity-dominate the latter if for every pair of agents $(i,j)$ such that $(x_{i},x_{j})=(y_{j},y_{i})$ it holds that $s_{i}(x_{i},x_{j})\geq s_{j}(y_{j},y_{i})$, and the inequality is strict for some pair.
 Definition 5.
-An allocation is intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ if it is Pareto-efficient at the induced preference profile $P_{S}\in\mathcal{P}^{V}$ and is not intensity-dominated by another Pareto-efficient allocation at $P_{S}$.
+An allocation is intensity-efficient at profile $S\in\widehat{\mathcal{S}}$ if it is Pareto-efficient at the induced preference profile $P_{S}\in\mathcal{P}$ and is not intensity-dominated by another Pareto-efficient allocation at $P_{S}$.
 It is intensity-dominant if it intensity-dominates every other Pareto-efficient allocation at $P_{S}$.
@@
 4.2 Intensity-Efficient and Utilitarian Allocations Are Distinct
-To enable a comparison between the concept introduced above and the benchmark Pareto-efficiency refinement afforded by classical utilitarianism, we assume throughout this subsection that, in addition to satisfying the strictness condition $(\mathbf{5_{\mathit{D}}})$, each $\dot{\succsim}_{l}$ is utility-difference representable as in (3) under some cardinal utility index $u_{l}$.
-This is the case, for example, when $u_{l}:A\rightarrow\mathbb{R}$ is the utility function associated with an expected-utility representation of preferences over lotteries over $A$ and the intensity order defined by the $u_{l}$-differences is strictly convex or strictly concave relative to the agent’s induced (strict) preferences.
-With this assumption in place, we finally invoke the Cardinal Unit Comparability (Roemer, 1996; Baccelli, 2023) informational basis to social welfare.
-More specifically, we assume that: (i) $u_{l}:A\rightarrow\mathbb{R}$ is unique up to a positive affine transformation and features utility differences that define an intensity relation $\dot{\succsim}_{l}$ that abides with $(\mathbf{5_{\mathit{D}}})$; (ii) interpersonal utility-difference comparisons are possible under such transformations of any utility profile $U=(u_{1},\ldots,u_{n})$ where each $u_{l}$ is of this kind, as long as the scale/multiplicative parameter of such transformations is the same for all $u_{l}$.
-As is well-known, this informational basis is sufficient [^1] The reason why it is not necessary is closely related to the complex uniqueness properties of general utility-difference representations in the sense of (3).
-See Bossert (1991) for more on this.
-for the classical utilitarian social welfare functional—recalled below—to be meaningful.
+[…]
 Fixing a utility profile $U$ with the above properties, we refer to it as strict if the ordering of every agent’s utility differences is strict (see Example 1).
@@
 The sets of intensity-efficient and classical utilitarian allocations associated with strict utility profiles are generally distinct.
-Proof.
-Consider the strict utility profile of Example 1.
-Observe that the Pareto-efficient allocations are $(a_{1},a_{2},a_{3})$ and $(a_{2},a_{1},a_{3})$, and that the former allocation is the utilitarian one.
-Yet this is intensity-dominated by the latter since $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$.
-∎
 This potential divergence between intensity efficiency and the hallmark cardinal-welfarist allocation criterion is important because it shows that the hereby proposed notion is indeed distinct even under the richest possible informational assumptions.
-Furthermore, while perhaps not obvious ex ante, as is clarified by Example 1 and the discussion at the end of in Section 3, the reason why such a divergence may occur is that the classical utilitarian criterion is guided by the utility differences across agents.
-By contrast, in this special environment where intensity orderings are defined by the relative ordering rather than the absolute magnitude of utility differences, the Interpersonal Comparisons of Ordinal Intensities postulate upon which the intensity efficiency criterion is built compares the relative ranking of those differences across agents’ intensity orderings.
+Furthermore, while perhaps not obvious ex ante, as is clarified by Example 1 and the discussion in Section 3, the reason why such a divergence may occur is that the classical utilitarian criterion is guided by the utility differences across agents.
+By contrast, in this special environment where intensity relations are defined by the relative ordering rather than the absolute magnitude of utility differences, the Interpersonal Comparisons of Ordinal Intensities postulate upon which the intensity efficiency criterion is built compares the relative ranking of those differences across agents’ intensity orderings.
 Naturally, one may inquire about the conditions under which classical utilitarian allocations are intensity-efficient.
@@
 Classical utilitarian allocations of strict and balanced utility profiles are intensity-efficient.
-Proof.
+It is well-known that utilitarian allocations are Pareto-efficient.
+Proposition 2 shows that, when the utility profile is strict and balanced, this benchmark criterion has the additional normative status conveyed by intensity-efficiency.
+Furthermore, Proposition 2 implies the following existence result that is perhaps worth stating explicitly:
+Corollary 1.
+An intensity-efficient allocation exists for every intensity profile that is induced by a strict and balanced utility profile.
+Example 1 (continued).
+Observe that the slightly modified strict utility profile in Example 1 where the second agent’s utility vector is marginally altered from $(46,33,21)$ to $(46,34,20)$ remains strict but now also becomes balanced.
+Under the modified profile the two Pareto-efficient allocations are both utilitarian and intensity-efficient.
+$\blacklozenge$
+4.3 Existence without Cardinality Assumptions
+We now return to the general environment in which intensity profiles come from the class of relations that were specified in Section 4.1, without imposing any further restrictions such as cardinal or even utility-difference representability of the individuals’ intensities.
+Theorem 1.
+An intensity-efficient allocation exists for every profile $S\in\mathcal{\widehat{S}}$ when $n=3$ but not in general when $n\geq 4$.
+Table 2: Example profile that has no intensity-efficient allocation.
+$s_{i}(a,b)=$ $i=$ 1 2 3 4
+6 $(a,d)$ $(a,d)$ $(a,d)$ $(a,d)$
+5 $(b,d)$ $(a,c)$ $(a,c)$ $(b,d)$
+4 $(a,c)$ $(b,d)$ $(b,d)$ $(a,c)$
+3 $(b,c)$ $(b,c)$ $(b,c)$ $(b,c)$
+2 $(a,b)$ $(c,d)$ $(c,d)$ $(a,b)$
+1 $(c,d)$ $(a,b)$ $(a,b)$ $(c,d)$
+We explain why non-existence is a potential issue when $n\geq 4$, as this will be useful for our subsequent analysis.
+To this end, consider the example strict intensity profile on $A=\{a,b,c,d\}$ whose canonical representation is shown in Table 2.
+Observe that $s_{1}=s_{4}$, $s_{2}=s_{3}$ and $a\succ_{i}b\succ_{i}c\succ_{i}d$, $i=1,\ldots,4$.
+This implies that all 24 possible allocations, defined and listed below, are Pareto efficient:
+$\begin{array}[]{llllllllllllllllll}x^{1}&=&(a,b,c,d)&&&x^{2}&=&(a,b,d,c)&&&x^{3}&=&(a,c,b,d)&&&x^{4}&=&(a,c,d,b)\\ x^{5}&=&(a,d,b,c)&&&x^{6}&=&(a,d,c,b)&&&x^{7}&=&(b,a,c,d)&&&x^{8}&=&(b,a,d,c)\\ x^{9}&=&(b,c,a,d)&&&x^{10}&=&(b,c,d,a)&&&x^{11}&=&(b,d,a,c)&&&x^{12}&=&(b,d,c,a)\\ x^{13}&=&(c,a,b,d)&&&x^{14}&=&(c,a,d,b)&&&x^{15}&=&(c,b,a,d)&&&x^{16}&=&(c,b,d,a)\\ x^{17}&=&(c,d,a,b)&&&x^{18}&=&(c,d,b,a)&&&x^{19}&=&(d,a,b,c)&&&x^{20}&=&(d,a,c,b)\\ x^{21}&=&(d,b,a,c)&&&x^{22}&=&(d,b,c,a)&&&x^{23}&=&(d,c,a,b)&&&x^{24}&=&(d,c,b,a)\end{array}$
+The following comparisons, whose validity can be readily established by the reader, demonstrate that for each $x^{i}$ in this set, $i=1,\ldots,24$, there is a distinct $x^{j}$ that intensity-dominates $x^{i}$—a situation denoted by $x^{j}Dx^{i}$.
+$\begin{array}[]{llllllll}x^{6}Dx^{1}&x^{1}Dx^{2}&x^{4}Dx^{3}&x^{13}Dx^{4}&x^{3}Dx^{5}&x^{17}Dx^{6}&x^{1}Dx^{7}&x^{2}Dx^{8}\\ x^{3}Dx^{9}&x^{8}Dx^{10}&x^{9}Dx^{11}&x^{11}Dx^{12}&x^{18}Dx^{13}&x^{13}Dx^{14}&x^{17}Dx^{15}&x^{22}Dx^{16}\\ x^{23}Dx^{17}&x^{24}Dx^{18}&x^{8}Dx^{19}&x^{22}Dx^{20}&x^{11}Dx^{23}&x^{12}Dx^{22}&x^{24}Dx^{23}&x^{10}Dx^{24}\end{array}$
+Therefore, no intensity-efficient allocation exists in this profile.
+The somewhat surprising fact that a plausible dominance concept may be cyclic and prevent an optimal entity to emerge invites an informal analogy between intensity-dominance cycles over allocations with at least 4 agents and Condorcet cycles over alternatives in pairwise majority-based preference aggregation with at least 3 agents (Condorcet, 1785).
+Importantly, however, although intensity-dominance cycles here may prevent refining the Pareto set, unlike the Condorcet social-welfare theoretic framework these do not lead to a “policy paralysis” problem because the Pareto set is always non-empty and, absent any distributively juster suggestions, one of them might be promoted by the social planner.
+That said, it is naturally of interest to understand better the conditions under which our proposed refinement of Pareto-efficiency is well-defined.
+We turn to this problem next.
+For any set $A^{\prime}\subseteq A$ and preference relation $\succ_{i}$ that is induced by an intensity relation $\dot{\succsim}_{i}\in\mathcal{I}$ we write $\succ_{i}^{A^{\prime}}$ for the restriction of $\succ_{i}$ on $A^{\prime}$.
+Definition 8.
+Given a profile $S\in\mathcal{\widehat{S}}$, a set of agents $N^{\prime}\subseteq N$ is said to be:
+  1. 1.
+    Envy-free with respect to a set of items $A^{\prime}\subseteq A$ if each $v_{i}\in N^{\prime}$ has a distinct $\succ_{i}$-best item in $A$, and this item (denoted $b_{i}^{\prime}$) belongs to $A^{\prime}$.
+  2. 2.
+    Top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ if $\succ_{i}^{A^{\prime}}=\succ_{j}^{A^{\prime}}$ for all $v_{i},v_{j}\in N^{\prime}$ and, when $A^{\prime}\subset A$, $a^{\prime}\in A^{\prime}$, $a\in A\setminus A^{\prime}$ holds, then $a^{\prime}\succ_{i}a$ for all $v_{i}\in N^{\prime}$.
+An envy-free set $N^{\prime}$ with respect to $A^{\prime}$ consists of agents in $N^{\prime}$ who have distinct overall-best items in $A^{\prime}$.
+This terminology is motivated by the fact that, if $N^{\prime}=N$ and $A^{\prime}=A$, then each agent receiving their most preferred option results in an allocation that is Pareto-efficient and also envy-free (Varian, 1974; Moulin, 2004).
+A top-preference-coincident set $N^{\prime}$ with respect to $A^{\prime}$ consists of agents whose preferences over $A^{\prime}$ are identical and such that every item in $A^{\prime}$ is preferred to everything in $A\setminus A^{\prime}$, with no further restrictions.
+We now introduce three interlinked concepts and illustrate them with two examples before explaining their content verbally.
+Definition 9.
+A set of agents $N^{\prime}\subseteq N$ that is top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ at profile $S\in\mathcal{\widehat{S}}$ is monotonic at $A^{\prime}$ if $|N^{\prime}|=|A^{\prime}|$ and
+$s_{1}(a_{1},a_{2})$ >> $s_{1}(a_{2},a_{3})$ >> $\cdots$ >> $s_{1}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
+$\geq$ $\geq$ $\cdots$ $\geq$
+$s_{2}(a_{1},a_{2})$ >> $s_{2}(a_{2},a_{3})$ >> $\cdots$ >> $s_{2}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
+$\geq$ $\geq$ $\cdots$ $\geq$
+$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
+$\geq$ $\geq$ $\cdots$ $\geq$
+$s_{|N^{\prime}|}(a_{1},a_{2})$ >> $s_{|N^{\prime}|}(a_{2},a_{3})$ >> $\cdots$ >> $s_{|N^{\prime}|}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
+It is strictly monotonic if every weak inequality above is replaced by a strict inequality.
+Finally, a monotonic or strictly monotonic such set is homeo-monotonic at $A^{\prime}$ if
+$s_{i}(a_{i},a_{j})$ $\geq$ $s_{j}(a_{i},a_{j})$ for all $1\leq i<j\leq|N^{\prime}|$.
+Definition 10.
+A profile $S\in\mathcal{\widehat{S}}$ is homeo-monotonic if, whenever $N$ and $A$ can be partitioned into subsets $N^{1},\ldots,N^{m}$ and $A^{1},\ldots,A^{m}$ where $N^{1}$ is envy-free with respect to $A^{1}$ and $N^{i>1}$ is monotonically top-preference-coincident with respect to $A^{i>1}$, then $N^{i>1}$ is homeo-monotonic with respect to $A^{i>1}$.
+Definition 11.
+A profile $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic if it induces partitions of agents and items relative to which it is homeo-monotonic.
+Remark 1.
+If $N$ is envy-free with respect to $A$, then $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
+Example 2.
+Table 3 describes an effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where $n=6$ and the envy-free and top-preference-coincident sets that partition $N$, AA are $(N^{1},A^{1})=(\{v_{1}\},\{a_{1}\})$, $(N^{2},A^{2})=(\{v_{2},v_{3},v_{4}\},\{a_{2},a_{3},a_{4}\})$, $(N^{3},A^{3})=(\{v_{5},v_{6}\},\{a_{5},a_{6}\})$.
+$\blacklozenge$
+Table 3: An effectively homeo-monotonic profile.
+$\overbrace{\hskip 7.0pt\text{$N^{1}$\hskip 7.0pt}}$ $\overbrace{\hskip 40.0pt\text{$N^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$N^{3}$}\hskip 20.0pt}$
+1 2 3 4 5 6
+$s_{i}(a,b)=$ $i=$ ${\color[rgb]{1,0,1}a_{1}}\succ_{1}\cdots$ ${\color[rgb]{1,0,0}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}}\succ_{i}\cdots$ ${\color[rgb]{0,0,1}a_{5}\succ_{i}a_{6}}\succ_{i}\cdots$
+15 $(a_{1},a_{6})$ $(a_{2},a_{1})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{5},a_{1})$ $(a_{5},a_{3})$
+14 $(a_{2},a_{6})$ $(a_{2},a_{6})$ $(a_{3},a_{5})$ $(a_{3},a_{6})$ $(a_{5},a_{4})$ $(a_{5},a_{1})$
+13 $(a_{1},a_{5})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{3})$ $(a_{5},a_{4})$
+12 $(a_{3},a_{6})$ $(a_{2},a_{4})$ $(a_{3},a_{6})$ $(a_{2},a_{1})$ $(a_{5},a_{2})$ $(a_{5},a_{2})$
+11 $(a_{2},a_{5})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$ $(a_{3},a_{1})$ $(a_{6},a_{1})$ $(a_{6},a_{3})$
+10 $(a_{1},a_{4})$ $(a_{3},a_{6})$ $(a_{3},a_{1})$ $(a_{4},a_{1})$ $(a_{6},a_{4})$ $(a_{6},a_{1})$
+9 $(a_{4},a_{6})$ $(a_{3},a_{5})$ $(a_{4},a_{5})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$
+8 $(a_{3},a_{5})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{5})$ $(a_{6},a_{3})$ $(a_{6},a_{4})$
+7 $(a_{2},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{3},a_{5})$ $(a_{2},a_{1})$ $(a_{2},a_{3})$
+6 $(a_{1},a_{3})$ $(a_{4},a_{1})$ $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{2},a_{4})$ $(a_{4},a_{3})$
+5 $(a_{5},a_{6})$ $(a_{4},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{1})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$
+4 $(a_{4},a_{5})$ $(a_{5},a_{1})$ $(a_{1},a_{5})$ $(a_{1},a_{6})$ $(a_{6},a_{2})$ $(a_{6},a_{2})$
+3 $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{4},a_{1})$ $(a_{2},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$
+2 $(a_{2},a_{3})$ $(a_{5},a_{6})$ $(a_{1},a_{6})$ $(a_{2},a_{3})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$
+1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
+An effectively homeo-monotonic intensity profile partitions the sets of agents and goods in certain clusters.
+These clusters can be of the envy-free type (by definition, at most one such cluster is possible) or of the top-preference-coincident type.
+Agents in the first cluster have completely heterogeneous preferences as far as their top items are concerned.
+Agents within each cluster of the second kind have identical preferences up to a certain rank-order position (say, the 3rd-ranked item for a cluster with 3 agents), whereas agents across different clusters have distinct such preferences.
+Within each top-preference-coincident cluster, moreover, agents can be weakly or strictly ordered in the way in which they rank their intensity of preference between any two items that appear consecutively in the relevant upper part of their preference ordering (monotonicity and strict monotonicity).
+Finally, this ordering over agents within any such cluster is required to be preserved over items that do not appear in consecutive positions in that cluster’s preference ordering (homeo-monotonicity).
+For example, if $s_{i}(a,b)\geq s_{i}(b,c)>0$ for all $i\leq 3$ and $s_{1}(a,b)\geq s_{2}(a,b)\geq s_{3}(a,b)$, $s_{1}(b,c)\geq s_{2}(b,c)\geq s_{3}(b,c)$ are also true, then homeo-monotonicity requires $s_{1}(a,c)\geq s_{2}(a,c)\geq s_{3}(a,c)$.
+Clearly, any set of agents can be partitioned into groups according to whether distinct or identical alternatives are the most preferred within each group.
+The additional structure imposed by effectively homeo-monotonic profiles amounts to requiring that: (i) the preferences of agents with the same top-ranked item also coincide in how they rank their second-, third-, $\ldots$, $k$th-most preferred item for some $k\leq n$ that coincides with the number of agents in that group; (ii) the preferences of distinct such groups are sufficiently heterogeneous.
+While this condition is restrictive and one can easily construct profiles where intensity-efficient allocations exist but the condition is violated, the proof of the next result shows that it is helpful toward clarifying some important mechanisms for the existence of such allocations in the general case.
+Theorem 2.
+Every effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation.
+Example 2 (continued).
+By Lemma 3, there are 12 Pareto-efficient allocations in this profile, namely those which assign $a_{1}$ to $v_{1}$ and alternate items $a_{2},a_{3},a_{4}$ among agents $v_{2},v_{3},v_{4}$ and $a_{5},a_{6}$ among $v_{5},v_{6}$.
+Out of these 12 allocations, only the following 2 are the (in fact, equivalent) intensity-efficient ones: $(a_{1},a_{2},a_{3},a_{4},a_{5},a_{6})$; $(a_{1},a_{2},a_{3},a_{4},a_{6},a_{5})$.
+$\blacklozenge$
+We note that the way in which the particular intensity-efficient allocation (labelled $x^{*}$) is constructed in the proof of Theorem 2 suggests the following stronger implication for a class of special cases:
+Corollary 2.
+An intensity-dominant (up to equivalence) allocation exists in any effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where every top-preference-coincident set has two agents.
+Evidently, the fine partitioning requirement of effective homeo-monotonicity is a restrictive sufficient condition for the existence of intensity-efficient allocations.
+Importantly, though, it is not necessary.
+Yet the intuitive structure that the homeo-monotonicity concept of Definition 9 imposes on a strictly monotonic top-preference-coincident set that may be present in some profile, even if that profile’s remaining structure rules out the possibility of partitioning agents and items in the way described in Definition 11, turns out to also be necessary for existence of intensity-efficient allocations.
+Theorem 3.
+If a profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation, then every strictly monotonic top-preference-coincident set in it is homeo-monotonic.
+The next example explains why “strictly monotonic” cannot be replaced by “monotonic” in the statement of Theorem 3.
+Example 3.
+For $n=4$ let the envy-free and top-preference-coincident sets that partition $N$ and $A$ be defined by $(N^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(N^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as in Table 4.
+The Pareto-efficient allocations are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$.
+Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$, $z$ (these are dominated by $p$, $y$, respectively).
+Yet $N^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$.
+$\blacklozenge$
+Table 4: A monotonic profile that is not homeo-monotonic but has an intensity-efficient allocation.
+$\overbrace{\hskip 40.0pt\text{$N^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 7.0pt\text{$N^{1}$\hskip 7.0pt}}$
+1 2 3 4
+$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}$ $a_{4}\succ_{4}\cdots\succ_{4}a_{1}$
+6 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{4},a_{1})$
+5 $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{1},a_{3})$
+4 $(a_{1},a_{3})$ $(a_{1},a_{3})$ $(a_{2},a_{4})$
+3 $(a_{1},a_{2})$ $(a_{1},a_{2})$ $(a_{3},a_{4})$
+2 $(a_{2},a_{3})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
+1 $(a_{3},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$
+Our last example, finally, shows that a strictly monotonic profile that is not homeo-monotonic may fail to have an intensity-efficient allocation.
+Example 4.
+For $n=5$ let the envy-free and top-preference-coincident sets that partition $N$ and $A$ be defined by $(N^{1},A^{1})=(\{v_{4},v_{5}\},\{a_{4},a_{5}\})$ and $(N^{2},A^{2})=(\{v_{1},v_{2},v_{3}\}$, $\{a_{1},a_{2},a_{3}\})$, as in Table 5.
+The Pareto-efficient allocations here are
+$\begin{array}[]{lllllll}\underbrace{(c,a,b,d,e)}_{o},&\underbrace{(c,b,a,d,e)}_{t},&\underbrace{(a,b,c,d,e)}_{x},&\underbrace{(a,c,b,d,e)}_{y},&\underbrace{(b,c,a,d,e)}_{w},&\underbrace{(b,a,c,d,e)}_{z}\end{array}$
+Yet the intensity-dominance relation $D$ is cyclic over this set:
+[…]
+$\blacklozenge$
+Table 5: A strictly monotonic but not homeo-monotonic profile that does not have an intensity-efficient allocation.
+$\overbrace{\hskip 40.0pt\text{$N^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$N^{1}$\hskip 20.0pt}}$
+1 2 3    4    5
+$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}\succ_{i}a_{5}$ $a_{4}\succ_{4}\cdots$ $|$ $a_{5}\succ_{5}\cdots$
+10 $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{4},\cdot)$ $(a_{5},\cdot)$
+9 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $\vdots$ $\vdots$
+8 $(a_{2},a_{5})$ $(a_{2},a_{5})$ $(a_{1},a_{3})$
+7 $(a_{2},a_{4})$ $(a_{1},a_{3})$ $(a_{2},a_{5})$
+6 $(a_{1},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$
+5 $(a_{3},a_{5})$ $(a_{3},a_{5})$ $(a_{3},a_{5})$
+4 $(a_{1},a_{2})$ $(a_{3},a_{4})$ $(a_{3},a_{4})$
+3 $(a_{2},a_{3})$ $(a_{1},a_{2})$ $(a_{4},a_{5})$
+2 $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
+1 $(a_{4},a_{5})$ $(a_{4},a_{5})$ $(a_{2},a_{3})$
+Together, Theorems 2-3 and Examples 2-4 highlight the importance of the intuitive structure imposed by homeo-monotonicity for the general existence of intensity-efficient allocations, both as a sufficient condition and, in a weaker sense, as a necessary one too.
+As the counterexample for $n=4$ in the proof of Theorem 1 shows, moreover, a profile may fail to have an intensity-efficient allocation even when homeo-monotonicity as defined above is satisfied, yet a subtle variation of it is not.
+More specifically, that profile features: (i) $n=4$; (ii) $a\succ_{i}b\succ_{i}c\succ_{i}d$ for $i\leq 4$ (hence a top-preference-coincident set $N$ with respect to $A$); (iii) $s_{1}=s_{4}$, $s_{2}=s_{3}$.
+Although the four agents’ intensities over their common preferences do not define a monotonic top-preference-coincident set, one observes that emergent here is a variant of this notion whereby each set in the partition $\{\{v_{1},v_{2}\},\{v_{3},v_{4}\}\}$ of $N$ is monotonically top-preference-coincident with respect to $\{a,b,c\}\subset A$: $s_{1}(a,b)>s_{2}(a,b)$, $s_{1}(b,c)=s_{2}(b,c)$ and $s_{4}(a,b)>s_{3}(a,b)$, $s_{4}(b,c)=s_{3}(b,c)$.
+However, both these pairs of agents violate what homeo-monotonicity would require here: $s_{1}(a,c)<s_{2}(a,c)$ and $s_{4}(a,c)<s_{3}(a,c)$.
+Thus, while the problem of fully characterizing the conditions under which strict profiles have intensity-efficient allocations remains open, the preceding analysis hopefully provides meaningful guidance in that direction.
+We end this section by remarking that the paper’s new analytical environment necessitates novel arguments in the proofs of this section’s main results.
+More specifically, viewed as an abstract mathematical problem, existence of intensity-efficient allocations amounts to identifying sufficient conditions for acyclicity of the intensity-dominance binary relation defined on the set of permutations over alternatives.
+The mathematical literature on the combinatorics of permutations has studied partial orders over such sets, which may even have the richer structure of a lattice that additionally ensures the existence of both a greatest (hence maximal) and a smallest element.
+These results, however, are not applicable in our environment.
+Indeed, the strong and weak Bruhat partial orders in question (Stanley, 2012, pp.
+399-400), relate two permutations if and only if there is a single transposition (“flip”) between elements in the two permutations, and this transposition corresponds to an inversion (Stanley, 2012, p.
+30) relative to some postulated fixed linear order over its elements.
+The weak Bruhat order (see Ceballos and Pons (2024) for a recent generalization) gives rise to a lattice, whose graph is known as the permut(a)(o)hedron, but under the additional assumption that the said single transposition occurs between adjacent elements of the two relevant permutations.
+Our problem, by contrast, does not assume a fixed linear order over items, and is defined by intensity-dominance comparisons that generally feature possibly multiple transpositions of non-adjacent elements.
+The latter, in turn, are defined by dominance between specific entries in the two relevant columns of an integer matrix, themselves dictated by the corresponding agents’ intensity orderings.
+5 Allocation via Aggregation of Intensity-Rank Scores
+5.1 A Generalization of Borda’s Criterion
+[…]
+A mapping $W:\mathcal{P}\rightarrow\mathcal{R}$ will be called a classic social welfare function, and a mapping $W:\mathcal{\widetilde{S}}\rightarrow\mathcal{R}$ an extended one.
+[^4] Harvey (1999), Weymark (2005) and Nebel (2024) study versions of the classical utilitarian social welfare function by also taking profiles of cardinal-utility representable intensity relations—rather than interpersonally comparable cardinal utility profiles—as the relevant domain.
+Using this terminology, and recalling the definition of $b_{l}$ in (10), the Borda count (Borda, 1781; Young, 1974; Mas-Colell et al., 1995; Maskin, 2025) is the classic social welfare function $\mathbb{Bo}:\mathcal{P}\rightarrow\mathcal{R}$ defined by
+$c\;\mathbb{Bo}(P)\;d$ $\Longleftrightarrow$ $\sum_{l=1}^{n}b_{l}^{P}(c)\geq\sum_{l=1}^{n}b_{l}^{P}(d)$ (15)
+Thus, when strict preference profiles act as inputs to the social welfare aggregation process, $a$ is socially preferred to $b$ under $\mathbb{Bo}$ if the aggregate Borda score of $a$ is larger than $b$.
+Definition 12.
+[…]
+[…]
+where $r_{l}$ is as defined in (11).
+[…]
+As was anticipated in Section 3, an alternative’s intensity-rank and Borda scores are generally distinct but subtly related concepts.
+Towards clarifying their connection further we introduce an additional notion.
+Definition 13.
+An intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widetilde${I}} is linear if, whenever
+$a_{\pi(1)}\succ_{l}a_{\pi(2)}\succ_{l}\cdots\succ_{l}a_{\pi(n)}$
+is true for the induced $\succ_{l}$ under some permutation $\pi$ on $N$, then
+$(a_{\pi(i)},a_{\pi(i+1)})$ $\dot{\sim}_{l}$ $(a_{\pi(i+1)},a_{\pi(i+2)})$ (17)
+holds for all $i=1,\ldots,n-2$.
+A profile $S\in\mathcal{\widetilde{S}}$ is linear if $\dot{\succsim}_{l}$ is linear for every $l\in N$.
+[…]
+The next result establishes that, on linear intensity profiles that induce strict preferences, the social ordering that is prescribed by the intensity-positional rule coincides with the social ordering prescribed by Borda on the corresponding induced preference profile.
+Theorem 4.
+[…]
+[…]
+5.2 Intensity-Positional Allocations
+The previous subsection clarified how intensity-positional social welfare generalizes Borda social welfare in the domain of intensity profiles that induce strict preferences and, in addition, feature the same number of intensity-equivalence classes for every agent.
+We now build on these insights and return our focus to the more restricted domain of strict (hence non-linear) intensity profiles, $\mathcal{\widehat{S}}\subset\mathcal{\widetilde{S}}$, in order to make the ensuing analysis of the assignment problem directly comparable to that of Section 4.
+The concept that we introduce next translates the intensity-positional scoring notation and formulae (16)(11) into the language of allocations.
+Definition 14.
+Let $S=\big(\dot{\succsim}_{1},\ldots,\dot{\succsim}_{n}\bigr)\in\widehat{\mathcal{S}}$ be a strict intensity profile, and let $s$ be its canonical representation.
+An allocation $x$ is intensity-positional at $S$ if
+$x$ $\in$ $\arg\max_{y\in\mathcal{A}}\sum_{l=1}^{n}r_{l}^{S}(y_{l})$
+$\equiv$ $\arg\max_{y\in\mathcal{A}}\sum_{l=1}^{n}\max_{y^{\prime}_{l}\in A}s_{l}^{S}(y_{l},y^{\prime}_{l})$
+$:=$ $\mathsf{IP}(S)$
+Analogously, a Borda allocation at profile $S\in\mathcal{\widehat{S}}$—henceforth denoted by $\mathsf{Bo}(S)\equiv\mathsf{Bo}(P_{S})$—is defined as in (14) but with the Borda scores $b_{l}^{P}$ at the strict preference profile $P_{S}$ replacing the intensity-rank scores $r_{l}^{S}$ at the strict intensity profile $S$.
+Example 5.
+Consider the example profile in Table 2 ($n=4$) for which we established that no intensity-efficient allocation exists.
+Since all agents’ preferences in this example are identical, all 24 (Pareto-efficient) allocations are also Borda allocations.
+As the reader can easily verify, however, only the following 8 such allocations are intensity-positional:
+$\begin{array}[]{llll}(a,c,d,b),&(a,d,c,b),&(b,a,c,d),&(b,c,a,d)\\ (b,c,d,a),&(b,d,c,a),&(d,a,c,b),&(d,c,a,b)\end{array}$
+$\blacklozenge$
+This example illustrates two important features of intensity-positional allocations: (i) they exist even when intensity-efficient allocations do not; (ii) they can refine the sets of Pareto-efficient and Borda-optimal allocations quite significantly (in this case, by discarding $2\over 3$} of them).
+This fact, in turn, naturally invites a formal investigation of general properties that such allocations have in the present analytical environment.
+Proposition 3.
+An intensity-positional allocation is a Borda allocation (hence Pareto-efficient): for all $S\in\mathcal{\widehat{S}}$, $\mathsf{IP}(S)\subseteq\mathsf{Bo}(S)$. The converse is false.
+In other words, Borda allocations—which, as is well-known, are Pareto-efficient—are refined by the intensity-positional criterion.
+That this refinement is generally strict can in turn be verified with the following $n=3$ example profile:
+$(a,c)\dot{\succ}_{1}(a,b)\dot{\succ}_{1}(b,c)$; $(a,c)\dot{\succ}_{2}(b,c)\dot{\succ}_{2}(a,b)$; $(c,a)\dot{\succ}_{3}(c,b)\dot{\succ}_{3}(b,a)$.
+The reader can readily check that both Pareto-efficient allocations $(a,b,c)$ and $(b,a,c)$ are selected at this profile under $\mathsf{Bo}$, whereas only the former is so under $\mathsf{IP}$ (and that this allocation is also intensity-efficient).
+This is intuitive: the first two agents have identical preferences, but the intensity rank of $b$ is higher for the second agent, reflecting the fact that preference intensity between $a$ and $b$ is higher for agent 1 than it is for agent 2.
+The intensity-positional criterion accounts for this interpersonal difference, whereas the Pareto and Borda criteria do not.
+In this example it so happened that the set of intensity-positional and intensity-efficient allocations coincided. Yet these criteria are distinct in general.
+Proposition 4.
+The sets of intensity-efficient and intensity-positional allocations at a profile $S\in\mathcal{\widehat{S}}$ are generally distinct.
+The first implication of this claim—namely, that an intensity-positional allocation may not be intensity-efficient—has already been established in Example 5.
+The converse implication can also be proved by modifying the profile of Example 3 (Table 4) so that $s_{2}(a_{1},a_{3})=5$, $s_{2}(a_{2},a_{4})=4$ and $s_{3}(a_{1},a_{2})=4$, $s_{3}(a_{2},a_{4})=3$, $s_{3}(a_{2},a_{3})=2$, $s_{3}(a_{3},a_{4})=1$.
+In this case, the set of intensity-efficient allocations contains three elements, one of which—namely, $(a_{3},a_{2},a_{1},a_{4})$—is not intensity-positional.
+This analysis uncovers that it is possible for allocations that are intensity-dominated to be intensity-positional.
+Intuitively, this happens in those situations where, despite the presence of a pairwise intensity dominance that favours $x$ over $y$, there are sufficiently many agents who do not belong to a “swapping” pair and who receive items under $y$ which they prefer more—according to the intensity-rank criterion—than those items’ beneficiaries under allocation $x$.
+In such cases, the welfare argument in favour of $x$ that is conveyed by the relevant pairwise intensity dominance comparisons—which are ordinal and ignore intensity-rank information—is offset by the welfare argument in favour of $y$ that are represented by the total differences in aggregate intensity rank scores.
+This also helps clarify how the two concepts that are studied in this and the preceding section jointly provide two distinct and complementary ways in which an analyst can utilise ordinal intensity information in order to improve upon the Pareto criterion.
+6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
+We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems.
+In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency, either in the intensity-dominating or intensity-positional direction.
+Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational.
+To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
+Proposition 5.
+Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in at most (n1)(n+1)(n-1)(n+1) steps, hence in $O(n^{2})$ time.
+Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available.
+The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm.
+The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run.
+TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment.
+Shapley and Scarf (1974) proved that TTC leads to a core allocation.
+Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique.
+Extending the results in Roth (1982), Ma (1994) established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
+Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes.
+This mechanism is not strategy-proof.
+Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism.
+The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available.
+These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority mechanism.
+Despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, Zhou (1990) showed that these are generally ex ante inefficient.
+Bogomolnaia and Moulin (2001), finally, demonstrated that they are also ordinally inefficient in the sense that they may be dominated by another random allocation under the stochastic-dominance relation that is defined by all agents’ preferences.
+In response to this realization, the latter authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
+Informed by the main insights of the above-cited literature, we now consider the following algorithm for the problem at hand:
+Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
+Step 1: Choose a linear order $\gg$ over $N$ from the uniform distribution with support $n!$.
+Step 2: Apply RP under $\gg$ on $P_{S}$ to arrive at a Pareto-efficient allocation $x$.
+Step 3: Reorder the entries in $x$ according to $\gg$, with $j>i$ for $x_{i},x_{j}$ if and only if $i\gg j$.
+Step 4: Search for the $\gg$-lowest agent $v_{j}\in N\setminus\{v_{1}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$.
+If such an agent exists, swap $x_{1}$ with $x_{j}$ and remove $v_{1},v_{j}$ from $N$ and $x_{1},x_{j}$ from $A$.
+If no such agent exists, remove $v_{1}$, $x_{1}$.
+Repeat Step 4 until no agents remain.
+The first two steps setup the RP algorithm.
+The third one merely relabels the items in the allocation that was arrived at earlier, so that the first (second, third, …) item belongs to the first-(second-, third-, …)ranked agent in the ordering that was drawn at the beginning.
+The fourth step is the essential new addition.
+It starts with agent 1 and searches—without loss, from the bottom up—for an agent $j$ whose RP assignment, $x_{j}$, is preferred by agent 1 to their own assignment, $x_{1}$, more than it is so by agent $j$.
+If such an agent exists, the algorithm implements the swap, removes the two agents and the respective items, then repeats this step with the remaining agents and items.
+If not, then agent 1 and $x_{1}$ are removed before the step is re-applied.
+This forced exit is important in two ways.
+First, it prevents potentially indefinite loops, for example in cases where the RP allocation is already intensity-efficient.
+Second, as is clarified below, it ensures that the computational complexity of the whole process falls in the manageable quadratic class.
+Proposition 6.
+If $S\in\mathcal{\widehat{S}}$, and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm ends in at most ${n}\choose{2}$} steps—hence in $O(n^{2})$ time—and either produces the Pareto-efficient allocation of its RP-implementing step or one that intensity-dominates it.
+The allocation delivered by this algorithm is not intensity-efficient in general.
+This is so because the forced exit of agents and items in the fourth step means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation.
+Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP.
+It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.555Suppose, instead, that an intensity-positional allocation is the welfare-improving criterion of interest.
+Such an allocation exists for any strict intensity profile and can be identified via simple summation once agents’ ordinal intensities are elicited.
+Owing to these facts, it is a straightforward implication of the discussion around Proposition 6 that, in order for the algorithm’s RP allocation to be converted into an intensity-positional one, it suffices to replace the last two steps by any deterministic or random criterion that selects one among the possibly many intensity-positional allocations corresponding to the given profile.
+Moreover, by Proposition 5 and the linear complexity of the RP algorithm (Manlove, 2013, Section 6.2), this process terminates in at most (n1)(n+1)(n-1)(n+1) as well.
+In light of the discussion preceding the statement of Proposition 6, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency.
+Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 5, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully.
+The answer is negative.
+To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$.
+The possible orderings over $N$ are: 1.
+$v_{1}\gg v_{2}\gg v_{3}$; 2.
+$v_{1}\gg v_{3}\gg v_{2}$; 3.
+$v_{2}\gg v_{1}\gg v_{3}$; 4.
+$v_{2}\gg v_{3}\gg v_{1}$; 5.
+$v_{3}\gg v_{1}\gg v_{2}$; 6.
+$v_{3}\gg v_{2}\gg v_{1}$.
+The two Pareto-efficient allocations are $(a,b,c)$, $(b,a,c)$, and the former intensity-dominates the latter.
+Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with prob.
+$1\over 2$}).
+If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it.
+But if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e.
+with prob.
+$1\over 2$}) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
+Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem.
+Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up.
+Abdulkadiroğlu et al.
+(2011), however, followed by Pycia (2011), Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for.
+In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools.
+In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible.
+In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.
+Appendix
+[…]
+[…]
+$r_{l}(a,b)$ $:=$ $\left\{\begin{array}[]{ll}|\{\mathbf{s}_{l}[a^{\prime},b^{\prime}]:s_{l}(a,b)>s_{l}(a^{\prime},b^{\prime})\geq s_{l}(e,e)\}|,&\text{if }s_{l}(a,b)>s_{l}(e,e)\\ s_{l}(a,b),&\text{if }s_{l}(a,b)=s_{l}(e,e)\\ -r_{l}(b,a),&\text{if }s_{l}(a,b)<s_{l}(e,e)\end{array}\right$.
+By construction, $r_{l}$ is canonical and $r_{l}(a,b)\geq r_{l}(c,d)\Leftrightarrow s_{l}(a,b)\geq s_{l}(c,d)$. $\square$
+Proof of Proposition 1.
+Consider the strict utility profile of Example 1.
+Observe that the Pareto-efficient allocations are $(a_{1},a_{2},a_{3})$ and $(a_{2},a_{1},a_{3})$, and the former is the utilitarian one.
+Yet this is intensity-dominated by the latter since $s_{2}(a_{1},a_{2})>s_{1}(a_{1},a_{2})$.
+$\square$
+Proof of Proposition 2.
 Under the postulated conditions, suppose to the contrary that $x(U)$ is a utilitarian but not intensity-efficient allocation.
@@
 It follows that $\sum_{l=1}^{n}u_{l}(x^{\prime}_{l})>\sum_{l=1}^{n}u_{l}\big(x_{l}(U)\bigr)$, which contradicts the postulated utilitarian-efficiency of $x(U)$.
-∎
-It is well-known that utilitarian allocations are Pareto-efficient.
-Proposition 2 shows that, when the utility profile is strict and balanced, this benchmark criterion has the additional normative status conveyed by intensity-efficiency.
-Furthermore, Proposition 2 implies the following existence result that is perhaps worth stating explicitly:
-Corollary 1.
-An intensity-efficient allocation exists for every intensity profile that is induced by a strict and balanced utility profile.
-4.3 Existence without Cardinality Assumptions
-We now return to the general environment in which intensity profiles come from the class of relations that were specified in Section 4.1, without imposing any further restrictions such as cardinal or even utility-difference representability of the individuals’ intensities.
-Theorem 1.
-An intensity-efficient allocation exists for every profile $S\in\mathcal{\widehat{S}}$ when $n=3$ but not in general when $n\geq 4$.
-Proof.
+$\square$
+Proof of Theorem 1.
 The argument proceeds by considering the possible ways in which an arbitrary profile $S\in\mathcal{\widehat{S}}$ when $n=3$ might generate a sequence of distinct Pareto-efficient allocations that are implicated in an intensity-dominance cycle.
@@
 We therefore have the following 4 cases to consider:
-[[Note: in what follows we make repeated use—often without explicit reference—of the Pareto-efficiency implications (15)–(17), the skew-symmetry property $s_{l}(a,b)=-s_{l}(b,a)$ of canonical representations and, whenever exact values of the $s_{i}$ functions are asserted, of the lateral-consistency property (6) together with the assumption that every $s_{i}$ is canonical and strict (cf Observations 34).]]
+[[Note: in what follows we make repeated use—often without explicit reference—of the Pareto-efficiency implications (22)(24), the skew-symmetry property $s_{l}(a,b)=-s_{l}(b,a)$ of canonical representations and, whenever exact values of the $s_{i}$ functions are asserted, of the lateral-consistency property (6) together with the assumption that every $s_{i}$ is canonical and strict (cf Observations 34).]]
 Case 1. $w^{3}=(b,c,a)$, $w^{4}=(c,b,a)$. By the definition of $D$, and by the above assumptions:
@@
 $s_{2}(a,c)>s_{3}(a,c)>s_{1}(a,c)$ (27)
-Since $w^{i}$ is Pareto efficient for $i\leq 4$, it follows from (15)–(17) that there are 4 subcases to consider:
+Since $w^{i}$ is Pareto efficient for $i\leq 4$, it follows from (22)(24) that there are 4 subcases to consider:
 Subcase 1-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{2}(b,c)>0$. By (6) and the fact that $s_{1},s_{2}$ are canonical, this implies $s_{1}(a,c)=3$, which contradicts (27).
@@
 $s_{1}(a,b)>s_{2}(a,b)>s_{3}(a,b)$ (30)
-In view of (15)–(17), we can now consider the following 4 possible subcases:
+In view of (22)(24), we can now consider the following 4 possible subcases:
 Subcase 2-i.
 $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$.
-By (23) and (15)–(17), the former postulate implies $s_{3}(a,b)>0$.
+By (30) and (22)(24), the former postulate implies $s_{3}(a,b)>0$.
 Since $s$ is canonical, this further implies $s_{3}(a,b)=1$, $s_{2}(a,b)=2$ and $s_{1}(a,b)=3$.
@@
 $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$.
-The latter, together with (15)–(17), (25) and canonicality, implies $s_{3}(a,b)=1$.
-But canonicality, (15)–(17) and (17) also implies $s_{3}(a,b)=3$, which contradicts Strictness.
+The latter, together with (22)(24), (32) and canonicality, implies $s_{3}(a,b)=1$.
+But canonicality, (22)(24) and (24) also implies $s_{3}(a,b)=3$, which contradicts Strictness.
 Subcase 3-ii.
@@
 We therefore have $s_{1}(b,a)=2$, $s_{1}(a,c)=1$ and, by (6) and canonicality, $s_{1}(b,c)=3$.
-From (15)–(17), (19), (25) and canonicality we also know that $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)>0$ implies $s_{3}(b,c)=2$ and $s_{2}(b,c)=1$.
-Thus, we have $s_{3}(b,a)>0$, $s_{3}(b,c)>0$ and, by (15)–(17) and $s_{1}(a,c)>0$, also $s_{3}(a,c)>0$.
+From (22)(24), (26), (32) and canonicality we also know that $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)>0$ implies $s_{3}(b,c)=2$ and $s_{2}(b,c)=1$.
+Thus, we have $s_{3}(b,a)>0$, $s_{3}(b,c)>0$ and, by (22)(24) and $s_{1}(a,c)>0$, also $s_{3}(a,c)>0$.
 But $s_{3}(b,a)>0$, $s_{3}(a,c)>0$ together with (6) and canonicality implies $s_{3}(b,c)=3$, a contradiction.
@@
 $w^{4}Dw^{1}$ $\Longrightarrow s_{3}(b,c)>s_{2}(b,c)$ (35)
-which, together with (18)–(19), imply
+which, together with (25)(26), imply
 $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)$ (36)
 Suppose first that $s_{2}(b,c)>0$.
-Then, by (29) and (15)–(17), $s_{1}(b,c)=3$, $s_{2}(b,c)=1$ and $s_{3}(b,c)=2$.
+Then, by (36) and (22)(24), $s_{1}(b,c)=3$, $s_{2}(b,c)=1$ and $s_{3}(b,c)=2$.
 From $s_{1}(b,c)=3$ and (6) we also get $s_{1}(b,a)>0$ and $s_{1}(a,c)>0$.
@@
 Suppose $s_{1}(a,c)>0$.
-From (27) and (15)–(17), $s_{2}(a,c)>0$.
+From (34) and (22)(24), $s_{2}(a,c)>0$.
 Since $s_{2}(a,c)>0$ and $s_{3}(c,b)>0$, by (6) we get $s_{2}(a,b)>s_{2}(c,b)=3$, which is impossible.
-Hence, $s_{1}(c,a)>0$ holds instead and, from (27) and (15)–(17), $s_{1}(c,a)>s_{2}(c,a)>0$ is also true.
+Hence, $s_{1}(c,a)>0$ holds instead and, from (34) and (22)(24), $s_{1}(c,a)>s_{2}(c,a)>0$ is also true.
 Suppose $s_{2}(a,b)>0$ holds too.
@@
 It has therefore been shown that $D$ is acyclic when $n=3$, establishing the existence of an intensity-efficient allocation in this case.
-For the claimed potential non-existence when $n\geq 4$, consider the example strict intensity profile on $A=\{a,b,c,d\}$ whose canonical representation is shown in the table below:
-$s_{i}(a,b)=$ $i=$ 1 2 3 4
-6 $(a,d)$ $(a,d)$ $(a,d)$ $(a,d)$
-5 $(b,d)$ $(a,c)$ $(a,c)$ $(b,d)$
-4 $(a,c)$ $(b,d)$ $(b,d)$ $(a,c)$
-3 $(b,c)$ $(b,c)$ $(b,c)$ $(b,c)$
-2 $(a,b)$ $(c,d)$ $(c,d)$ $(a,b)$
-1 $(c,d)$ $(a,b)$ $(a,b)$ $(c,d)$
-We start by observing that $s_{1}=s_{4}$, $s_{2}=s_{3}$ and $a\succ_{i}b\succ_{i}c\succ_{i}d$, $i=1,\ldots,4$.
-This implies that all 24 possible allocations, defined and listed below, are Pareto efficient.
-$\begin{array}[]{llllllllllllllllll}x^{1}&=&(a,b,c,d)&&&x^{2}&=&(a,b,d,c)&&&x^{3}&=&(a,c,b,d)&&&x^{4}&=&(a,c,d,b)\\ x^{5}&=&(a,d,b,c)&&&x^{6}&=&(a,d,c,b)&&&x^{7}&=&(b,a,c,d)&&&x^{8}&=&(b,a,d,c)\\ x^{9}&=&(b,c,a,d)&&&x^{10}&=&(b,c,d,a)&&&x^{11}&=&(b,d,a,c)&&&x^{12}&=&(b,d,c,a)\\ x^{13}&=&(c,a,b,d)&&&x^{14}&=&(c,a,d,b)&&&x^{15}&=&(c,b,a,d)&&&x^{16}&=&(c,b,d,a)\\ x^{17}&=&(c,d,a,b)&&&x^{18}&=&(c,d,b,a)&&&x^{19}&=&(d,a,b,c)&&&x^{20}&=&(d,a,c,b)\\ x^{21}&=&(d,b,a,c)&&&x^{22}&=&(d,b,c,a)&&&x^{23}&=&(d,c,a,b)&&&x^{24}&=&(d,c,b,a)\end{array}$
-The following comparisons, whose validity can be readily established by the reader, demonstrate that for each $x^{i}$, $i=1,\ldots,24$, in this set there is a distinct $x^{j}$ that intensity-dominates $x^{i}$—a situation denoted by $x^{j}Dx^{i}$.
-$\begin{array}[]{llllllll}x^{6}Dx^{1}&x^{1}Dx^{2}&x^{4}Dx^{3}&x^{13}Dx^{4}&x^{3}Dx^{5}&x^{17}Dx^{6}&x^{1}Dx^{7}&x^{2}Dx^{8}\\ x^{3}Dx^{9}&x^{8}Dx^{10}&x^{9}Dx^{11}&x^{11}Dx^{12}&x^{18}Dx^{13}&x^{13}Dx^{14}&x^{17}Dx^{15}&x^{22}Dx^{16}\\ x^{23}Dx^{17}&x^{24}Dx^{18}&x^{8}Dx^{19}&x^{22}Dx^{20}&x^{11}Dx^{23}&x^{12}Dx^{22}&x^{24}Dx^{23}&x^{10}Dx^{24}\end{array}$
-Therefore, no intensity-efficient allocation exists in this profile. ∎
-The somewhat surprising fact that a plausible dominance concept may by cyclic and prevent an optimal entity to emerge invites an informal analogy between intensity-dominance cycles over allocations with at least 4 agents and Condorcet cycles over alternatives in pairwise majority-based preference aggregation with at least 3 agents (Condorcet, 1785).
-Importantly, however, although intensity-dominance cycles here may prevent refining the Pareto set, unlike the Condorcet social-welfare theoretic framework these do not lead to a “policy paralysis” problem because the Pareto set is always non-empty and, absent any distributively juster suggestions, one of them might be promoted by the social planner.
-That said, it is naturally of interest to understand better the conditions under which our proposed refinement of Pareto-efficiency is well-defined.
-This is the focus of the next subsection.
-4.4 Existence Condition for General $n$: Homeo-Monotonicity
-For any set $A^{\prime}\subseteq A$ and preference relation $\succ_{i}$ that is induced by an intensity relation $\dot{\succsim}_{i}\in\mathcal{I}$ we write $\succ_{i}^{A^{\prime}}$ for the restriction of $\succ_{i}$ on $A^{\prime}$.
-Definition 8.
-Given a profile $S\in\mathcal{\widehat{S}}$, a set of agents $V^{\prime}\subseteq V$ is said to be:
-  1. 1.
-    Envy-free with respect to a set of items $A^{\prime}\subseteq A$ if each $v_{i}\in V^{\prime}$ has a distinct $\succ_{i}$-best item in $A$, and this item (denoted $b_{i}^{\prime}$) belongs to $A^{\prime}$.
-  2. 2.
-    Top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ if $\succ_{i}^{A^{\prime}}=\succ_{j}^{A^{\prime}}$ for all $v_{i},v_{j}\in V^{\prime}$ and, when $A^{\prime}\subset A$, $a^{\prime}\in A^{\prime}$, $a\in A\setminus A^{\prime}$ holds, then $a^{\prime}\succ_{i}a$ for all $v_{i}\in V^{\prime}$.
-An envy-free set $V^{\prime}$ with respect to $A^{\prime}$ consists of agents in $V^{\prime}$ who have distinct overall-best items in $A^{\prime}$.
-This terminology is motivated by the fact that, if $V^{\prime}\equiv V$ and $A^{\prime}\equiv A$, then each agent receiving their most preferred option results in an allocation that is Pareto-efficient and also envy-free (Varian, 1974; Moulin, 2004).
-A top-preference-coincident pair consists of agents whose preferences over $A^{\prime}$ are identical and also such that every item in $A^{\prime}$ is preferred to everything in $A\setminus A^{\prime}$, with no restrictions on preferences in the latter set.
-Definition 9.
-A set of agents $V^{\prime}\subseteq V$ that is top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ at profile $S\in\mathcal{\widehat{S}}$ is monotonic with respect to $A^{\prime}$ if $|V^{\prime}|=|A^{\prime}|$ and
-$s_{1}(a_{1},a_{2})$ >> $s_{1}(a_{2},a_{3})$ >> $\cdots$ >> $s_{1}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
-$\geq$ $\geq$ $\cdots$ $\geq$
-$s_{2}(a_{1},a_{2})$ >> $s_{2}(a_{2},a_{3})$ >> $\cdots$ >> $s_{2}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
-$\geq$ $\geq$ $\cdots$ $\geq$
-$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
-$\geq$ $\geq$ $\cdots$ $\geq$
-$s_{|V^{\prime}|}(a_{1},a_{2})$ >> $s_{|V^{\prime}|}(a_{2},a_{3})$ >> $\cdots$ >> $s_{|V^{\prime}|}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
-It is strictly monotonic if every weak inequality above is replaced by a strict inequality.
-Finally, a (strictly) monotonic such set is homeo-monotonic with respect to $A^{\prime}$ if
-$s_{i}(a_{i},a_{j})$ $\geq$ $s_{j}(a_{i},a_{j})$ for all $1\leq i<j\leq|V^{\prime}|$.
-Verbally, when $k$ agents have identical preferences over their top $k$ items, and agents can be (strictly) ordered in how much they prefer any two consecutively-ranked items in this subset [cf (strict) monotonicity], then this ordering cannot be violated in how much they prefer any non-consecutively ranked items in this subset (cf homeo-monotonicity).
-Definition 10.
-A profile $S\in\mathcal{\widehat{S}}$ is homeo-monotonic if, whenever $V$ and $A$ can be partitioned into subsets $V^{1},\ldots,V^{m}$ and $A^{1},\ldots,A^{m}$ where $V^{1}$ is envy-free with respect to $A^{1}$ and $V^{i>1}$ is monotonically top-preference-coincident with respect to $A^{i>1}$, then $V^{i>1}$ is homeo-monotonic with respect to $A^{i>1}$.
-Definition 11.
-A profile $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic if it induces partitions of agents and items relative to which it is homeo-monotonic.
-Example 2.
-The table below describes an effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where $n=6$ and where the envy-free and top-preference-coincident sets that partition $V$ and $A$ are $(V^{1},A^{1})=(\{v_{1}\},\{a_{1}\})$, $(V^{2},A^{2})=(\{v_{2},v_{3},v_{4}\},\{a_{2},a_{3},a_{4}\})$, $(V^{3},A^{3})=(\{v_{5},v_{6}\},\{a_{5},a_{6}\})$:
-$\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$ $\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{3}$}\hskip 20.0pt}$
-1 2 3 4 5 6
-$s_{i}(a,b)=$ $i=$ ${\color[rgb]{1,0,1}a_{1}}\succ_{1}\cdots$ ${\color[rgb]{1,0,0}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}}\succ_{i}\cdots$ ${\color[rgb]{0,0,1}a_{5}\succ_{i}a_{6}}\succ_{i}\cdots$
-15 $(a_{1},a_{6})$ $(a_{2},a_{1})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{5},a_{1})$ $(a_{5},a_{3})$
-14 $(a_{2},a_{6})$ $(a_{2},a_{6})$ $(a_{3},a_{5})$ $(a_{3},a_{6})$ $(a_{5},a_{4})$ $(a_{5},a_{1})$
-13 $(a_{1},a_{5})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{3})$ $(a_{5},a_{4})$
-12 $(a_{3},a_{6})$ $(a_{2},a_{4})$ $(a_{3},a_{6})$ $(a_{2},a_{1})$ $(a_{5},a_{2})$ $(a_{5},a_{2})$
-11 $(a_{2},a_{5})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$ $(a_{3},a_{1})$ $(a_{6},a_{1})$ $(a_{6},a_{3})$
-10 $(a_{1},a_{4})$ $(a_{3},a_{6})$ $(a_{3},a_{1})$ $(a_{4},a_{1})$ $(a_{6},a_{4})$ $(a_{6},a_{1})$
-9 $(a_{4},a_{6})$ $(a_{3},a_{5})$ $(a_{4},a_{5})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$
-8 $(a_{3},a_{5})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{6},a_{3})$ $(a_{6},a_{4})$
-7 $(a_{2},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$
-6 $(a_{1},a_{3})$ $(a_{4},a_{1})$ $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{2},a_{4})$ $(a_{4},a_{3})$
-5 $(a_{5},a_{6})$ $(a_{4},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{1})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$
-4 $(a_{4},a_{5})$ $(a_{5},a_{1})$ $(a_{1},a_{5})$ $(a_{1},a_{6})$ $(a_{6},a_{2})$ $(a_{6},a_{2})$
-3 $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{4},a_{1})$ $(a_{2},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$
-2 $(a_{2},a_{3})$ $(a_{5},a_{6})$ $(a_{1},a_{6})$ $(a_{2},a_{3})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$
-1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
-$\blacklozenge$
-Remark 1.
-If $V$ is envy-free with respect to $A$, then $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
-Clearly, any set of agents can be partitioned into groups according to whether distinct or identical alternatives are the most preferred within each group.
-The additional structure imposed by effectively homeo-monotonic profiles amounts to requiring that: (i) the preferences of agents with the same top-ranked item also coincide in how they rank their second-, third-, $\ldots$, $k$th-most preferred item for some $k\leq n$ that coincides with the number of agents in that group; (ii) the preferences of distinct such groups are sufficiently heterogeneous.
-While this condition is restrictive and one can easily construct profiles where intensity-efficient allocations exist but the condition is violated, the proof of the next result shows that it is helpful toward clarifying some important mechanisms for the existence of such allocations in the general case.
-Moreover, a weaker version of this condition that we consider later turns out to be necessary for such existence.
-Theorem 2.
-Every effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation.
-Proof.
+The potential non-existence when $n\geq 4$ is proved in the main text.
+(cf Table 2).
+$\square$
+Proof of Theorem 2.
 Consider a profile $S\in\mathcal{\widehat{S}}$ with the postulated structure.
-Let $V^{1}:=\{v^{1}_{1},\ldots,v^{1}_{|V^{1}|}\}$ be the envy-free set with respect to $A^{1}:=\{a^{1}_{1},\ldots,a^{1}_{|A^{1}|}\}$, and let $V^{i>1}:=\{i_{1},\ldots,i_{|V^{i}|}\}$ be the $i$th top-preference-coincident set with respect to $A^{i>1}:=\{i_{1},\ldots,i_{|A^{i}|}\}$, $i=2,\ldots,m$.
-By the profile assumption, $|A^{i}|=|V^{i}|$ for all $i\leq m$, and the collections $(V^{i})_{i=1}^{m}$ and $(A^{i})_{i=1}^{m}$ partition the sets $V$ and $A$, respectively.
-Without loss of generality we may assume that $a^{1}_{j}$ is the top-ranked item of agent $v^{1}_{j}$, $j=1,\ldots,|A^{1}|$ (for the general case we apply a permutation over $\{1,\ldots,|A^{1}|\}$ on the $a^{1}$-subscripts).
+Let $N^{1}:=\{N^{1}_{1},\ldots,N^{1}_{|N^{1}|}\}$ be the envy-free set with respect to $A^{1}:=\{a^{1}_{1},\ldots,a^{1}_{|A^{1}|}\}$, and let $N^{i>1}:=\{i_{1},\ldots,i_{|N^{i}|}\}$ be the $i$th top-preference-coincident set with respect to $A^{i>1}:=\{i_{1},\ldots,i_{|A^{i}|}\}$, $i=2,\ldots,m$.
+By the profile assumption, $|A^{i}|=|N^{i}|$ for all $i\leq m$, and the collections $(N^{i})_{i=1}^{m}$ and $(A^{i})_{i=1}^{m}$ partition the sets $N$ and $A$, respectively.
+Without loss of generality we may assume that $a^{1}_{j}$ is the top-ranked item of agent $N^{1}_{j}$, $j=1,\ldots,|A^{1}|$ (for the general case we apply a permutation over $\{1,\ldots,|A^{1}|\}$ on the $a^{1}$-subscripts).
 Recalling our earlier notation, this amounts to writing $a_{1}^{1}\equiv b_{1}^{1}$, $a_{2}^{1}\equiv b_{2}^{1}$, etc.
-Also without loss, for $A^{i>1}$, $V^{i>1}$ we may assume that $a^{i}_{1}\succ_{j}a^{i}_{2}\succ_{j}\cdots\succ_{j}a^{i}_{|A^{i}|}$ for all $v_{j}\in V^{i>1}$.
-(Note: the special cases where $V^{1}=\emptyset$, $V^{1}=V$ or $V^{i}=V$ for some $i>1$ are allowed by the postulated $S\in\mathcal{\widehat{S}}$, and so will they be in the proof.)
+Also without loss, for $A^{i>1}$, $N^{i>1}$ we may assume that $a^{i}_{1}\succ_{j}a^{i}_{2}\succ_{j}\cdots\succ_{j}a^{i}_{|A^{i}|}$ for all $v_{j}\in N^{i>1}$.
+(Note: the special cases where $N^{1}=\emptyset$, $N^{1}=N$ or $N^{i}=N$ for some $i>1$ are allowed by the postulated $S\in\mathcal{\widehat{S}}$, and so will they be in the proof.)
 Lemma 3.
 If $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic, then an allocation $x$ is Pareto-efficient with respect to $P_{S}$ if and only if it satisfies the following:
-  1. (i)
-    $v_{j}\in V^{1}$ $\Rightarrow$ $x_{j}=b_{j}^{1}$.
-  2. (ii)
-    $v_{j}\in V^{i>1}$ $\Rightarrow$ $x_{j}\in A^{i>1}$.
+(i) $v_{j}\in N^{1}$ $\Rightarrow$ $x_{j}=b_{j}^{1}$.
+(ii) $v_{j}\in N^{i>1}$ $\Rightarrow$ $x_{j}\in A^{i>1}$.
 Proof.
@@
 This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$.
-Let $V^{*}\subsetneq V$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$.
+Let $N^{*}\subsetneq N$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$.
 By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$.
-Suppose $V^{*}=\emptyset$.
+Suppose $N^{*}=\emptyset$.
 From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$.
-If the envy-free set $V^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in V^{1}$.
-If $V^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $V^{i>1}$ is non-empty.
-Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in V^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction.
-Now suppose $V^{*}\neq\emptyset$.
-Since $y$ Pareto-dominates $x$, the set $V^{**}:=V\setminus V^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty.
-From the preceding argument we know that for every $v_{l}\in V^{**}$ it must be that $v_{l}\not\in V^{1}$ and $v_{l}\neq v_{1}^{i}\in V^{i>1}$, for all $i>1$.
-By $s_{l}(y_{l},x_{l})>0$ and $|A|=|V|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$.
+If the envy-free set $N^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in N^{1}$.
+If $N^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $N^{i>1}$ is non-empty.
+Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in N^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction.
+Now suppose $N^{*}\neq\emptyset$.
+Since $y$ Pareto-dominates $x$, the set $N^{**}:=N\setminus N^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty.
+From the preceding argument we know that for every $v_{l}\in N^{**}$ it must be that $v_{l}\not\in N^{1}$ and $v_{l}\neq v_{1}^{i}\in N^{i>1}$, for all $i>1$.
+By $s_{l}(y_{l},x_{l})>0$ and $|A|=|N|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$.
 By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$.
-Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in V^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $V^{i>1}$ and $V^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$.
+Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in N^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $N^{i>1}$ and $N^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$.
 By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$.
@@
 Suppose to the contrary that (i) is false.
-That is, $x_{j}\neq b_{j}^{1}$ for some $v_{j}\in V^{1}$.
-Let $b_{j}^{1}=x_{l}$ for the relevant distinct agent $v_{l}\in V$ (such an agent exists because $|V|=|A|$).
-Either $v_{l}\in V^{1}$ or $v_{l}\in V^{i>1}$.
+That is, $x_{j}\neq b_{j}^{1}$ for some $v_{j}\in N^{1}$.
+Let $b_{j}^{1}=x_{l}$ for the relevant distinct agent $v_{l}\in N$ (such an agent exists because $|N|=|A|$).
+Either $v_{l}\in N^{1}$ or $v_{l}\in N^{i>1}$.
 In the former case, swapping $x_{j}$ and $x_{l}$ between the two agents Pareto-improves upon $x$, contradicting its postulated Pareto optimality.
-In the latter case, from the assumption that $V^{i>1}$ is top-preference-coincident with respect to $A^{i>1}$, and that $S$ is homeo-monotonic, we have $a^{i}_{t}\succ_{l}x_{l}\equiv b^{1}_{j}$ for every $a^{i}_{t}\in A^{i>1}$.
+In the latter case, from the assumption that $N^{i>1}$ is top-preference-coincident with respect to $A^{i>1}$, and that $S$ is homeo-monotonic, we have $a^{i}_{t}\succ_{l}x_{l}\equiv b^{1}_{j}$ for every $a^{i}_{t}\in A^{i>1}$.
 Furthermore, in this case (again because $|V|=|A|$) there must also exist another agent $k\in V^{p}$, $p\neq i$, such that $x_{k}\in A^{i>1}$.
@@
 If $p=1$, then the above contradiction is obtained now for $x_{l}$ and $x_{k}$.
-If $p>1$, then, by homeo-monotonicity of $S$ and the implied structure of $(V^{l>1},A^{l>1})$, $(V^{p>1},A^{p>1})$, we have $s_{l}(x_{k},x_{l})>0$ and $s_{k}(x_{l},x_{k})>0$, which also contradicts the postulated Pareto-efficiency of $x$.
+If $p>1$, then, by homeo-monotonicity of $S$ and the implied structure of $(N^{l>1},A^{l>1})$, $(N^{p>1},A^{p>1})$, we have $s_{l}(x_{k},x_{l})>0$ and $s_{k}(x_{l},x_{k})>0$, which also contradicts the postulated Pareto-efficiency of $x$.
 The argument establishing necessity of (ii) is analogous and omitted.
 ∎
-Next, consider a top-preference-coincident set $V^{i>1}$ with respect to $A^{i>1}$.
-By effective homeo-monotonicity, either such a non-empty set exists or the envy-free set $V^{1}$ coincides with $V$.
+Next, consider a top-preference-coincident set $N^{i>1}$ with respect to $A^{i>1}$.
+By effective homeo-monotonicity, either such a non-empty set exists or the envy-free set $N^{1}$ coincides with $N$.
 In the latter case, the allocation $x^{*}\in\mathcal{A}$ defined by $x^{*}_{i}=b^{1}_{i}$ is obviously the unique Pareto-efficient allocation at $S$ and therefore, trivially, intensity-efficient.
-So, in the sequel we assume that $V^{i>1}\neq\emptyset$.
-Denote by $a_{1}^{i}\succ_{i}a_{2}^{i}\succ_{i}\cdots\succ_{i}a_{|A^{i}|}^{i}$ the induced preference order on $A^{i>1}$ which, by the definition of $(V^{i>1},A^{i>1})$, is common across all agents in $V^{i>1}$.
-Define the binary relation $\trianglerighteq^{i}$ on $V^{i>1}$ by
+So, in the sequel we assume that $N^{i>1}\neq\emptyset$.
+Denote by $a_{1}^{i}\succ_{i}a_{2}^{i}\succ_{i}\cdots\succ_{i}a_{|A^{i}|}^{i}$ the induced preference order on $A^{i>1}$ which, by the definition of $(N^{i>1},A^{i>1})$, is common across all agents in $N^{i>1}$.
+Define the binary relation $\trianglerighteq^{i}$ on $N^{i>1}$ by
 $v_{j}\trianglerighteq^{i}v_{k}$ $\Longleftrightarrow$ $s_{j}(a^{i}_{l},a^{i}_{l+1})\geq s_{k}(a^{i}_{l},a^{i}_{l+1})\text{ for all }l=1,\ldots,|A^{i>1}|-1$
-By the postulated monotonicity of $(V^{i>1},A^{i>1})$, $\trianglerighteq^{i}$ is a weak order on $V^{i>1}$.
-Let $\trianglerighteq_{i}^{*}$ be some extension of $\trianglerighteq^{i}$ into a linear order over $V^{i>1}$, and denote by $\trianglerighteq^{*i}(j)$ the $\trianglerighteq^{*i}$-rank of agent $v_{j}\in V^{i>1}$.
+By the postulated monotonicity of $(N^{i>1},A^{i>1})$, $\trianglerighteq^{i}$ is a weak order on $N^{i>1}$.
+Let $\trianglerighteq_{i}^{*}$ be some extension of $\trianglerighteq^{i}$ into a linear order over $N^{i>1}$, and denote by $\trianglerighteq^{*i}(j)$ the $\trianglerighteq^{*i}$-rank of agent $v_{j}\in N^{i>1}$.
 By construction, $\trianglerighteq^{*i}$ coincides with $\trianglerighteq^{i}$ except where it breaks any equivalence ties that the latter relation may feature.
 Now define $x^{*}\in\mathcal{A}$ by
-$x_{j}^{*1}:=b_{j}^{1},\;j=1,\ldots,|V^{1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{1}$
-$x_{j}^{*i>1}:=a^{i}_{l},\;j=1,\ldots,|V^{i>1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in V^{i>1}\text{ and }\trianglerighteq^{*i}(j)=l$
-That is, each agent in the set $V^{1}$ (which could be empty) is assigned their most preferred item, while the $\trianglerighteq^{*i}$-$l$th-ranked agent in $V^{i>1}$ is assigned their $l$th most preferred item.
-Recalling that $V^{1}$ is envy-free and each $V^{i>1}$ is top-preference-coincident, it follows from Lemma 3 that $x^{*}$ is Pareto-efficient.
+$x_{j}^{*1}:=b_{j}^{1},\;j=1,\ldots,|N^{1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in N^{1}$
+$x_{j}^{*i>1}:=a^{i}_{l},\;j=1,\ldots,|N^{i>1}|\qquad$ $\Longleftrightarrow$ $\qquad v_{j}\in N^{i>1}\text{ and }\trianglerighteq^{*i}(j)=l$
+That is, each agent in the set $N^{1}$ (which could be empty) is assigned their most preferred item, while the $\trianglerighteq^{*i}$-$l$th-ranked agent in $N^{i>1}$ is assigned their $l$th most preferred item.
+Recalling that $N^{1}$ is envy-free and each $N^{i>1}$ is top-preference-coincident, it follows from Lemma 3 that $x^{*}$ is Pareto-efficient.
 We will prove that it is also intensity-efficient.
 To this end, suppose to the contrary that some Pareto-efficient allocation $y\in\mathcal{A}$ intensity-dominates $x^{*}$.
-Then, there are agents $v_{i},v_{j}\in V$ such that $(x_{i}^{*},x_{j}^{*})=(y_{j},y_{i})$ and $s_{j}(y_{j},y_{i})\equiv s_{j}(x_{i}^{*},x_{j}^{*})>s_{i}(x^{*}_{i},x_{j}^{*})>0$.
-By construction of $x^{*}$, $v_{i},v_{j}$ cannot belong to the same $V^{l\geq 1}$.
+Then, there are agents $v_{i},v_{j}\in N$ such that $(x_{i}^{*},x_{j}^{*})=(y_{j},y_{i})$ and $s_{j}(y_{j},y_{i})\equiv s_{j}(x_{i}^{*},x_{j}^{*})>s_{i}(x^{*}_{i},x_{j}^{*})>0$.
+By construction of $x^{*}$, $v_{i},v_{j}$ cannot belong to the same $N^{l\geq 1}$.
 For $l=1$ this is obvious because $x^{*i}=b^{1}_{i}$ and $x^{*j}=b^{1}_{j}$, contradicting the above postulate.
-For $l>1$, moreover, one has $x^{*}_{i},x_{j}^{*}\in A^{l>1}$ while items in $A^{l>1}$ are assigned to agents in $V^{l>1}$ according to the linear order $\trianglerighteq^{*l}$, which, by homeo-monotonicity of $S$, ensures that $s_{i}(x_{i}^{*},x_{j}^{*})\geq s_{j}(x_{i}^{*},x_{j}^{*})$, thereby contradicting that postulate.
-The remaining possibility is $v_{i}\in V^{l>1}$ and $v_{j}\in V^{k>1}$ for some $k\neq l$.
-By construction of $x^{*}$, however, we have $x_{i}^{*}\in A^{l>1}$ and $x_{j}^{*}\in A^{k>1}$ which, because $V^{l>1}$, $V^{k>1}$ are top-preference-coincident with respect to $A^{l>1}$, $A^{k>1}$ and $A^{l>1}\cap A^{k>1}=\emptyset$, implies $s_{i}(x_{i}^{*},x_{j}^{*})\equiv s_{i}(y_{j},y_{i})>0$ and $s_{j}(x_{i}^{*},x_{j}^{*})\equiv s_{j}(y_{j},y_{i})<0$.
+For $l>1$, moreover, one has $x^{*}_{i},x_{j}^{*}\in A^{l>1}$ while items in $A^{l>1}$ are assigned to agents in $N^{l>1}$ according to the linear order $\trianglerighteq^{*l}$, which, by homeo-monotonicity of $S$, ensures that $s_{i}(x_{i}^{*},x_{j}^{*})\geq s_{j}(x_{i}^{*},x_{j}^{*})$, thereby contradicting that postulate.
+The remaining possibility is $v_{i}\in N^{l>1}$ and $v_{j}\in N^{k>1}$ for some $k\neq l$.
+By construction of $x^{*}$, however, we have $x_{i}^{*}\in A^{l>1}$ and $x_{j}^{*}\in A^{k>1}$ which, because $N^{l>1}$, $N^{k>1}$ are top-preference-coincident with respect to $A^{l>1}$, $A^{k>1}$ and $A^{l>1}\cap A^{k>1}=\emptyset$, implies $s_{i}(x_{i}^{*},x_{j}^{*})\equiv s_{i}(y_{j},y_{i})>0$ and $s_{j}(x_{i}^{*},x_{j}^{*})\equiv s_{j}(y_{j},y_{i})<0$.
 Thus, the postulate that $y$ intensity-dominates $x^{*}$ is contradicted in all situations.
-∎
-Example 2 (continued).
-By Lemma 3, there are 12 Pareto-efficient allocations in this profile, namely those which assign $a_{1}$ to $v_{1}$ and alternate items $a_{2},a_{3},a_{4}$ among agents $v_{2},v_{3},v_{4}$ and $a_{5},a_{6}$ among $v_{5},v_{6}$.
-Out of these 12 allocations, only the following 2 are the (in fact, equivalent) intensity-efficient ones: $(a_{1},a_{2},a_{3},a_{4},a_{5},a_{6})$; $(a_{1},a_{2},a_{3},a_{4},a_{6},a_{5})$.
-$\blacklozenge$
-The way in which the intensity-efficient allocation $x^{*}$ is constructed in the proof of Theorem 2 readily suggests the following stronger implication for a class of special cases:
-Corollary 2.
-An intensity-dominant—up to equivalence—allocation exists in any effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where every top-preference-coincident set has two agents.
-Importantly, although effective homeo-monotonicity is a sufficient but not necessary condition for the existence of intensity-efficient allocations, we show next that the restriction it imposes within monotonic top-preference-coincident sets—if the profile induces such sets—is also necessary whenever these sets are strictly monotonic.
-Theorem 3.
-If a profile $S\in\mathcal{\widehat{S}}$ has an intensity-efficient allocation, then every strictly monotonic top-preference-coincident set in it is homeo-monotonic.
-Proof.
+$\square$
 […]
@@
 Proof.
-The only difference between allocations $x$ and $x^{\prime}$ lies in how they permute items in $A^{\prime}$ to agents in $V^{\prime}$.
+The only difference between allocations $x$ and $x^{\prime}$ lies in how they permute items in $A^{\prime}$ to agents in $N^{\prime}$.
 Denote their restrictions on these subsets by $\widehat{x}$ and $\widehat{x}^{\prime}$, respectively.
@@
 Since $x$ is intensity-efficient, it is not intensity-dominated by $x^{\prime}$.
-Hence, there is no pair $(a^{\prime},a^{\prime\prime})\in A^{\prime}\times A^{\prime}$ such that $(a^{\prime},a^{\prime\prime})=(\widehat{x}_{i},\widehat{x}_{j})=(\widehat{x}^{\prime}_{j},\widehat{x}^{\prime}_{i})$ for some $v^{\prime}_{i},v^{\prime}_{j}\in V^{\prime}$, making $x$ and $x^{\prime}$ incomparable by intensity-dominance.
+Hence, there is no pair $(a^{\prime},a^{\prime\prime})\in A^{\prime}\times A^{\prime}$ such that $(a^{\prime},a^{\prime\prime})=(\widehat{x}_{i},\widehat{x}_{j})=(\widehat{x}^{\prime}_{j},\widehat{x}^{\prime}_{i})$ for some $N^{\prime}_{i},N^{\prime}_{j}\in N^{\prime}$, making $x$ and $x^{\prime}$ incomparable by intensity-dominance.
 The argument in the proof of Theorem 2 following Lemma 3 establishes that there is no other Pareto-efficient allocation that intensity-dominates $x^{\prime}$ either.
@@
 […]
-$\begin{array}[]{lllllllll}s^{\prime}_{1}(a^{\prime}_{i},a^{\prime}_{i+1})&>&s^{\prime}_{2}(a^{\prime}_{i},a^{\prime}_{i+1})&>&\cdots&>&s^{\prime}_{|V^{\prime}|}(a^{\prime}_{i},a^{\prime}_{i+1})&&\text{for all }i\leq|A^{\prime}|-1\end{array}$
+$\begin{array}[]{lllllllll}s^{\prime}_{1}(a^{\prime}_{i},a^{\prime}_{i+1})&>&s^{\prime}_{2}(a^{\prime}_{i},a^{\prime}_{i+1})&>&\cdots&>&s^{\prime}_{|N^{\prime}|}(a^{\prime}_{i},a^{\prime}_{i+1})&&\text{for all }i\leq|A^{\prime}|-1\end{array}$
 Now recall that the allocation $x^{\prime}$ constructed above is intensity-efficient.
-This assigns $x^{\prime}_{m}=a^{\prime}_{m}\in A^{\prime}$ to each $v^{\prime}_{m}\in V^{\prime}$.
-Since $(V^{\prime},A^{\prime})$ is strictly monotonic but not homeo-monotonic, there are $v^{\prime}_{j},v^{\prime}_{l}\in V^{\prime}$ such that $v^{\prime}_{j}\trianglerighteq^{\prime}v^{\prime}_{l}$, jl>1j-l>1 and $s^{\prime}_{l}(a^{\prime}_{j},a^{\prime}_{l})>s^{\prime}_{j}(a^{\prime}_{j},a^{\prime}_{l})$.
+This assigns $x^{\prime}_{m}=a^{\prime}_{m}\in A^{\prime}$ to each $N^{\prime}_{m}\in N^{\prime}$.
+Since $(N^{\prime},A^{\prime})$ is strictly monotonic but not homeo-monotonic, there are $N^{\prime}_{j},N^{\prime}_{l}\in N^{\prime}$ such that $N^{\prime}_{j}\trianglerighteq^{\prime}N^{\prime}_{l}$, jl>1j-l>1 and $s^{\prime}_{l}(a^{\prime}_{j},a^{\prime}_{l})>s^{\prime}_{j}(a^{\prime}_{j},a^{\prime}_{l})$.
 Since $(x^{\prime}_{j},x^{\prime}_{l})=(a^{\prime}_{j},a^{\prime}_{l})$, this implies that the allocation $x^{\prime\prime}$ which is identical to $x^{\prime}$ except that $(x^{\prime\prime}_{j},x^{\prime\prime}_{l})=(a^{\prime}_{l},a^{\prime}_{j})$ intensity-dominates $x^{\prime}$, contradicting the postulated intensity-efficiency of the latter allocation.
+$\square$
+Proof of Theorem 4.
+To ease notation, we write $r_{l}$ instead of $r_{l}^{S}$ in much of the proof.
+We start with the following lemma, which is valid regardless of whether the intensity relation $\dot{\succsim}_{l}$ is linear or whether it induces a strict or weak preference relation on $A$:
+Lemma 5.
+[…]
+Proof.
+By definition,
+$\begin{array}[]{ccccc}r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})&\geq&r_{l}(a_{j})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{j},a_{k})&\Longleftrightarrow\\ s_{l}(a_{i},a_{k})&\geq&s_{l}(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow\\ (a_{i},a_{k})&\dot{\succsim}_{l}&(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow&\\ (a_{i},a_{j})&\dot{\succsim}_{l}&(a_{j},a_{i})&\Longleftrightarrow&a_{i}\succsim_{l}a_{j},\end{array}$
+where the penultimate equivalence immediately follows from $(\mathbf{4_{\mathit{D}}})$, which in turn is implied by the postulate that $s_{l}$ represents $\dot{\succsim}_{l}$ in profile $S\in\mathcal{\widetilde{S}}$.
 ∎
-The next example explains why “strictly monotonic” cannot be replaced by “monotonic” in the statement of Theorem 3.
-Example 3.
-For $n=4$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4}\},\{a_{4}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\},\{a_{1},a_{2},a_{3}\})$, as follows:
-$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$
-1 2 3 4
-$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}$ $a_{4}\succ_{4}\cdots\succ_{4}a_{1}$
-6 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{4},a_{1})$
-5 $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{1},a_{3})$
-4 $(a_{1},a_{3})$ $(a_{1},a_{3})$ $(a_{2},a_{4})$
-3 $(a_{1},a_{2})$ $(a_{1},a_{2})$ $(a_{3},a_{4})$
-2 $(a_{2},a_{3})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
-1 $(a_{3},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$
-The Pareto-efficient allocations here are $p:=(a_{1},a_{2},a_{3},a_{4})$, $q:=(a_{1},a_{3},a_{2},a_{4})$, $r:=(a_{2},a_{3},a_{1},a_{4})$, $x:=(a_{2},a_{1},a_{3},a_{4})$, $y:=(a_{3},a_{1},a_{2},a_{4})$ and $z:=(a_{3},a_{2},a_{1},a_{4})$.
-Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$ and $z$ (these are dominated by $p$ and $y$, respectively).
-Yet $V^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$.
-$\blacklozenge$
-Our last example, finally, shows that a strictly monotonic profile that is not homeo-monotonic may fail to have an intensity-efficient allocation.
-Example 4.
-For $n=5$ let the envy-free and top-preference-coincident sets that partition $V$ and $A$ be defined by $(V^{1},A^{1})=(\{v_{4},v_{5}\},\{a_{4},a_{5}\})$ and $(V^{2},A^{2})=(\{v_{1},v_{2},v_{3}\}$, $\{a_{1},a_{2},a_{3}\})$, as follows:
-$\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{1}$\hskip 20.0pt}}$
-1 2 3    4    5
-$s_{i}(a,b)=$ $i=$ $a_{1}\succ_{i}a_{2}\succ_{i}a_{3}\succ_{i}a_{4}\succ_{i}a_{5}$ $a_{4}\succ_{4}\cdots$ $|$ $a_{5}\succ_{5}\cdots$
-10 $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{1},a_{5})$ $(a_{4},\cdot)$ $(a_{5},\cdot)$
-9 $(a_{1},a_{4})$ $(a_{1},a_{4})$ $(a_{1},a_{4})$ $\vdots$ $\vdots$
-8 $(a_{2},a_{5})$ $(a_{2},a_{5})$ $(a_{1},a_{3})$
-7 $(a_{2},a_{4})$ $(a_{1},a_{3})$ $(a_{2},a_{5})$
-6 $(a_{1},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$
-5 $(a_{3},a_{5})$ $(a_{3},a_{5})$ $(a_{3},a_{5})$
-4 $(a_{1},a_{2})$ $(a_{3},a_{4})$ $(a_{3},a_{4})$
-3 $(a_{2},a_{3})$ $(a_{1},a_{2})$ $(a_{4},a_{5})$
-2 $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{1},a_{2})$
-1 $(a_{4},a_{5})$ $(a_{4},a_{5})$ $(a_{2},a_{3})$
-The Pareto-efficient allocations here are
-$\begin{array}[]{lllllll}\underbrace{(c,a,b,d,e)}_{o},&\underbrace{(c,b,a,d,e)}_{t},&\underbrace{(a,b,c,d,e)}_{x},&\underbrace{(a,c,b,d,e)}_{y},&\underbrace{(b,c,a,d,e)}_{w},&\underbrace{(b,a,c,d,e)}_{z}\end{array}$
-Yet the intensity-dominance relation $D$ is cyclic over this set:
+Considering now the stronger postulate whereby $\dot{\succsim}_{l}\in\mathcal$$\widetilde${I}} is linear, note first that this implies that the relation $\succsim_{l}$ induced by $\dot{\succsim}_{l}$ is a linear order.
+To ease notation, in the rest of the proof we assume without loss that the $n$ elements of $A$ are $\succ_{l}$-ranked by every agent $l\in N$ as follows:
+$\begin{array}[]{lllllllllll}a_{1}&\succ_{l}&a_{2}&\succ_{l}&a_{3}&\succ_{l}&\cdots&\succ_{l}&a_{n-1}&\succ_{l}&a_{n}\end{array}$
+(Dealing with heterogeneous preferences in $P_{S}$ is straightforward and involves replacing the index $i$ in $a_{i}$ with $\pi^{l}(i)$ for a suitable agent-specific permutation $\pi^{l}$ on $\{1,\ldots,n\}$.)
+Lemma 6.
 […]
-$\blacklozenge$
-Together, Theorems 2-3 and Examples 2-4 highlight the importance of the intuitive structure imposed by homeo-monotonicity for the general existence of intensity-efficient allocations, both as a sufficient condition and, in a weaker sense, as a necessary one too.
-As the counterexample for $n=4$ in the proof of Theorem 1 shows, moreover, a profile may fail to have an intensity-efficient allocation even when homeo-monotonicity as defined above is satisfied, yet a subtle variation of it is not.
-More specifically, that profile features: (i) $n=4$; (ii) $a\succ_{i}b\succ_{i}c\succ_{i}d$ for $i\leq 4$ (hence a top-preference-coincident set $V$ with respect to $A$); (iii) $s_{1}=s_{4}$, $s_{2}=s_{3}$.
-Although the four agents’ intensities over their common preferences do not define a monotonic top-preference-coincident set, one observes that emergent here is a variant of this notion whereby each set in the partition $\{\{v_{1},v_{2}\},\{v_{3},v_{4}\}\}$ of $V$ is monotonically top-preference-coincident with respect to $\{a,b,c\}\subset A$: $s_{1}(a,b)>s_{2}(a,b)$, $s_{1}(b,c)=s_{2}(b,c)$ and $s_{4}(a,b)>s_{3}(a,b)$, $s_{4}(b,c)=s_{3}(b,c)$.
-However, both these pairs of agents violate what homeo-monotonicity would require here: $s_{1}(a,c)<s_{2}(a,c)$ and $s_{4}(a,c)<s_{3}(a,c)$.
-Thus, while the problem of fully characterizing the conditions under which strict profiles have intensity-efficient allocations remains open, the preceding analysis hopefully provides non-trivial and useful guidance in that direction.
-Figure 1: Simulations-based histograms showing the proportions of Pareto-efficient allocations that are also intensity-efficient when the number of agents and items is small.
-Additional motivation for further study is also contributed by the analysis of uniform-randomly drawn intensity profiles when $n=3,4,5$.
-Figure 1 summarizes this output and displays how the distribution of the per-profile proportion of Pareto-efficient allocations that are also intensity-efficient varies with $n$.
-Interestingly, while the average number of Pareto-efficient allocations increases rapidly from 2.8 to 5.9 to 14.1 when $n=3,4,5$, respectively, the corresponding averages for intensity-efficient allocations are 1.6, 1.9 and 2.4.
-The refinement gains afforded by intensity-efficiency are therefore significant and appear to be increasing in $n$, with the mean proportion of intensity-efficient out of all Pareto-efficient allocations falling from 0.63 ($n=3$) to 0.2 ($n=5$).
-Finally, non-existence of intensity-efficient allocations is a rare occurrence in these simulations, with only 1 and 15 out of the respective 100,000 randomly drawn profiles being associated with a cyclic intensity-dominance relation over Pareto-efficient allocations when $n=4$ and $n=5$, respectively.
-We end this section by remarking that the paper’s new analytical environment necessitated novel arguments in the proofs of this section’s main results.
-More specifically, viewed as an abstract mathematical problem, existence of intensity-efficient allocations amounts to identifying sufficient conditions for acyclicity of the intensity-dominance binary relation defined on the set of permutations over alternatives.
-The mathematical literature on the combinatorics of permutations has studied partial orders over such sets, which may even have the richer structure of a lattice that additionally ensures the existence of both a greatest (hence maximal) and a smallest element.
-These results, however, are not applicable in our environment.
-The reason is that those partial orders, namely the strong and weak Bruhat orders (Stanley, 2012, pp.
-399-400), relate two permutations if and only if there is a single transposition (“flip”) between elements in the two permutations, and this transposition corresponds to an inversion (Stanley, 2012, p.
-30) relative to some postulated fixed linear order over its elements.
-The weak Bruhat order (see Ceballos and Pons (2024) for a recent generalization) gives rise to a lattice—whose graph is known as the permut(a)(o)hedron—but under the additional assumption that the said single transposition occurs between adjacent elements of the two relevant permutations.
-Our problem on the other hand does not assume a fixed linear order over items and is defined by intensity-dominance comparisons that generally feature possibly multiple transpositions of non-adjacent elements, which are defined by dominance between specific entries in the two relevant columns of an integer matrix, defined in turn by the corresponding agents’ intensity orderings.
-5 Eliciting Ordinal Intensities and Finding Intensity-Dominating Allocations
-We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems.
-In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency in the intensity-dominating direction.
-Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational.
-To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
-Proposition 3.
-Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in exactly $\frac{1}{2}(n-1)(n-2)$ steps, hence in $O(n^{2})$ time.
+$r_{l}(A)$ $=$ $\{0,1,\ldots,n-1\}$ (38)
 Proof.
-Define $f(n):=\frac{1}{2}(n-1)(n-2)$. Observe that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains
-$|\dot{\succsim}_{l}|$ $=$ $2{n\choose 2}+n$ (31)
-comparisons between pairs in $A\times A$. We proceed by induction.
-Let $n=3$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$.
-By (31), $|\dot{\succsim}_{l}|=9$.
-By (6) and the assumption that $\succ_{l}$ is induced by $\dot{\succsim}_{l}$ we readily get $(a_{1},a_{3})\dot{\succsim}_{l}(a_{1},a_{2})$ and $(a_{1},a_{3})\dot{\succsim}_{l}(a_{2},a_{3})$.
-By (5), moreover, we have $(a_{1},a_{1})\dot{\sim}_{l}(a_{2},a_{2})\dot{\sim}_{l}(a_{3},a_{3})$.
-By (5), $(\mathbf{5}_{D})$ and $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}$ we also get $(a_{1},a_{2})\dot{\succ}_{l}(a_{i},a_{i})$ and $(a_{2},a_{3})\dot{\succ}_{l}(a_{i},a_{i})$, $i\leq 3$.
-By (5) and the above comparisons, finally, we also have $(a_{i},a_{i})\dot{\succ}_{l}(a_{2},a_{1})\dot{\succ}_{l}(a_{3},a_{1})$ and $(a_{i},a_{i})\dot{\succ}_{l}(a_{3},a_{2})\dot{\succ}_{l}(a_{3},a_{1})$.
-Together, these account for 7 out of the 9 comparisons in $\dot{\succsim}_{l}$, with those between $(a_{1},a_{2})$, $(a_{2},a_{3})$ and between $(a_{3},a_{2})$, $(a_{2},a_{1})$ being the two missing ones.
-By (5) once again, knowing either the first or the second such comparison also pins down the other.
-Hence, $f(3)=1$ is indeed the number of steps needed to elicit $\dot{\succsim}_{l}$ from $\succ_{l}$ when $n=3$.
-Now let $n=4$ and, without loss, $a_{1}\succ_{l}a_{2}\succ_{l}a_{3}\succ_{l}a_{4}$.
-By the preceding arguments we know that, out of the $|\dot{\succsim}_{l}|=16$ total comparisons here, missing are the following 3: $(a_{1},a_{2})$ vs $(a_{2},a_{3})$; $(a_{2},a_{3})$ vs $(a_{3},a_{4})$; $(a_{1},a_{3})$ vs $(a_{2},a_{4})$.
-Hence, $f(4)=3$ is the number of necessary steps here too.
-For the induction step of the proof when $n>4$, suppose $f(m)$ is the correct number for every $n\leq m$.
-We will show that $f(m+1)$ is the correct number when $n=m+1$.
-To this end, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m}$ and notice that, from the induction hypothesis and the information derived from the $n=3$ and $n=4$ cases above, we can decompose the $f(m)$ comparisons as follows:
-(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-2$ (m2m-2 comparisons);
-(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-3$ (m3m-3 comparisons);
-(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-4$ (m4m-4 comparisons);
-$\cdots$
-(m2m-2):  $(a_{i},a_{i+m-1})$ vs $(a_{i+1},a_{i+m})$ (1 comparison).
-Indeed, summing the terms inside brackets across the right hand sides of $(1)\cdots(m-2)$ we get
-$\sum\limits_{i=2}^{m-1}(m-i)$ $=$ $\sum\limits_{i=1}^{m-1}(m-i)-(m-1)$
-$=$ $\frac{1}{2}(m-1)m-(m-1)$
-$=$ $(m-1)\left(\frac{1}{2}m-1\right)$
-$=$ $\frac{1}{2}(m-1)(m-2)$
-$=$ $f(m)$
-Consider now $n=m+1$ and, once again without loss, let $a_{1}\succ_{l}a_{2}\succ_{l}\ldots\succ_{l}a_{m+1}$.
-From the preceding decomposition of the required comparisons when $n=m$ we can deduce that, for $n=m+1$,
-(1):  $(a_{i},a_{i+1})$ vs $(a_{i+1},a_{i+2})$ for $i\leq m-1$ (m1m-1 comparisons);
-(2):  $(a_{i},a_{i+2})$ vs $(a_{i+1},a_{i+3})$ for $i\leq m-2$ (m2m-2 comparisons);
-(3):  $(a_{i},a_{i+3})$ vs $(a_{i+1},a_{i+4})$ for $i\leq m-3$ (m3m-3 comparisons);
-$\cdots$
-(m1m-1):  $(a_{i},a_{i+m})$ vs $(a_{i+1},a_{i+m+1})$ (1 comparison).
-Summing across categories now yields $\sum\limits_{i=1}^{m}(m-i)=\frac{1}{2}(m-1)m=f(m+1)$. ∎
-Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available.
-The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm.
-The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run.
-TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment.
-Shapley and Scarf (1974) proved that TTC leads to a core allocation.
-Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique.
-Extending the results in Roth (1982), Ma (1994) later established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
-Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes.
-This mechanism is not strategy-proof.
-Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism.
-The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available.
-These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority under some ranking over agents.
-Bogomolnaia and Moulin (2001), finally, demonstrated that, despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, these are generally ex ante inefficient in the sense that, based on the stochastic-dominance relation defined by all agents’ preferences, they may be dominated by another random allocation.
-In response to this realization, the authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
-Informed by the main insights of the above-cited literature, we now consider the following algorithm for the problem at hand:
-Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
-Step 1: Choose a linear order $\gg$ over $V$ from the uniform distribution with support $n!$.
-Step 2: Apply RP under $\gg$ on $P_{S}$ to arrive at a Pareto-efficient allocation $x$.
-Step 3: Reorder the entries in $x$ according to $\gg$, with $j>i$ for $x_{i},x_{j}$ if and only if $i\gg j$.
-Step 4: Search for the $\gg$-lowest agent $v_{j}\in\{v_{2},\ldots,v_{n}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$.
-If such an agent exists, swap $x_{1}\leftrightarrow x_{j}$ and remove $v_{1},v_{j}$ from $V$ and $x_{1},x_{j}$ from $A$.
-If no such agent exists, remove $v_{1}$, $x_{1}$.
-Repeat Step 4 until no agents remain.
-Proposition 4.
-If $S\in\mathcal{\widehat{S}}$ and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm terminates in $O(n^{2})$ time and either produces the Pareto-efficient allocation that it finds in its RP-implementing Step 2 or an allocation that intensity-dominates it.
-Proof.
+Consider a linear $\dot{\succsim}_{l}\in\mathcal$$\widetilde${I}} and assume that it induces the linear order $\succ_{l}$ above.
+We will show that the “upper half” of this intensity ordering—in which the left item in each pair is preferred to its right counterpart—exhibits the following pattern:
+$(a_{1},a_{n})$
+$\dot{\succ}_{l}$
+$(a_{1},a_{n-1})\,\dot{\sim}_{l}\,(a_{2},a_{n})$
+$\vdots$
+$(a_{1},a_{3})\,\dot{\sim}_{l}\,(a_{2},a_{4})\,\dot{\sim}_{l}\,\cdots\,\dot{\sim}_{l}\,(a_{n-2},a_{n})$
+$\dot{\succ}_{l}$
+$(a_{1},a_{2})\,\dot{\sim}_{l}\,(a_{2},a_{3})\,\dot{\sim}_{l}\,(a_{3},a_{4})\,\dot{\sim}_{l}\,\cdots\,\dot{\sim}_{l}\,(a_{n-1},a_{n})$
+$\dot{\succ}_{l}$
+$(a_{1},a_{1})\,\dot{\sim}_{l}\,(a_{2},a_{2})\,\dot{\sim}_{l}\,(a_{3},a_{3})\,\dot{\sim}_{l}\,\cdots\,\dot{\sim}_{l}\,(a_{n-1},a_{n-1})\,\dot{\sim}_{l}\,(a_{n},a_{n})$
+The bottom two lines are true by reflexivity of $\dot{\succsim}_{l}$ and by the postulated linearity of $\dot{\succsim}_{l}$, respectively, while the $\dot{\succ}_{l}$ sign between them is implied by $a_{i}\succ_{l}a_{i+1}$ for all $i\leq n-1$.
+Consider now the preceding line.
+We know from the above that $s_{l}(a_{i},a_{i+1})=s_{l}(a_{i+1},a_{i+2})=s_{l}(a_{i+2},a_{i+3})>0$.
+Hence, by (6), $s_{l}(a_{i},a_{i+2})=s_{l}(a_{i+1},a_{i+2})$ is also true.
+Since $s_{l}(a_{i+1},a_{i+2})=s_{l}(a_{i+2},a_{i+3})$ holds too, (6) further implies $s_{l}(a_{i+1},a_{i+3})=s_{l}(a_{i+1},a_{i+2})$.
+Combined, these yield $s_{l}(a_{i},a_{i+2})=s_{l}(a_{i+1},a_{i+3})$ or, equivalently, $(a_{i},a_{i+2})\dot{\sim}_{l}$ $(a_{i+1},a_{i+3})$.
+This proves that the intensity-equivalence statements in that line are valid too, while the $\dot{\succ}_{l}$ sign following it is a direct consequence of (6).
+By finiteness, proceeding upwards in this fashion and applying the same argument repeatedly establishes the claimed pattern in the structure of $\dot{\succsim}_{l}$.
+(Equivalently, this pyramidal scheme shows how the top-half intensity-equivalence classes of $\dot{\succsim}_{l}$ are formed and ordered by $\dot{\succ}_{l}$.)
+Since there are $n$ alternatives in $A$, the above shows that there are n1n-1 intensity-equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
+By ($1$), $r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})$.
+This and the above-postulated $\succ_{l}$ ordering on AA jointly imply
+$s_{l}(a_{1},a_{n})$ $=$ $n-1$
+$s_{l}(a_{2},a_{n})=s_{l}(a_{1},a_{n-1})$ $=$ $n-2$
+$\vdots$
+$s_{l}(a_{1},a_{2})=\cdots=s_{l}(a_{n-1},a_{n})$ $=$ $1$
+$s_{l}(a_{1},a_{1})=\cdots=s_{l}(a_{n},a_{n})$ $=$ $0$
+Thus, in line with Lemma 5, we have $a_{i}\succ_{l}a_{j}\Leftrightarrow r_{l}(a_{i})>r_{l}(a_{j})$ and, in addition,
+[…]
+from which the claim follows. ∎
+Now consider the induced profile $P_{S}\in\mathcal{P}$.
+By definition, $b_{l}^{P_{S}}(a_{i})\equiv|\{a_{k}\in A:a_{i}\succ_{l}a_{k}\}|$.
+This and the given ordering $\succ_{l}$ on $A$ readily imply $b_{l}^{P_{S}}(a_{i})=n-i$.
+Therefore, for every linear intensity profile $S\in\mathcal{\widetilde{S}}$, every agent $l\leq n$, and every alternative $a_{i}\in A$,
+$r_{l}^{S}(a_{i})$ $=$ $b_{l}^{P_{S}}(a_{i})$
+$=$ $n-i$
+[…]
+Proof of Proposition 3.
+By the definitions of $b_{l}$ and $r_{l}$ in (10) and (11), for any $S\in\mathcal{\widehat{S}}$ and every $l\leq n$ there is a strictly increasing function $f_{l}:\{0,1,2,\ldots,n-1\}\rightarrow\{0,1,2,\ldots,{{n}\choose{2}}\}$ such that, for all $a\in A$,
+$r_{l}^{S}(a)$ $\equiv$ $f_{l}(b_{l}^{P_{S}}(a))$
+Hence, for $x,y\in\mathcal{A}$ it holds that
+$\sum_{i=1}^{n}b_{l}^{P_{S}}(x_{i})>\sum_{i=1}^{n}b_{l}^{P_{S}}(y_{i})$ $\Longrightarrow$ $\sum_{i=1}^{n}r_{l}^{S}(x_{i})>\sum_{i=1}^{n}r_{l}^{S}(y_{i})$ (39)
+To prove the first claim, suppose to the contrary that $x\in\mathsf{IP}(S)$ but $x\not\in\mathsf{Bo}(S)$.
+Then, there must be $y\in\mathcal{A}$ such that $\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(y_{i})>\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(x_{i})$.
+By (39), this contradicts the postulate $x\in\mathsf{IP}(S)$.
+The second claim is proved in the main text.
+$\square$
+Proof of Proposition 4. In the main text.
+Proof of Proposition 5.
+It suffices to note that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains $|\dot{\succsim}_{l}|=2{n\choose 2}+n=(n-1)(n+1)$ comparisons between pairs in $A\times A$.
+$\square$
+Proof of Proposition 6.
 Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step.
@@
 Worst-case performance is therefore associated with a total number of searches equalling
-$\sum_{m=1}^{n-1}(n-m)$ $=$ $n^{2}-n-\sum_{m=1}^{n-1}m$
-$=$ $n^{2}-\sum_{m=1}^{n}m$
-Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in $O(n^{2})$ time.
+$\begin{array}[]{llclcllll}\sum\limits_{m=1}^{n-1}(n-m)&=&n^{2}-n-\sum\limits_{m=1}^{n-1}m&=&n^{2}-\sum\limits_{m=1}^{n}m&=&&&\\ &&&&&&\\ &&n^{2}-\dfrac{n}{2}(n+1)&=&\dfrac{1}{2}n(n-1)&=&\displaystyle{{n\choose k}}&&\end{array}$
+Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in at most ${n}\choose{2}$} steps, hence in $O(n^{2})$ time.
 It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$.
@@
 By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap.
-Thus, $y$ must be one of the following:
-(i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$
-(ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$
-(iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$
+Thus, $y$ must be one of the following: (i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$; (ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$; (iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$.
 In each case, $yDx$.
 Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
-Denote by $(V^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $V$ and $A$ into a (possibly empty) top-preference diverse set $V^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $V^{i>1}$ with respect to $A^{i>1}$, $i\leq k$.
-By the properties of these sets profiles, and by the definition of algorithm’s Step 4, irrespective of the $\gg$-order on $V$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in V^{1}$.
-Thus, $y_{i}=x_{i}$ for all $v_{i}\in V^{1}$.
-By the properties of $(V^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in V^{l>1}$ and $v_{j}\in V^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct.
-It follows that $y$ differs from $x$ by at most $\frac{n}{2}$ or $\frac{n-1}{2}$ pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such swap $x_{i}\leftrightarrow x_{j}$ occurring within the same top-preference-coincident pair $(V^{l>1},A^{l>1})$.
-By Step 4, finally, every such swap reassigns the jointly preferred (by Pareto-efficiency of $x$) item within the relevant pair in the direction of intensity-dominance.
+Denote by $(N^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $N$ and $A$ into a (possibly empty) top-preference diverse set $N^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $N^{i>1}$ with respect to $A^{i>1}$, $i\leq k$.
+By the properties of these sets and profile, and by the fourth step of the algorithm, irrespective of the $\gg$-order on $N$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in N^{1}$.
+Thus, $y_{i}=x_{i}$ for all $v_{i}\in N^{1}$.
+By the properties of $(N^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in N^{l>1}$ and $v_{j}\in N^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct.
+It follows that $y$ differs from $x$ by at most $n\over 2$} or $n-1\over 2$} pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such $x_{i}\leftrightarrow x_{j}$ swap occurring within the same top-preference-coincident pair $(N^{l>1},A^{l>1})$.
+By Step 4, finally, every such swap reassigns the—by Pareto-efficiency of xx—jointly preferred item within the relevant pair in the direction of intensity-dominance.
 Hence, $yDx$.
-∎
-The allocation delivered by this algorithm is not intensity-efficient in general.
-This is so because the forced exit of agents and items in the fourth step—a crucial feature for the algorithm’s manageable quadratic-time complexity—means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation.
-Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP.
-It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.
-In light of the discussion preceding the statement of Proposition 4, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency.
-Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 3, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully.
-The answer is negative.
-To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$.
-The possible orderings over $V$ are: 1.
-$v_{1}\gg v_{2}\gg v_{3}$; 2.
-$v_{1}\gg v_{3}\gg v_{2}$; 3.
-$v_{2}\gg v_{1}\gg v_{3}$; 4.
-$v_{2}\gg v_{3}\gg v_{1}$; 5.
-$v_{3}\gg v_{1}\gg v_{2}$; 6.
-$v_{3}\gg v_{2}\gg v_{1}$.
-The two Pareto-efficient allocations are $(a,b,c)$ and $(b,a,c)$, and the former intensity-dominates the latter.
-Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with probability $\frac{1}{2}$).
-If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it.
-However, if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e.
-with probability $\frac{1}{2}$) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
-Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem.
-Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up.
-Abdulkadiroğlu et al.
-(2011), however, followed by Pycia (2011); Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for.
-In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools.
-In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible.
-In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.
+$\square$

### v8

@@
-Intensity-Efficient and Intensity-Positional Allocations
+Efficient Allocation with Ordinal Preference Intensities
 Abstract
-This paper studies the allocation problem in environments where, in addition to having ordinal preferences, agents also have ordinal intensities: they can say “I prefer $a$ to $b$ more than I prefer $c$ to $d$ without necessarily being able to claim how much more.
-A rank-based criterion for interpersonal comparisons of ordinal intensities is introduced for this analytical environment.
-Building on it, the concepts of intensity-dominated, intensity-efficient and intensity-positional allocations are defined.Two allocations are related by intensity dominance if they assign the same two items to the same two agents but in a “flipped” way, and one allocation assigns the commonly preferred item in any such pair to the agent who prefers it more.
-An allocation is intensity-efficient if it is Pareto-efficient and not intensity-dominated.
-Distinctly, an allocation is intensity-positional if it maximizes the agents’ total intensity-rank scores.
-The latter extend standard Borda scores to environments where consecutively-ranked items do not necessarily feature identical intensity differences.
-The paper analyses intensity-efficient and intensity-positional allocations, and clarifies how they differ from those based on the utilitarian or Borda rules.
+Standard ordinal allocation methods ignore how strongly agents value different improvements, while cardinal methods require additional assumptions that are often considered too demanding.
+This paper studies assignment problems in the middle ground environment of ordinal preference intensities where agents can rank alternatives as well as preference improvements.
+The two criteria it proposes—intensity-efficient and intensity-positional allocations—use this combined information to refine Pareto efficiency via a novel dominance concept and a generalization of Borda-style scoring, respectively.
+These criteria point in new directions where stronger welfare gains are possible without cardinal utility or monetary transfers assumptions.
 “Suppose I am left with a ticket to a Mozart concert I am unable to attend and decide to give it to one of my closest friends.
@@
 1 Introduction
-Distributional principles that are used in the assignment of goods to individuals can be evaluated jointly on the realism of their assumptions and the appeal—according to some justice criterion—of their prescribed allocations.
-Arguably the most important such economic principle is Pareto-efficiency: assuming only information on agents’ preference rankings over the relevant goods, its notion of optimality is defined by the requirement that no changes can make some individuals better off without hurting others.
+1.1 Motivation
+Distributional criteria for the assignment of goods to individuals can be evaluated jointly on the realism of their assumptions and the appeal—according to some justice principle—of their prescribed allocations.
+Arguably the most important such criterion is Pareto efficiency: assuming only information on agents’ preferences over the relevant goods, its notion of optimality is defined by the requirement that no changes can make some individuals better off without hurting others.
 Yet it is well-known that among the typically many Pareto-efficient allocations there are often some which violate intuitive notions of distributive justice.
-Appreciation of these facts raises the question of whether one can narrow down and improve upon Pareto’s criterion by considering alternative principles which, in addition to agents’ preferences, also incorporate information on their preference intensities.
+Appreciation of these facts raises the question of whether Pareto’s criterion might be improved on by considering alternative principles which, in addition to agents’ preferences, also incorporate information on their preference intensities.
 Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities.
-Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered.
-Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion, formalized in Harsanyi (1955), is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Moulin, 1988, 2004).
-Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Alaoui and Penta, 2025; Samuelson, 2025).
-More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982).
-This stands in contrast to some contemporary views in the field of cognitive psychology and decision-making, according to which “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” (Vlaev et al., 2011, p.
-551).
-To the extent that comparing options is easier and more reliable than assigning values to them, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
-Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made.
-Specifically, instead of cardinal-welfarist or Arrowian profiles that, respectively consist of interpersonally comparable cardinal utility functions—possibly not arising from expected-utility preferences (Nebel, 2024; Weymark, 2005; Harvey, 1999)—or preference orderings, this domain comprises profiles of generally ordinal preference intensity orderings, i.e.
-collections of quaternary relations that reflect the agents’ preference intensity comparisons.
-This analytical choice assumes that, similar to the way in which preferences are elicited, e.g.
-in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
-This framework where both preference and—generally only ordinal—intensity information is available to the social planner naturally raises the question of how it might be used, as per the goals that we set out above, to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ intensities and mitigate the multiplicity of Pareto-efficient allocations.
-Similar to utilitarian and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons.
-Unlike those notions, however, where such comparisons are based on agents’ utilities (Fleurbaey and Hammond, 2004; d’Aspremont, 2018; Echenique et al., 2023; Baccelli, 2023), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
-To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful.
-In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more.
-In Section 3 we formalize and motivate this interpretational assumption concerning interpersonal comparisons of ordinal intensities in the present analytical environment.
-We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781; Young, 1974; Mas-Colell et al., 1995; Maskin, 2025) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (d’Aspremont, 2018; Baccelli, 2023).
-In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just.
-Specifically, we define an allocation $x$ to be intensity-efficient if it is Pareto-efficient and also such that, whenever another Pareto-efficient allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e.
-when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more.
-Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile when there are three agents and alternatives.
-We also show that existence is not guaranteed when there are more agents and items, because the underlying dominance relation may be then be cyclic.
-Motivated by this, for the general case of $n$ agents and as many items we establish a sufficient condition for existence, requiring profiles to be (effectively) homeo-monotonic.
-The key part of this condition could be summarized informally as follows: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any item in this subset over that item’s immediate successor, then this ordering cannot be violated when it comes to comparing the agents’ intensities between any non-consecutively ranked items in this subset.
-Importantly, although the presence of such groups—and the corresponding restrictions within them—is not necessary for existence, we show that if a profile does give rise to such a group, and if the above-mentioned ordering over its agents is strict, then homeo-monotonicity in that group is indeed necessary for an intensity-efficient allocation to exist.
-Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible.
-Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons do emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide.
-They do so, however, when the said utility profile satisfies an intuitive additional condition that we introduce.
-This requires any interpersonal dominance that emerges by comparing two agents’ utility differences to be accompanied by an interpersonal dominance in the same direction in the rank-order positions of these utility differences in the respective agents’ intensity orderings.
-In Section 5 we build on a stronger interpersonal comparisons assumption to propose and analyse the concept of intensity-positional allocations.
-Toward defining them, we recall first than an item’s Borda score for some agent is the number of items that are ranked below it by that agent.
-Although Borda scores give rise to a specific normalized ordinal utility representation of the agent’s preferences, in practice they are often used with a cardinal interpretation (see, for example, Maskin, 2025) due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where the agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart.
+Economists have traditionally interpreted these to be synonymous to the existence of a cardinal utility function that represents an agent’s preferences over alternatives.
+Typically, this has been of the von Neumann and Morgenstern (1947) (vNM) kind, derived from expected-utility preferences over lotteries over alternatives.
+Assuming that utility differences from such functions are interpersonally comparable, the classical utilitarian summation criterion, formalized in Harsanyi (1955), has been the benchmark distributional principle of the cardinal-welfarist tradition (Moulin, 1988; Roemer, 1996; Fleurbaey and Hammond, 2004; d’Aspremont, 2018; Echenique, Immorlica, and Vazirani, 2023), refining the set of Pareto-efficient allocations in intensity-inclusive directions.
+Despite the undeniable tractability and widespread application of this and related criteria in various domains of economic analysis, it is well-understood that cardinal utility functions emerging from expected-utility preferences over lotteries confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Alaoui and Penta, 2025; Samuelson, 2025).
+This fact, in turn, has generated concerns—such as the one expressed by Kenneth Arrow in the cited quotation—about the epistemic origins of such functions.
+In addition, and no less importantly, when cardinal utility differences are used to define an agent’s intensity comparisons, the person in question is portrayed as having a supernaturally high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982).
+Acknowledging the logical implication of this fact—namely, that elicitation of agents’ utilities in reality may be subject to “noise”—diminishes a planner’s confidence that application of utilitarian or related criteria in practical allocation problems is guaranteed to achieve the intended distributive outcome.
+In the distinct but related domain of evaluative voting problems, for example, where agents rate the candidates on numerical scales, Baujard, Gavrel, Igersheim, Laslier, and Lebona (2018); Baujard, Igersheim, and Lebon (2021) provide experimental evidence suggesting that the use of different scales leads to non-equivalent voting results.
+1.2 Contribution
+This skepticism notwithstanding, intensity comparisons such as those in Harsanyi’s (1987) opening quotation are common in everyday life and, in fact, do not presuppose the existence of interpersonally comparable utilities.
+They do, however, implicitly assume that:
+(i) agents can rank alternatives according to their preferences and rank preference improvements according to the intensity of these preferences;
+(ii) systematic interpersonal comparisons of such second-order data are possible.
+Given all the above, this paper’s central question can now be stated: “When agents cannot provide credible cardinal utilities, but can report which preference improvements feel larger than others, can a planner use that information to improve upon Pareto-efficient allocations?” We answer this question by focusing on the classic assignment problem where finitely many indivisible goods must be assigned to as many agents.
+To address requirement (i) we assume that agents’ preference intensity orderings belong to the broad class of those that admit a generally ordinal numerical representation by means of a preference intensity function (Gerasimou, 2021).
+In effect, this portrays each individual as being able to rank the above-mentioned preference improvements in a coherent way, yet without necessarily being able to quantify them.
+This modeling approach is one possible way of dealing with the concern expressed earlier regarding cardinal utilities, and is in line with some contemporary views in cognitive psychology that are elegantly summarized in Vlaev, Chater, Stewart, and Brown (2011, p.
+551) with the remark that “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” To address requirement (ii), moreover, we employ an intuitive normalization of agents’ preference intensity functions and introduce an interpersonal comparability assumption that is suitable for this kind of ordinal intensity data.
+With this interpretational assumption in the background, finally, we propose two refinements of Pareto efficiency—intensity-efficiency and intensity-positionality—that incorporate such data and discard allocations that are intuitively inferior to others in the light of this additional information.
+To illustrate both the interpersonal comparability assumption and the intensity-efficiency criterion, imagine, for example, that Anna and Baba have identical preferences that put ($a$)pple above ($b$)anana, and the latter above ($o$)range.
+Suppose Anna thinks the preference improvement from b to a is bigger than the improvement from o to b, and that the opposite is true for Baba (clearly, the improvement from o to a is the biggest one for both).
+Who should get the apple if it is also known that the third person in the group likes the orange most?
+The Pareto criterion, as well as the Pareto-refining Borda criterion (Borda, 1781; Young, 1974; Mas-Colell et al., 1995) that we recall two paragraphs below, delivers a tied recommendation based solely on the agents’ preferences: either Anna or Baba should get it.
+By contrast, the intensity-efficiency criterion assigns it to Anna.
+The reason is that the preference improvement from b to a lies in a higher position in Anna’s intensity ordering than in Baba’s.
+The interpersonal comparability condition translates this fact into Borda-style intensity scores on a common scale (ranging from 1 to 3 here) and, in that spirit, declares Anna’s improvement to be stronger than Baba’s because pair $(a,b)$ has a score of 2 for Anna and 1 for Baba.
+Intensity-efficiency requires that, other things equal, whenever two Pareto allocations assign the same two items to the same two agents but in a “flipped” way, and one allocation assigns the commonly preferred item in every such pair to the agent who prefers it more, then that allocation should be favored over the alternative, which is discarded as being intensity-dominated.
+Assuming strict preferences and ordinal intensities, in Theorem 1 we show that an intensity-efficient allocation exists under every profile when there are three agents and alternatives.
+Existence is not guaranteed when there are more because the underlying intensity-dominance relation may be cyclic.
+For the general case of $n$ agents and items, then, we establish a sufficient condition for existence, requiring profiles to be (effectively) homeo-monotonic (Theorem 2).
+The key part of this condition can be summarized informally as follows: when a group of $k\leq n$ agents have identical preferences over their top $k$ items, and the $k$ agents can be weakly ordered in how much they prefer any item in this subset over that item’s immediate successor, this ordering cannot be violated when agents’ intensities between any non-consecutively ranked items are compared.
+For example, consider a group that consists of agents $L_{1}$, $L_{2}$, $L_{3}$ and items $l_{1}$, $l_{2}$, $l_{3}$ over which agents have the identical preferences $l_{1}\succ l_{2}\succ l_{3}$ but $L_{1}$ ranks the improvements from $l_{3}$ to $l_{2}$ and from $l_{2}$ to $l_{1}$ higher than $L_{2}$, who ranks them higher than $L_{3}$.
+Then, homeo-monotonicity requires the improvement from $l_{3}$ to $l_{1}$ be ranked weakly higher by $L_{1}$ too, followed by $L_{2}$ and then by $L_{3}$.
+Although the presence of such groups is not necessary for intensity-efficient allocations to exist, if a profile gives rise to a group with the antecedent structure over consecutively-ranked items, and if the relevant ordering over its agents is strict, then homeo-monotonicity within that group is indeed necessary for existence (Theorem 3).
+Intensity-efficiency is the first intensity-inclusive refinement of Pareto-efficiency that operates in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible.
+Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons do emerge from cardinal utility functions that are interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide (Proposition 1).
+The reason is that an agent’s utility difference between two items may be greater than another agent’s in absolute terms but still have a lower intensity score.
+When no such discrepancy arises, utilitarian allocations are indeed intensity-efficient (Proposition 2).
+By its definition, the intensity-dominance relation on which the intensity-efficiency criterion rests is not always applicable.
+Motivated by this fact, we later introduce a stronger interpersonal comparability assumption and, building on it, propose the concept of intensity-positional allocations.
+Toward defining it, recall that an item’s Borda score for some agent is the number of items that this person ranks below it.
+Importantly, although Borda scores give rise to a specific normalized ordinal utility representation of the agent’s preferences, in practice these are often used with a cardinal interpretation (see, for example, Maskin, 2025 and, for recent experimental evidence, Ambuehl and Bernheim, 2026).
+This is due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where the agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart.
 We will refer to these intensity relations as “linear”.
-Similar to Borda allocations, which are defined as those that maximize the aggregate Borda score, intensity-positional allocations are defined as the maximisers of the aggregate intensity-rank score.
-Unlike an item’s Borda score, however, which reveals the position of that item on the agent’s preference ordering, its intensity-rank score corresponds to the highest rank-order position in the agent’s intensity ordering where this item appears as preferred to something else.
-Hence, it encodes information both about the number of alternatives that are ranked below this item and about the relative intensity differences between it and the items it is preferred to.
+Similar to Borda allocations, which are defined by maximization of the aggregate Borda score, intensity-positional allocations are those that maximize the aggregate intensity-rank score.
+Unlike, say, item $a$’s Borda score, however, which reveals the relative position of $a$ in the agent’s preference ordering, $a$’s intensity-rank score corresponds to the highest-ranked preference improvement towards item $a$.
+The latter, in turn, is simply the position of pair $(a,w)$ in the agent’s intensity ordering, where $w$ is that person’s bottom-ranked item.
+An item’s intensity-rank score, therefore, encodes information both about the number of alternatives that are ranked below it and about the relative intensity differences between it and the items it is preferred to.
 This information, in turn, is reflected in the prescribed allocation process.
-In the paper’s focal case of strict intensity profiles we show that intensity-positional allocations refine standard Borda allocations.
-We further show that, in a larger domain where a like-for-like comparison remains meaningful, intensity-rank and Borda-score aggregation diverge in general but coincide at linear profiles.
-The intuitive reason is that, unlike intensity-positional aggregation, Borda aggregation treats every profile as it if were linear.
-Thus, the intensity-positional criterion is a genuine generalization of the standard Borda-count criterion in this extended domain.
+In the paper’s focal case of strict intensity profiles we show that intensity-positional allocations refine Pareto-efficient allocations and are distinct from Borda allocations (Propositions 3 and 4).
+We further show that, in a larger domain where a like-for-like comparison remains meaningful, intensity-rank and Borda-score aggregation diverge in general but coincide at linear profiles (Theorem 4).
+The intuitive reason is that, unlike intensity-positional aggregation, Borda aggregation treats every profile as if it were linear.
+Thus, the intensity-positional criterion is a genuine generalization of the standard Borda-count criterion in this extended domain, enabling the analyst to select Pareto-efficient allocations with desirable welfare properties which might otherwise have been discarded.
 Intensity-positional and intensity-efficient allocations are logically distinct.
@@
 A disadvantage is that making intensity-positional allocations operational implicitly invokes stronger interpersonal comparability assumptions.
-In Section 6, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon the Pareto-efficiency desideratum.
-We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited in quadratic time by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class that we focus on.
-Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) mechanism’s algorithm and produces—in quadratic time in the number of items/agents—an allocation which intensity-dominates the Pareto-efficient one that is found by RP.
-While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm.
-We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
-As far as the broader related literature is concerned, this paper’s contribution lies at the intersection of welfare economics, social choice and matching theory, in all three cases building bridges between the existing ordinal- and cardinal-welfarist paradigms.
-Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Moulin, 2004) and others, the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process.
-Unlike these refinements, the hereby proposed ones are the first that are operational in the general class of ordinal preference-intensity profiles, and build on novel interpersonal-comparability criteria that are distinct from interpersonal comparisons of utility differences which the above notions rely on.
-Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain.
-Intensity-positional allocations on the other hand are indeed derived from a social welfare function.
-The problem of characterizing this function axiomatically is important and deserves separate treatment.
-In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs into the allocation process toward maximizing utilitarian welfare.
-Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer.
-While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
-2 Measurement-Theoretic Foundations
+1.3 Related Literature
+The paper contributes to welfare economics, social choice and allocation theory, in all three cases building bridges between the ordinal- and cardinal-welfarist paradigms.
+In particular, similar to cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), the hereby proposed refinements of Pareto efficiency are inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process.
+Unlike the former refinements, those studied here are the first that are applicable in principle in the general class of ordinal preference-intensity profiles.
+Furthermore, they build on a novel analytical environment and interpersonal-comparability criteria that are distinct from interpersonal comparisons of cardinal utility differences.
+The previous subsection also clarifies how they differ from the ordinal but “interpretationally cardinalized” classic Borda scoring method.
+Although the building blocks differ in important ways, conceptually this paper is closely related to Abdulkadiroğlu, Che, and Yasuda (2011) (see also Miralles, 2009; Pycia, 2011; Carroll, 2018) and the several studies in matching theory with one- and two-sided preferences that have taken agents’ interpersonally comparable vNM cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs into the allocation process toward maximizing utilitarian welfare.
+Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching by incorporating intensity-relevant information.
+Its distinguishing point is that it is the first to pursue this goal in the unexplored environment where only ordinal intensity comparisons are generally available to the designer.
+While this environment is less tractable for the modelling analyst, it is cognitively simpler for the economic agent and, as such, we believe worthy of further investigation.
+2 Preliminaries
 By $A:=\{a_{1},\ldots,a_{n}\}$ and $N:=\{1,\ldots,n\}$ we denote, respectively, the finite sets of choice alternatives and agents.
@@
 When no ambiguity arises, we will interpret $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ more succinctly as suggesting that the intensity difference between $a$ and $b$ is weakly larger than that between $c$ and $d$.
-In line with these interpretations and the literature, agent $l$’s binary preference relation $\succsim_{l}$ on $A$ is derived from their intensity relation $\dot{\succsim}_{l}$ by $a\succsim_{l}b\Leftrightarrow(a,b)\,\dot{\succsim}_{l}\,(b,a)$.
+In line with these interpretations, and with the literature, agent $l$’s binary preference relation $\succsim_{l}$ on $A$ is derived from their intensity relation $\dot{\succsim}_{l}$ by $a\succsim_{l}b\Leftrightarrow(a,b)\,\dot{\succsim}_{l}\,(b,a)$.
 We will refer to $\succsim_{l}$ as the preference relation that is induced by $\dot{\succsim}_{l}$.
@@
 $(A\times A)_{\dot{\sim}_{l}}$ $:=$ $\{[a,b]:[a,b]\text{ is an intensity-equivalence class under }\dot{\sim}_{l}\}$ (2)
-The definitions that follow omit universal quantifiers, yet they should be understood as imposing conditions that apply to all objects in the respective domains.
-Definition 1.
 […]
@@
 Baccelli (2024) has recently clarified, however, that every ordinal utility representation defines a certain subset of utility-difference comparisons whose order is, in fact, preserved by all ordinal transformations of the original representation.
-Scott (1964), completing and extending the analysis that was initiated in Scott and Suppes (1958), characterized the existence of a utility-difference representation over a finite set by means of the following three axioms on $\dot{\succsim}_{l}$:
-$(\mathbf{1_{\mathit{D}}})$ [Completeness]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ or $(c,d)\,\dot{\succsim}_{l}\,(a,b)$.
-$(\mathbf{2_{\mathit{D}}})$ [Cancellation].
-$(b_{i},c_{i})\,\dot{\succsim}_{l}\,(b_{\pi(i)},c_{\sigma(i)})$ for all $i<t$, $i>0$, implies $(b_{\pi(0)},c_{\sigma(0)})\,\dot{\succsim}_{l}\,(b_{0},c_{0})$ for all sequences $b_{0},b_{1},\ldots,b_{t-1},c_{0},\ldots,c_{t-1}\in A$ and all permutations $\pi,\sigma$ of $\{0,\ldots,t-1\}$, where $t>0$.
-$(\mathbf{3_{\mathit{D}}})$ [Reversal]. $(a,b)\,\dot{\succsim}_{l}\,(c,d)$ implies $(d,c)\,\dot{\succsim}_{l}\,(b,a)$.
-The first imposes universal comparability on $\dot{\succsim}_{l}$.
-The third is essential for $\dot{\succsim}_{l}$ to be interpreted as a preference intensity relation rather than as an arbitrary preference relation over pairs.
-The second is known to have “no simple way of summarizing what it says about choice behavior” (Luce and Suppes, 1965, p.
-277).
-Scott (1964, p.
-244-5) explained how $(\mathbf{2_{\mathit{D}}})$ implies transitivity of $\dot{\succsim}_{l}$, but also remarked that $(\mathbf{2_{\mathit{D}}})$ is an infinite bundle of conditions (for each $t>0$, each $\pi,\sigma$), and it was shown in Scott and Suppes (1958) that no finite number of them could be sufficient”.
-Once the number $n$ of alternatives in $A$ is fixed, there are obviously finitely many such conditions to be satisfied.
-Scott’s remark, however, clarifies that it is impossible to reduce this condition to a finite number of statements/axioms that, together with $(\mathbf{1_{\mathit{D}}})$ and $(\mathbf{3_{\mathit{D}}})$, would be sufficient to characterize (3) for any finite $n$.
-In Gerasimou (2021, 2022) this author recently proposed and characterized the class of ordinal intensity relations that admit the following more general numerical representation.
-Definition 2.
 […]
@@
 In words, $\widehat{s}_{l}$ represents the agent’s intensity weak ordering [(4)] and the preferences induced by it [(5)], further ensuring that these preferences are themselves weakly ordered and intensities are monotonically increasing in the preference ordering [(6)].
-Such a representation is possible if and only if $\dot{\succsim}_{l}$ is a weak order on $A\times A$ that additionally satisfies $(\mathbf{3_{\mathit{D}}})$ and the following condition, which states that $a$ is preferred to $b$ if and only if the intensity difference between $a$ and $c$ is higher than that between $b$ and $c$.
-$(\mathbf{4_{\mathit{D}}})$ [Translocation Consistency]. $(a,c)\,\dot{\succsim}_{l}\,(b,c)$ if and only if $(a,b)\,\dot{\succsim}_{l}\,(b,a)$.
 We denote by $\mathcal{I}$ the class of intensity relations that admit this ordinal representation.
@@
 Lemma 1.
-A relation $\dot{\succsim}_{l}$ that is represented by $\widehat{s}_{l}$ is utility-difference representable if and only if some strictly increasing transformation $\widetilde{s}_{l}$ of $\widehat{s}_{l}$ satisfies
+A relation $\dot{\succsim}_{l}$ that is represented by a preference intensity function $\widehat{s}_{l}$ is utility-difference representable if and only if a strictly increasing transformation $\widetilde{s}_{l}$ of $\widehat{s}_{l}$ satisfies
 $\widetilde{s}_{l}(a,c)$ $=$ $\widetilde{s}_{l}(a,b)+\widetilde{s}_{l}(b,c)$ (7)
-The proof of this and most other results appear in the Appendix.
-The normalization of $\widehat{s}_{l}$ introduced next is important for the rest of this paper.
-Definition 3.
-A preference intensity function $\widehat{s}_{l}$ represents $\dot{\succsim}_{l}$ canonically if
+The proof of this and most other results appears in the Appendix.
+Next, we will say that a preference intensity function $\widehat{s}_{l}$ represents $\dot{\succsim}_{l}$ canonically if
 $\widehat{s}_{l}(A\times A)$ $=$ $\{-k,-k+1,\ldots,-1,0,1,\ldots,k-1,k\},$ (8)
-with $k$ the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
+where $k$ is the number of intensity equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
 We will denote by $s_{l}$ the canonical intensity function that represents $\dot{\succsim}_{l}$.
-The usefulness of this normalization for our purposes lies in the fact that the integer it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
+As will be clarified later, the usefulness of this normalization for our purposes lies in the fact that the integer value it associates with a pair reveals the rank position of that pair in the agent’s intensity ordering.
 Lemma 2.
@@
 Our subsequent analysis will further impose the following additional property on $\dot{\succsim}_{l}$:
-$(\mathbf{5_{\mathit{D}}})$ [Strictness]. $(a,b)\,\mathord{\not\mathrel{\dot{\sim}_{l}}}\,(c,d)$ for all distinct off-diagonal pairs $(a,b),(c,d)\in A\times A$.
+Strictness. $(a,b)\,\mathord{\not\mathrel{\dot{\sim}_{l}}}\,(c,d)$ for all distinct off-diagonal pairs of items $(a,b)$ and $(c,d)$.
 This condition rules out the possibility of the agent preferring $a$ to $b$ exactly as much as they prefer $c$ to $d$ for distinct pairs of distinct alternatives $(a,b)$ and $(c,d)$.
@@
 As we elaborate in Section 3, strict intensities are useful for our purposes because they ensure that the canonical intensity functions of all agents are onto the same set $\{-k,\ldots,-2,-1,0,1,2,\ldots,k\}$, where $k\equiv{n\choose 2}$.
-In particular, the only non-trivial intensity-equivalence class when $\dot{\succsim}_{l}$ satisfies $(\mathbf{5_{\mathit{D}}})$ is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
-For illustration, the novel computational output in columns 3–4 of Table 1 lists the number of distinct strict intensity orderings that admit,respectively, a (pseudo-cardinal) utility-difference representation and an ordinal representation when $n\leq 5$.
-Table 1: The number of distinct intensity relations of different classes for small values of $n$.
-Strict & utility-difference Strict Strict ordinal
-Strict representable ordinally representable intensity
-$n$ orders intensity relations intensity relations profiles
-3 6 12 12 $12^{3}$
-4 24 240 384 $384^{4}$
-5 120 13,680 92,160 $92,160^{5}$
-Source: author’s computations with the Minizinc constraint-programming tool (Nethercote et al., 2007).
+In particular, the only non-trivial intensity-equivalence class when $\dot{\succsim}_{l}$ satisfies Strictness is the diagonal subset of $A\times A$, with every pair $(a,a)$ in this set being mapped to 0 by the canonical representation of $\dot{\succsim}_{l}$.
 We conclude this section by commenting on possible sources of intensity information.
@@
 This is different from a lab, field or other environment that might, in principle, provide the analyst with agents’ choices at several menus at different points in time, together with intensity-revealing information at those menus.
-As we elaborate formally in Section 6, for this specific environment we assume that agents’ ordinal preferences are available, and that their ordinal intensities are elicited by asking them directly—and conditional on their submitted preferences—the kinds of simple binary questions that were mentioned earlier.
+For this specific environment we assume that agents’ ordinal preferences are available, and that their ordinal intensities are elicited by asking them directly—and conditional on their submitted preferences—the kinds of simple binary questions that were mentioned earlier and which pertain to ranking preference improvements.
 3 Interpersonal (Non-)Comparability of Ordinal Intensities and Intensity-Rank Scores
 […]
-Assumption 1 (for Sections 4 & 5)
+Assumption 1 (for Sections 4–5)
 Given an intensity profile $S\in\mathcal{\widehat{S}}$ and the corresponding canonical profile $s$, the statement
@@
 4 Intensity-Dominance and Efficiency
-4.1 Intensity-Dominance and Efficiency
 Since the number of agents and items are assumed to coincide, an allocation of the $n$ goods is a permutation on $A$.
@@
 Among them, $r$ is intensity-efficient: it dominates $x$; is incomparable to $p$ and $y$; and is equivalent to $q$, $z$ (these are dominated by $p$, $y$, respectively).
-Yet $N^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{1}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$.
+Yet $N^{2}$ is (non-strictly) monotonic but not homeo-monotonic with respect to $A^{2}$: $s_{1}(a_{1},a_{2})=s_{2}(a_{1},a_{2})>s_{3}(a_{1},a_{2})$ and $s_{1}(a_{2},a_{3})=s_{2}(a_{2},a_{3})>s_{3}(a_{2},a_{3})$, but $s_{1}(a_{1},a_{3})=s_{2}(a_{1},a_{3})<s_{3}(a_{1},a_{3})$.
 $\blacklozenge$
@@
 $:=$ $\mathsf{IP}(S)$
-Analogously, a Borda allocation at profile $S\in\mathcal{\widehat{S}}$—henceforth denoted by $\mathsf{Bo}(S)\equiv\mathsf{Bo}(P_{S})$—is defined as in (14) but with the Borda scores $b_{l}^{P}$ at the strict preference profile $P_{S}$ replacing the intensity-rank scores $r_{l}^{S}$ at the strict intensity profile $S$.
+An intensity-positional allocation maximizes the total of agents’ ordinal intensity-rank scores at their assigned items, and thus resolves trade-offs across agents globally rather than only through pairwise dominance comparisons.
+Indeed, as was noted earlier, an item’s intensity-rank score contains information about the number of alternatives that are ranked below it as well as about the relative intensity difference between that item and the one ranked at the bottom.
+This double channel—and trade-offs thereof—guides the intensity-positional allocation criterion and makes it distinct from Borda’s criterion that is guided by the former channel only.
+By analogy, a Borda allocation at profile $S\in\mathcal{\widehat{S}}$—henceforth denoted by $\mathsf{Bo}(S)\equiv\mathsf{Bo}(P_{S})$—is defined as in (11) but with the Borda scores $b_{l}^{P}$ at the strict preference profile $P_{S}$ replacing the intensity-rank scores $r_{l}^{S}$ at the strict intensity profile $S$.
 Example 5.
@@
 Proposition 3.
-An intensity-positional allocation is a Borda allocation (hence Pareto-efficient): for all $S\in\mathcal{\widehat{S}}$, $\mathsf{IP}(S)\subseteq\mathsf{Bo}(S)$. The converse is false.
-In other words, Borda allocations—which, as is well-known, are Pareto-efficient—are refined by the intensity-positional criterion.
-That this refinement is generally strict can in turn be verified with the following $n=3$ example profile:
-$(a,c)\dot{\succ}_{1}(a,b)\dot{\succ}_{1}(b,c)$; $(a,c)\dot{\succ}_{2}(b,c)\dot{\succ}_{2}(a,b)$; $(c,a)\dot{\succ}_{3}(c,b)\dot{\succ}_{3}(b,a)$.
-The reader can readily check that both Pareto-efficient allocations $(a,b,c)$ and $(b,a,c)$ are selected at this profile under $\mathsf{Bo}$, whereas only the former is so under $\mathsf{IP}$ (and that this allocation is also intensity-efficient).
-This is intuitive: the first two agents have identical preferences, but the intensity rank of $b$ is higher for the second agent, reflecting the fact that preference intensity between $a$ and $b$ is higher for agent 1 than it is for agent 2.
-The intensity-positional criterion accounts for this interpersonal difference, whereas the Pareto and Borda criteria do not.
-In this example it so happened that the set of intensity-positional and intensity-efficient allocations coincided. Yet these criteria are distinct in general.
+An intensity-positional allocation is Pareto-efficient.
 Proposition 4.
-The sets of intensity-efficient and intensity-positional allocations at a profile $S\in\mathcal{\widehat{S}}$ are generally distinct.
-The first implication of this claim—namely, that an intensity-positional allocation may not be intensity-efficient—has already been established in Example 5.
-The converse implication can also be proved by modifying the profile of Example 3 (Table 4) so that $s_{2}(a_{1},a_{3})=5$, $s_{2}(a_{2},a_{4})=4$ and $s_{3}(a_{1},a_{2})=4$, $s_{3}(a_{2},a_{4})=3$, $s_{3}(a_{2},a_{3})=2$, $s_{3}(a_{3},a_{4})=1$.
+The sets of intensity-efficient, intensity-positional and Borda allocations are generally distinct.
+That an intensity-positional allocation may not be intensity-efficient has already been shown in Example 5.
+The converse implication can also be seen by modifying the profile of Example 3 (Table 3) so that $s_{2}(a_{1},a_{3})=5$, $s_{2}(a_{2},a_{4})=4$ and $s_{3}(a_{1},a_{2})=4$, $s_{3}(a_{2},a_{4})=3$, $s_{3}(a_{2},a_{3})=2$, $s_{3}(a_{3},a_{4})=1$.
 In this case, the set of intensity-efficient allocations contains three elements, one of which—namely, $(a_{3},a_{2},a_{1},a_{4})$—is not intensity-positional.
+As far as the comparison between intensity-positional and Borda allocations is concerned, in light of Example 5 it suffices to show that there are profiles where some allocations are of the former but not of the latter kind.
+To this end, let $n=3$ and suppose $s_{1}(a,c)=3$, $s_{1}(b,c)=2$, $s_{1}(a,b)=1$; $s_{2}(a,b)=3$, $s_{2}(a,c)=2$, $s_{2}(c,b)=1$; and $s_{3}(b,a)=3$, $s_{3}(c,a)=2$, $s_{3}(b,c)=1$.
+Allocation $(b,a,c)$ is intensity-positional but not Borda.
+$\square$
 This analysis uncovers that it is possible for allocations that are intensity-dominated to be intensity-positional.
@@
 This also helps clarify how the two concepts that are studied in this and the preceding section jointly provide two distinct and complementary ways in which an analyst can utilise ordinal intensity information in order to improve upon the Pareto criterion.
-6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
-We conclude with an illustration of the potential usefulness of the preceding analysis in matching-theoretic problems.
-In particular, we study a matching rule/mechanism that takes agents’ ordinal intensities alongside their ordinal preferences as its input and produces an allocation that improves upon Pareto-efficiency, either in the intensity-dominating or intensity-positional direction.
-Before doing so, we clarify how ordinal intensities can be elicited alongside preferences in the first place in order for such a mechanism to become operational.
-To this end, we show below how the properties of intensity relations that are representable by ordinal preference intensity functions can simplify this elicitation problem very considerably.
-Proposition 5.
-Given a strict preference order $\succ_{l}$, an intensity relation $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} that induces $\succ_{l}$ is elicited in at most (n1)(n+1)(n-1)(n+1) steps, hence in $O(n^{2})$ time.
-Turning now to the problem of using as input a strict intensity profile SS—and the strict preference profile $P_{S}$ nested within it—to find an allocation that improves upon Pareto-efficiency, we proceed by recalling first some important landmarks from the matching-theoretic literature with one-sided preferences where no intensity information is available.
-The study of such rules goes back to Shapley and Scarf (1974) and David Gale’s Top-Trading Cycles (TTC) algorithm.
-The premise there is that each agent has an initial endowment from the set of items that will be (re-)assigned once the algorithm is run.
-TTC then operates sequentially by implementing swaps between these initial assignments in accordance with the agents’ preferences, giving priority to an item’s current assignee in situations where additional agents prefer this over their own endowment.
-Shapley and Scarf (1974) proved that TTC leads to a core allocation.
-Roth and Postlewaite (1977) later showed that this core allocation, identifiable by TTC, is unique.
-Extending the results in Roth (1982), Ma (1994) established that TTC is the unique rule that leads to a Pareto-efficient allocation and, in addition, satisfies strategy-proofness and individual-rationality.
-Hylland and Zeckhauser (1979) introduced randomness to study the assignment problem without initial endowments and, assuming that all agents have an equal amount of some artificial currency and expected-utility preferences over lotteries over allocations, studied that which is associated with (or approximated by) Competitive Equilibrium from Equal Incomes.
-This mechanism is not strategy-proof.
-Also introducing randomness, but without assuming expected-utility preferences or employing the competitive-equilibrium apparatus, Abdulkadiroğlu and Sönmez (1998) showed that TTC with random initial endowments is equivalent to the Random Priority (RP) mechanism.
-The latter does not involve initial endowments; ranks agents under a random priority order; and lets them choose, accordingly, their most preferred item from those still available.
-These authors also clarified that a deterministic matching mechanism leads to a Pareto-efficient allocation if and only if is a Random Priority mechanism.
-Despite the ex post Pareto-efficiency of RP allocations, hence of core/TTC allocations under random endowments, Zhou (1990) showed that these are generally ex ante inefficient.
-Bogomolnaia and Moulin (2001), finally, demonstrated that they are also ordinally inefficient in the sense that they may be dominated by another random allocation under the stochastic-dominance relation that is defined by all agents’ preferences.
-In response to this realization, the latter authors introduced ordinally-efficient allocations as those that are not first-order stochastically dominated by any other random allocation, and proved that their Probabilistic Serial (PS) mechanism, although not strategy-proof, leads to an ordinally-efficient allocation with desirable—under expected-utility preferences over allocations—ex ante fairness properties.
-Informed by the main insights of the above-cited literature, we now consider the following algorithm for the problem at hand:
-Input: A strict intensity profile $S\in\mathcal{\widehat{S}}$.
-Step 1: Choose a linear order $\gg$ over $N$ from the uniform distribution with support $n!$.
-Step 2: Apply RP under $\gg$ on $P_{S}$ to arrive at a Pareto-efficient allocation $x$.
-Step 3: Reorder the entries in $x$ according to $\gg$, with $j>i$ for $x_{i},x_{j}$ if and only if $i\gg j$.
-Step 4: Search for the $\gg$-lowest agent $v_{j}\in N\setminus\{v_{1}\}$ such that $s_{1}(x_{j},x_{1})>s_{j}(x_{j},x_{1})$.
-If such an agent exists, swap $x_{1}$ with $x_{j}$ and remove $v_{1},v_{j}$ from $N$ and $x_{1},x_{j}$ from $A$.
-If no such agent exists, remove $v_{1}$, $x_{1}$.
-Repeat Step 4 until no agents remain.
-The first two steps setup the RP algorithm.
-The third one merely relabels the items in the allocation that was arrived at earlier, so that the first (second, third, …) item belongs to the first-(second-, third-, …)ranked agent in the ordering that was drawn at the beginning.
-The fourth step is the essential new addition.
-It starts with agent 1 and searches—without loss, from the bottom up—for an agent $j$ whose RP assignment, $x_{j}$, is preferred by agent 1 to their own assignment, $x_{1}$, more than it is so by agent $j$.
-If such an agent exists, the algorithm implements the swap, removes the two agents and the respective items, then repeats this step with the remaining agents and items.
-If not, then agent 1 and $x_{1}$ are removed before the step is re-applied.
-This forced exit is important in two ways.
-First, it prevents potentially indefinite loops, for example in cases where the RP allocation is already intensity-efficient.
-Second, as is clarified below, it ensures that the computational complexity of the whole process falls in the manageable quadratic class.
-Proposition 6.
-If $S\in\mathcal{\widehat{S}}$, and either $n=3$ or $S$ is effectively homeo-monotonic, then the algorithm ends in at most ${n}\choose{2}$} steps—hence in $O(n^{2})$ time—and either produces the Pareto-efficient allocation of its RP-implementing step or one that intensity-dominates it.
-The allocation delivered by this algorithm is not intensity-efficient in general.
-This is so because the forced exit of agents and items in the fourth step means that not all pairwise possibilities for intensity-dominance-driven improvements are considered in the transition from the RP-mandated Pareto-efficient allocation in Step 2 to the final allocation.
-Clearly, however, the algorithm improves upon RP and Pareto-efficiency—even if partially so—in the direction of intensity-efficiency, without adding intractable computational complexity relative to RP.
-It is presently unknown if this algorithm can be modified to deliver an intensity-efficient allocation—whenever one is known to exist—without requiring factorial search over all possible priority orderings over agents in Step 2.555Suppose, instead, that an intensity-positional allocation is the welfare-improving criterion of interest.
-Such an allocation exists for any strict intensity profile and can be identified via simple summation once agents’ ordinal intensities are elicited.
-Owing to these facts, it is a straightforward implication of the discussion around Proposition 6 that, in order for the algorithm’s RP allocation to be converted into an intensity-positional one, it suffices to replace the last two steps by any deterministic or random criterion that selects one among the possibly many intensity-positional allocations corresponding to the given profile.
-Moreover, by Proposition 5 and the linear complexity of the RP algorithm (Manlove, 2013, Section 6.2), this process terminates in at most (n1)(n+1)(n-1)(n+1) as well.
-In light of the discussion preceding the statement of Proposition 6, it is clear that RP can be replaced by PS in Step 2, depending on the desired resolution between strategy-proofness and ordinal-efficiency.
-Regardless of whether RP, PS or some other mechanism is used to elicit the preference component of the input intensity profile in the manner dictated in the proof of Proposition 5, however, one is naturally interested to know if, when operating under the matching rule defined by the above algorithm, agents are incentivized to report their intensities truthfully.
-The answer is negative.
-To see this, suppose $n=3$ and let $S\in\mathcal{\widehat{S}}$ be such that $a\succ_{1}b\succ_{1}c$; $a\succ_{2}b\succ_{2}c$; $c\succ_{3}b\succ_{3}a$; $(a,b)\dot{\succ}_{1}(b,c)$; $(b,c)\dot{\succ}_{2}(a,b)$; $(c,b)\dot{\succ}_{3}(b,a)$.
-The possible orderings over $N$ are: 1.
-$v_{1}\gg v_{2}\gg v_{3}$; 2.
-$v_{1}\gg v_{3}\gg v_{2}$; 3.
-$v_{2}\gg v_{1}\gg v_{3}$; 4.
-$v_{2}\gg v_{3}\gg v_{1}$; 5.
-$v_{3}\gg v_{1}\gg v_{2}$; 6.
-$v_{3}\gg v_{2}\gg v_{1}$.
-The two Pareto-efficient allocations are $(a,b,c)$, $(b,a,c)$, and the former intensity-dominates the latter.
-Suppose the algorithm’s second step delivers this intensity-efficient allocation (under RP, this happens with prob.
-$1\over 2$}).
-If intensities are reported truthfully, there is no order $\gg$ under which the algorithm’s subsequent steps alter it.
-But if $(b,c)\dot{\succ}^{\prime}_{2}(a,b)$ is the second agent’s misreported intensity, then under the 3rd, 4th and 6th $\gg$-orders (i.e.
-with prob.
-$1\over 2$}) the algorithm reassigns $b$ to $v_{1}$ and $a$ to $v_{2}$, making the latter agent better off.
-Interestingly, the incentives to misreport one’s intensities in this way, namely toward increasing one’s chances of getting their most preferred alternative, act in the opposite direction to the incentives toward misreporting one’s preferences in the Boston/Immediate-Acceptance mechanism of the distinct two-sided (School Choice) matching problem.
-Specifically, as has been discussed by Pathak and Sönmez (2008); Abdulkadiroğlu, Che, and Yasuda (2011) and other authors, under this mechanism students may (and often do) submit their second- or third-ranked school as their supposedly most preferred one if their actual top-ranked school is very popular, out of concern that in the event they are not assigned to it they may end up at one of their lower-ranked schools if their second or third most preferred ones have in the meantime filled up.
-Abdulkadiroğlu et al.
-(2011), however, followed by Pycia (2011), Featherstone and Niederle (2016) and other studies, argued that this drawback on incentive grounds of the Boston mechanism may be outweighed by the total welfare gains associated with such misreporting once the students’ intensities are accounted for.
-In line with the prior relevant literature, intensity comparisons in that paper’s Bayesian-Nash equilibrium analysis are derived from intra- and inter-personal differences in cardinal utility functions that emerge from expected-utility preferences over lotteries over schools.
-In that framework too, finally, the intensity information that is conveyed to the social planner is assumed to exist without having been elicited incentive-compatibly, and it is unclear if such elicitation is possible.
-In relation to this, Pycia and Ünver (2025) recently showed that, in environments without monetary transfers such as the one considered above, only ordinal mechanisms are group strategy-proof.
+6 Concluding Remarks
+In this paper we study the classic one-to-one assignment problem in an environment where, in addition to standard ordinal preferences, agents are also assumed to have ordinal preference intensities.
+These are internally consistent comparisons between preference improvements that do not require any quantification of the kind that is implied by cardinal utility functions.
+Two interpersonal comparability assumptions for such ordinal intensities are introduced and motivated.
+Building on this analytical environment, which provides a novel and potentially fertile middle ground between purely ordinal and fully cardinal environments for the study of allocation problems, we introduce and analyze the concepts of intensity-efficient and intensity-positional allocations.
+While distinct and with different points of strength and weakness, they share the feature that they refine the set of Pareto-efficient allocations in ways that incorporate this cognitively less demanding information about agents’ intensities into the allocation process with the goal—shared by notions defined in the cardinal-welfarist tradition—of achieving greater distributive justice.
 Appendix
@@
 Let $N^{*}\subsetneq N$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$.
-By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$.
+By Strictness, $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$.
 Suppose $N^{*}=\emptyset$.
@@
 By $s_{l}(y_{l},x_{l})>0$ and $|A|=|N|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$.
-By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$.
+By Strictness and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$.
 Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in N^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $N^{i>1}$ and $N^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$.
@@
 Proof.
-By definition,
-$\begin{array}[]{ccccc}r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})&\geq&r_{l}(a_{j})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{j},a_{k})&\Longleftrightarrow\\ s_{l}(a_{i},a_{k})&\geq&s_{l}(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow\\ (a_{i},a_{k})&\dot{\succsim}_{l}&(a_{j},a_{k})\quad\forall\;a_{k}\in A&\Longleftrightarrow&\\ (a_{i},a_{j})&\dot{\succsim}_{l}&(a_{j},a_{i})&\Longleftrightarrow&a_{i}\succsim_{l}a_{j},\end{array}$
-where the penultimate equivalence immediately follows from $(\mathbf{4_{\mathit{D}}})$, which in turn is implied by the postulate that $s_{l}$ represents $\dot{\succsim}_{l}$ in profile $S\in\mathcal{\widetilde{S}}$.
-∎
+Let $w^{l}\in A$ be any $\succsim_{l}$-minimal alternative, i.e. such that $a\succsim_{l}w^{l}$ for all $a\in A$. We first show that
+$r_{l}(a)=\max_{b\in A}s_{l}(a,b)=s_{l}(a,w^{l})\qquad\text{for all }a\in A.$
+Fix $a,b\in A$. If $a\succsim_{l}b$, then $s_{l}(a,b)\geq 0$. Since $w^{l}$ is $\succsim_{l}$-minimal, $b\succsim_{l}w^{l}$ is true, so $s_{l}(b,w^{l})\geq 0$ too. By lateral consistency,
+$s_{l}(a,w^{l})\geq\max\{s_{l}(a,b),\,s_{l}(b,w^{l})\}\geq s_{l}(a,b).$
+If, instead, $b\succ_{l}a$, then $s_{l}(a,b)<0$, while $a\succsim_{l}w^{l}$ implies $s_{l}(a,w^{l})\geq 0$.
+Hence, again $s_{l}(a,w^{l})\geq s_{l}(a,b)$.
+Since this holds for every $b\in A$, it follows that $s_{l}(a,w^{l})$ is the maximum value of $s_{l}(a,\cdot)$ on $A$, i.e.
+$r_{l}(a)=s_{l}(a,w^{l})$ for all $a\in A$.
+Now let $a_{i},a_{j}\in A$.
+Then
+$r_{l}(a_{i})\geq r_{l}(a_{j})$ $\Longleftrightarrow$ $s_{l}(a_{i},w^{l})\geq s_{l}(a_{j},w^{l})$ (38)
+$\Longleftrightarrow$ $(a_{i},w^{l})\,\dot{\succsim}_{l}\,(a_{j},w^{l}).$ (39)
+We will show that $(a_{i},w^{l})\,\dot{\succsim}_{l}\,(a_{j},w^{l})$ $\Leftrightarrow$ $(a_{i},a_{j})\,\dot{\succsim}_{l}\,(a_{j},a_{i})$.
+Suppose $(a_{i},w^{l})\,\dot{\succsim}_{l}\,(a_{j},w^{l})$.
+This is equivalent to $s_{l}(a_{i},w^{l})\geq s_{l}(a_{j},w^{l})$.
+By the definition of $w^{l}$, moreover, $s_{l}(a_{j},w^{l})\geq 0$.
+Suppose to the contrary that $(a_{j},a_{i})\,\dot{\succ}_{l}\,(a_{i},a_{j})$.
+This is equivalent to $s_{l}(a_{j},a_{i})>0>s_{l}(a_{i},a_{j})$.
+By (6), $s_{l}(a_{j},a_{i})>0$ and $s_{l}(a_{i},w^{l})\geq 0$ implies $s_{l}(a_{j},w^{l})>s_{l}(a_{i},w^{l})$, which is equivalent to $(a_{j},w^{l})\,\dot{\succ}_{l}\,(a_{i},w^{l})$.
+This contradicts our original postulate.
+Conversely, suppose $(a_{i},a_{j})\,\dot{\succsim}_{l}\,(a_{j},a_{i})$, and assume to the contrary that $(a_{j},w^{l})\,\dot{\succ}_{l}\,(a_{i},w^{l})$; that is, $s_{l}(a_{i},a_{j})\geq s_{l}(a_{j},a_{i})$ and $s_{l}(a_{j},w^{l})>s_{l}(a_{i},w^{l})$.
+By (6) and a similar reasoning we are now led to the absurd conclusion that $s_{l}(a_{i},w^{l})>s_{l}(a_{j},w^{l})$.
+Now, by the definition of the preference relation induced by $\dot{\succsim}_{l}$, $(a_{i},a_{j})\,\dot{\succsim}_{l}\,(a_{j},a_{i})$ is equivalent to $a_{i}\succsim_{l}a_{j}$.
+Therefore, from (39) we get
+$r_{l}(a_{i})\geq r_{l}(a_{j})\Longleftrightarrow a_{i}\succsim_{l}a_{j},$
+so $r_{l}$ is an ordinal utility representation of the preference relation induced by $\dot{\succsim}_{l}$. ∎
 Considering now the stronger postulate whereby $\dot{\succsim}_{l}\in\mathcal$$\widetilde${I}} is linear, note first that this implies that the relation $\succsim_{l}$ induced by $\dot{\succsim}_{l}$ is a linear order.
@@
 Since there are $n$ alternatives in $A$, the above shows that there are n1n-1 intensity-equivalence classes $a_{i},a_{j}]\in(A\times A)_{\dot{\sim}_{l}}$ such that $a_{i}\succ_{l}a_{j}$.
-By ($1$), $r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})$.
+By (11), $r_{l}(a_{i})\equiv\max\limits_{a_{k}\in A}s_{l}(a_{i},a_{k})$.
 This and the above-postulated $\succ_{l}$ ordering on AA jointly imply
@@
 Proof of Proposition 3.
-By the definitions of $b_{l}$ and $r_{l}$ in (10) and (11), for any $S\in\mathcal{\widehat{S}}$ and every $l\leq n$ there is a strictly increasing function $f_{l}:\{0,1,2,\ldots,n-1\}\rightarrow\{0,1,2,\ldots,{{n}\choose{2}}\}$ such that, for all $a\in A$,
-$r_{l}^{S}(a)$ $\equiv$ $f_{l}(b_{l}^{P_{S}}(a))$
-Hence, for $x,y\in\mathcal{A}$ it holds that
-$\sum_{i=1}^{n}b_{l}^{P_{S}}(x_{i})>\sum_{i=1}^{n}b_{l}^{P_{S}}(y_{i})$ $\Longrightarrow$ $\sum_{i=1}^{n}r_{l}^{S}(x_{i})>\sum_{i=1}^{n}r_{l}^{S}(y_{i})$ (39)
-To prove the first claim, suppose to the contrary that $x\in\mathsf{IP}(S)$ but $x\not\in\mathsf{Bo}(S)$.
-Then, there must be $y\in\mathcal{A}$ such that $\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(y_{i})>\sum\limits_{i=1}^{n}b_{l}^{P_{S}}(x_{i})$.
-By (39), this contradicts the postulate $x\in\mathsf{IP}(S)$.
-The second claim is proved in the main text.
+Let $x$ be intensity-positional at $S\in\widehat{S}$.
+Suppose, toward a contradiction, that $x$ is not Pareto-efficient.
+Then there exists some $y\in\mathcal{A}$ that Pareto-dominates $x$ at the induced preference profile $P_{S}$.
+That is, $y_{i}\succsim_{i}x_{i}$ for all $i\leq n$, with $y_{j}\succ_{j}x_{j}$ for some $j\leq n$.
+For each agent $i\leq n$, the Borda score function $b_{i}^{P_{S}}$ is an ordinal utility representation of $\succ_{i}$, and so is the intensity-rank score $r_{i}^{S}$ (Lemma 5).
+Hence there exists a strictly increasing function $f_{i}:\{0,1,\dots,n-1\}\to\{0,1,\dots,{n\atopwithdelims( )2}\}$ such that $r_{i}^{S}(a)=f_{i}\bigl(b_{i}^{P_{S}}(a)\bigr)$ for all $a\in A$.
+Since $y_{i}\succsim_{i}x_{i}$ for all $i$ and $y_{j}\succ_{j}x_{j}$ for some $j$, it follows that $b_{i}^{P_{S}}(y_{i})\geq b_{i}^{P_{S}}(x_{i})$ for all $i$, with strict inequality for at least one $i$.
+Because each $f_{i}$ is strictly increasing, we obtain $r_{i}^{S}(y_{i})\geq r_{i}^{S}(x_{i})$ for all $i$, with strict inequality for at least one $i$.
+Therefore, $\sum\limits_{i=1}^{n}r_{i}^{S}(y_{i})>\sum\limits_{i=1}^{n}r_{i}^{S}(x_{i})$, contradicting the assumption that $x$ is intensity-positional at $S$.
+Hence, any intensity-positional allocation is Pareto-efficient.
 $\square$
-Proof of Proposition 4. In the main text.
-Proof of Proposition 5.
-It suffices to note that (4) and $(\mathbf{5}_{D})$ together imply that any $\dot{\succsim}_{l}\in\mathcal$$\widehat${I}} contains $|\dot{\succsim}_{l}|=2{n\choose 2}+n=(n-1)(n+1)$ comparisons between pairs in $A\times A$.
-$\square$
-Proof of Proposition 6.
-Let $x$ be the Pareto-efficient allocation that is found in Step 2 under the random priority ordering $\gg$ drawn in the algorithm’s first step.
-In the worst-case scenario, Step 4 always—with the possible exception of the agent ranked second-last by $\gg$—performs exhaustive search and removes one agent and one object in every iteration: Starting with $v_{1},x_{1}$, after n1n-1 searches it removes $v_{1},x_{1}$.
-Continuing with $v_{2},x_{2}$, after n2n-2 searches it removes $v_{2},x_{2}$.
-$\cdots$ Continuing with $v_{n-1},x_{n-1}$, after one search it removes both $v_{n-1},x_{n-1}$ and $v_{n},x_{n}$, either by implementing the swap $x_{n-1}\leftrightarrow x_{n}$ or otherwise.
-Worst-case performance is therefore associated with a total number of searches equalling
-$\begin{array}[]{llclcllll}\sum\limits_{m=1}^{n-1}(n-m)&=&n^{2}-n-\sum\limits_{m=1}^{n-1}m&=&n^{2}-\sum\limits_{m=1}^{n}m&=&&&\\ &&&&&&\\ &&n^{2}-\dfrac{n}{2}(n+1)&=&\dfrac{1}{2}n(n-1)&=&\displaystyle{{n\choose k}}&&\end{array}$
-Given that a Pareto-efficient allocation can be found by RP in Step 2 in $O(n)$ time (Manlove, 2013, Section 6.2), it follows that the algorithm terminates in at most ${n}\choose{2}$} steps, hence in $O(n^{2})$ time.
-It remains to be shown that, if the final allocation is $y\neq x$, then $yDx$.
-Suppose $n=3$ first.
-Write $x:=(x_{1},x_{2},x_{3})$.
-By the definition of Step 4, the transition from $x$ to $y$ involves a single pairwise swap.
-Thus, $y$ must be one of the following: (i) $(x_{1},x_{3},x_{2})$, implying $s_{2}(x_{3},x_{2})>s_{3}(x_{3},x_{2})>0$ or $s_{3}(x_{2},x_{3})>s_{2}(x_{2},x_{3})>0$; (ii) $(x_{3},x_{2},x_{1})$, implying $s_{1}(x_{3},x_{1})>s_{3}(x_{3},x_{1})>0$ or $s_{3}(x_{1},x_{3})>s_{1}(x_{1},x_{3})>0$; (iii) $(x_{2},x_{1},x_{3})$, implying $s_{2}(x_{1},x_{2})>s_{1}(x_{1},x_{2})>0$ or $s_{1}(x_{2},x_{1})>s_{2}(x_{2},x_{1})>0$.
-In each case, $yDx$.
-Now suppose $n>3$ and $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
-Denote by $(N^{i})_{i=1}^{k}$ and $(A^{i})_{i=1}^{k}$ the respective partitions of $N$ and $A$ into a (possibly empty) top-preference diverse set $N^{1}$ with respect to $A^{1}$ and top-preference-coincident sets $N^{i>1}$ with respect to $A^{i>1}$, $i\leq k$.
-By the properties of these sets and profile, and by the fourth step of the algorithm, irrespective of the $\gg$-order on $N$ there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i},v_{j}\in N^{1}$.
-Thus, $y_{i}=x_{i}$ for all $v_{i}\in N^{1}$.
-By the properties of $(N^{i>1},A^{i>1})$, moreover and the instruction in Step 4, there can be no swaps between $x_{i}$ and $x_{j}$ for $v_{i}\in N^{l>1}$ and $v_{j}\in N^{m>1}$ when $m,l\in\{2,\ldots,k\}$ are distinct.
-It follows that $y$ differs from $x$ by at most $n\over 2$} or $n-1\over 2$} pairwise swaps/transpositions if $n$ is even or odd, respectively, with each such $x_{i}\leftrightarrow x_{j}$ swap occurring within the same top-preference-coincident pair $(N^{l>1},A^{l>1})$.
-By Step 4, finally, every such swap reassigns the—by Pareto-efficiency of xx—jointly preferred item within the relevant pair in the direction of intensity-dominance.
-Hence, $yDx$.
-$\square$
+Proof of Proposition 4. In the main text. $\square$